### v1

# The Critical LYZ equation in Kähler Geometry

Abstract.
We establish the existence of smooth solutions for the LYZ equation at the critical phase $\theta=(n-2)\frac{\pi}{2}$, thereby solving the critical case of a problem posed by Collins-Jacob-Yau [7] and Li [25] concerning the solvability for phase $\theta\leq(n-2)\frac{\pi}{2}$. As applications, we solve the 3D Hessian equation $\sigma_{2}=1$ and the 4D Hessian quotient equation $\sigma_{3}=\sigma_{1}$ under weaker assumptions than previously required [34, 36].
Keywords: critical LYZ equation; subsolution; uniform estimates; blow-up argument
1. Introduction
In Kähler geometry, the LYZ (Leung-Yau-Zaslow) equation, also known as the dHYM (deformed Hermitian Yang-Mills) equation, is a fully nonlinear elliptic PDE that arises from mirror symmetry [24, 30]. This equation is of current interest in both mathematics and physics.
Let $(M,\omega)$ be a compact Kähler manifold and $\chi$ a closed real $(1,1)$-form. Suppose that $\mathcal{Z}=\int_{M}(\chi+\sqrt{-1}\omega)^{n}\neq 0$ and let $\hat{\theta}$ be the principal argument of $\mathcal{Z}$. The LYZ equation seeks a smooth function $u:M\rightarrow\mathbb{R}$ such that
$\mathrm{Im}\big(e^{-\sqrt{-1}\hat{\theta}}(\chi_{u}+\sqrt{-1}\omega)^{n}\big)=0.$ (1)
Let $\lambda=(\lambda_{1},\cdots,\lambda_{n})$ be the eigenvalues of $\chi_{u}$ with respect to $\omega$ and define the special Lagrangian operator as
$\theta_{\omega}(\chi_{u}):=\sum\limits_{i=1}^{n}\arctan\lambda_{i}.$
Then equation (1) can be rewritten in the form
$\theta_{\omega}(\chi_{u})=\theta$ (2)
where $\theta=n\frac{\pi}{2}-\hat{\theta}$. We refer to (2) as the supercritical LYZ equation if $\theta\in\big((n-2)\frac{\pi}{2},n\frac{\pi}{2}\big)$, and as the critical LYZ equation if $\theta=(n-2)\frac{\pi}{2}$, i.e., $\hat{\theta}=\pi$.
A key requirement for solving the LYZ equation is the existence of a subsolution: A smooth function $\underline{u}$ satisfying
$\min\limits_{1\leq j\leq n}\sum\limits_{\begin{subarray}{c}i=1\\ i\neq j\end{subarray}}^{n}\mathrm{arctan}\lambda_{i}(\chi_{\underline{u}})>\theta-\frac{\pi}{2}.$ (3)
This notion of subsolution is equivalent to that introduced by Guan [12] and Székelyhidi [36].
1.1. Previous Results
The 2-dimensional case of the LYZ equation was solved by Jacob-Yau [22], who also obtained partial results in higher dimensions. Collins-Jacob-Yau [7] then solved the supercritical case under the assumption of a subsolution $\underline{u}$ and an additional condition $\theta_{\omega}(\chi_{\underline{u}})>(n-2)\frac{\pi}{2}$. The authors [11] introduced a geometric flow to solve the equation under the same conditions. The extra condition was subsequently removed by Pingali [33] for $n=3$, by Lin [27] for $n=3,4$, and finally by Lin [26] in full generality. An alternative proof, based on Chen’s significant result [3], was given by Sun [35]. For further recent developments, we refer to [2, 4, 5, 6, 8, 9, 13, 14, 15, 18, 19, 20, 21, 23, 28, 37].
1.2. Our main results
Collins-Jacob-Yau [7] suggested studying the LYZ equation when $\theta\leq(n-2)\frac{\pi}{2}$. This was later explicitly posed as a question by Li [25], as resolving it is essential for the LYZ equation to induce a Bridgeland stability condition on $D^{b}\text{Coh}(M)$.
In this paper, we solve the critical case; that is, we establish the existence of smooth solutions for the critical LYZ equation:
$\theta_{\omega}(\chi_{u})=(n-2)\frac{\pi}{2}.$ (4)
In this case, the principal argument of $\int_{M}(\chi+\sqrt{-1}\omega)^{n}$ is $\pi$, and the subsolution condition becomes
$\min\limits_{1\leq j\leq n}\sum\limits_{\begin{subarray}{c}i=1\\ i\neq j\end{subarray}}^{n}\mathrm{arctan}\lambda_{i}(\chi_{\underline{u}})>\frac{n-3}{2}\pi.$ (5)
Theorem 1.1.
Let $(M,\omega)$ be a compact Kähler manifold and $\chi$ a closed real (1,1)-form such that the principal argument of $\int_{M}(\chi+\sqrt{-1}\omega)^{n}$ is $\pi$. If there exists a smooth function $\underline{u}$ satisfying (5), then the critical LYZ equation (4) admits a unique smooth solution u with $\sup_{M}u=0$.
Remark 1.1.
For the special Lagrangian equation in $\mathbb{R}^{n}$, interior Hessian estimates for critical and supercritical cases were proved by Warren-Yuan [41] and Wang-Yuan [39]. For $\theta<(n-2)\frac{\pi}{2}$, $C^{1,\alpha}$ solutions were constructed by Nadirashvili-Vlădut [32] and Wang-Yuan [38]. Very recently, Mooney-Savin [31] obtained a Lipschitz solution that is not $C^{1}$.
To solve the critical LYZ equation (4), we consider the family of equations:
$\theta_{\omega}(\chi_{u^{t}}+t\omega)=\theta(t).$ (6)
We verify that there exists a uniform constant $C>0$ such that
$\theta(t)\in((n-2)\frac{\pi}{2}+Ct,(n-2)\frac{\pi}{2}+2Ct)$
for small $t>0$ (see Lemma 2.3 in Section 2). Hence equations (6) are the supercritical LYZ equations. Then we prove the subsolution $\underline{u}$ of the critical LYZ equation (4) is also a subsolution of equations (6) for small $t>0$. Thus we can apply Lin’s result [26] (see also Sun [35] ) to obtain the smooth solution $u^{t}$ to equation (6) with $\sup_{M}u^{t}=0$.
To prove Theorem 1.1, we establish uniform $C^{2,\alpha}$ estimates for $u^{t}$ that are independent of $t$ and $(\theta(t)-(n-2)\frac{\pi}{2})^{-1}$.
Theorem 1.2.
Assume the same conditions as in Theorem 1.1 hold. Let $u^{t}$ be the smooth solution of (6) with $\sup_{M}u^{t}=0$. Then there exist uniform constants $C$ and $t_{0}$ depending only on $(M,\omega,\chi)$ but independent of $t$ such that for any $t\in(0,t_{0})$, we have
$|u^{t}|_{C^{2,\alpha}(M)}\leq C.$ (7)
According to Collins-Jacob-Yau [7], $\sup_{M}|u^{t}|$ is uniform bounded. Their results then imply that $C^{2,\alpha}$ estimates can be reduced to prove the uniform complex Hessian estimates. Our main contribution is the proof of the uniform complex Hessian estimates of Hou-Ma-Wu type
$\sup\limits_{M}|\sqrt{-1}\partial\bar{\partial}u|_{\omega}\leq C(1+\sup\limits_{M}|\partial u|^{2}_{\omega})$ (8)
and the uniform gradient estimates via the Dinew-Kolodziej blow-up argument [10].
However, the situation here is more complicated. If the gradient estimates fail, we obtain a $C^{1}$ bounded non-constant function $v$ on $\mathbb{C}^{n}$ which, in the sense of currents, is a weak solution to one of the following equations: $\sigma_{n}=0$, $\sigma_{n-1}=0$, or $\sigma_{n-1}+\sigma_{n}=0$. The last case is new and requires a new Liouville-type theorem. Note that $v+|z|^{2}$ is plurisubharmonic and $\Delta v$ is uniformly bounded in the weak sense (see Remark 4.2). When we follow the argument in [10, 36], we find the first case can be handled similarly. For the second case, we need to prove the standard mollification $v]_{r}$ of $v$ and $\frac{1}{3}[(v^{\epsilon})^{2}]_{r}-\tau_{0}|z|^{2}$ are subsolutions, where $v^{\epsilon}=[v]_{\epsilon}$ and $\tau_{0}>0$ is small enough. To address this issue, our starting point is that a smooth function $v$ satisfies equation $\sigma_{n-1}+\sigma_{n}=0$ (resp. $\geq 0$) on $\mathbb{C}^{n}$ is equivalent to the function $v+|z_{n+1}|^{2}$ satisfies equation $\sigma_{n}=0$ (resp. $\geq 0$) on $\mathbb{C}^{n+1}$, where $z_{n+1}$ is a new complex variable.
Finally, we find some applications of the main result. In general, the critical LYZ equation has the form
$\sigma_{n-1}(\chi_{u})=\sigma_{n-3}(\chi_{u})-\sigma_{n-5}(\chi_{u})+\cdots,$
where $\sigma_{k}(\chi_{u})$ is the $k$-Hessian operator of $\chi_{u}$ with respect to the metric $\omega$ (see (11)).
When $n=3$, the critical LYZ equation is just the 2-Hessian equation
$\sigma_{2}(\chi_{u})=1.$
As a consequence of our result, we can solve the 2-Hessian equation in dimension 3 under weaker conditions than those by Sun [34].
Corollary 1.1.
Let $(M,\omega)$ be a compact Kähler manifolds with dimension $n=3$ and $\chi$ be a closed real $(1,1)$ form satisfying $\chi\wedge\omega>0$ as a positive $(2,2)$ form and
$3\int_{M}\chi^{2}\wedge\omega=\int_{M}\omega^{3},\ \ \int_{M}\chi^{3}<3\int_{M}\chi\wedge\omega^{2}.$
Then there exists a unique smooth solution $u$ solving
$\sigma_{2}(\chi_{u})=1,$
$\chi_{u}\in\Gamma_{2}(M),\ \sup_{M}u=0.$
Note that Hou-Ma-Wu [17] proved the complex Hessian estimate of Hou-Ma-Wu type for the $k$-Hessian equation $\sigma_{k}(\omega_{u})=f$ on a compact Kähler manifold. Then Dinew-Kolodziej [10] obtained the gradient estimate by a Liouville theorem for $\sigma_{k}=0$ in $\mathbb{C}^{n}$ and thus solved the equation. The Hermitian case was solved independently by Székelyhidi [36] and Zhang [43].
When $n=4$, the critical LYZ equation is the Hessian quotient equation
$\sigma_{3}(\chi_{u})=\sigma_{1}(\chi_{u}).$
Another consequence of our result is to solve the above Hessian quotient equation in dimension 4 under weaker conditions than the one by Sun [34] and Székelyhidi [36].
Corollary 1.2.
Let $(M,\omega)$ be a compact Kähler manifold with dimension $n=4$. Let $\chi$ be a closed real $(1,1)$-form satisfying $3\chi^{2}\wedge\omega-\omega^{3}>0\ \text{as a positive $(3,3)$ form}$ and
$\int_{M}\chi^{3}\wedge\omega=$ $\int_{M}\chi\wedge\omega^{3},$
$\int_{M}\chi\wedge\omega^{3}>$ $0,\ \mathrm{Re}\int_{M}(\chi+\sqrt{-1}\omega)^{4}<0.$
Then there exists a unique smooth function $u$ solving
$\sigma_{3}(\chi_{u})=\sigma_{1}(\chi_{u}),$ (9)
$\chi_{u}\in\Gamma_{3}(M),\ \sup_{M}u=0.$
The Hessian quotient equation $\sigma_{k}(\chi_{u})=c\sigma_{l}(\chi_{u})$ with $0\leq l<k\leq n$ was solved by Sun [34] and Székelyhidi [36] by different continuity methods.
The organization of this paper is as follows. In Section 2, we provide preliminaries on $k$-Hessian operators, the special Lagrangian operator $\theta_{\omega}$, and the approximating family $\theta(t)$. We also give the $C^{0}$-estimate for $u^{t}$ using results from [7]. In Section 3, we first proves a key inequality implied by the subsolution condition, then we establish the uniform complex Hessian estimate of Hou-Ma-Wu-type which is independent of $(\theta(t)-(n-2)\frac{\pi}{2})^{-1}$). In Section 4, based on the complex Hessian estimate, we prove the gradient estimate by a blow up argument. Then we prove the main theorems. In Section 5, we apply the main results to the 3-dimensional 2-Hessian equation and 4-dimensional Hessian quotient equation.
2. preliminaries
2.1. The k-Hessian operators and the special Lagrangian operator
For any vector $\lambda\in\mathbb{R}^{n}$ and $1\leq k\leq n$, we define
$\sigma_{k}(\lambda)=\sum\limits_{1\leq i_{1}<\cdots<i_{k}\leq n}\lambda_{i_{1}}\cdots\lambda_{i_{k}}.$ (10)
For a real $(1,1)$-form $\alpha$, the $k$-Hessian operator of $\alpha$ with respect to $\omega$ is defined by
$\sigma_{k}(\alpha):=C_{n}^{k}\frac{\alpha^{k}\wedge\omega^{n-k}}{\omega^{n}}.$ (11)
In local coordinates, let $\alpha=\sqrt{-1}\alpha_{i\bar{j}}dz_{i}\wedge d\bar{z}_{j}$ and $\omega=\sqrt{-1}g_{i\bar{j}}dz_{i}\wedge d\bar{z}_{j}$, we have $\sigma_{k}{(\alpha)}=\sigma_{k}(\lambda(g^{j\bar{l}}\alpha_{i\bar{l}}))$. For simplicity, we also write $\sigma_{k}(g^{j\bar{l}}\alpha_{i\bar{l}})$ to represent $\sigma_{k}(\lambda(g^{j\bar{l}}\alpha_{i\bar{l}}))$ .
We need the following convex cones:
$\Gamma_{k}:=$ $\{\lambda\in\mathbb{R}^{n}\mid\sigma_{m}(\lambda)>0,1\leq m\leq k\},$
[…]
We have the concavity property for Hessian operators.
Lemma 2.1.
The function ${\sigma_{n-1}^{\frac{1}{n-1}}}(\lambda)$ is concave in $\Gamma_{n-1}$. The function ${\sigma_{n-1}^{\frac{1}{n-1}}}(A)$ is concave in $\mathcal{M}(\Gamma_{n-1})$.
Let
$\theta(\lambda):=\sum_{i=1}^{n}\arctan\lambda_{i}\quad\textup{for}\ \lambda=(\lambda_{1},\dots,\lambda_{n})\in\mathbb{R}^{n}$
and
$\Gamma^{\tau}:=\{\lambda\in\mathbb{R}^{n}\mid\theta(\lambda)>\tau\}\subset\mathbb{R}^{n}\quad\textup{for $\tau\in[(n-2)\frac{\pi}{2},n\frac{\pi}{2})$}.$
We have the following useful facts.
Lemma 2.2 (Yuan [42], Wang-Yuan [40]).
If $\theta(\lambda)\geq\tau\in[(n-2)\frac{\pi}{2},n\frac{\pi}{2})$ for $\lambda=(\lambda_{1},\dots,\lambda_{n})$ with $\lambda_{1}\geq\lambda_{2}\geq\cdots\geq\lambda_{n}$, then we have
  1. (1)
    $\lambda_{1}\geq\lambda_{2}\geq\cdots\geq\lambda_{n-1}>0$, and $\lambda_{n-1}\geq|\lambda_{n}|$,
  2. (2)
    $\lambda_{1}+(n-1)\lambda_{n}\geq 0$,
  3. (3)
    $\lambda\in\Gamma_{n-1}$, and
  4. (4)
    $\Gamma^{\tau}$ is convex.
$2.$ $C^{0}$-estimates
Let $\hat{\theta}(t)$ be the principal argument of the complex number $\int_{M}(\chi+t\omega+\sqrt{-1}\omega)^{n}$ and $\theta(t)=(n-2)\frac{\pi}{2}-\hat{\theta}(t)$. We first prove $\hat{\theta}(t)$ is sufficiently close to $\pi$ for sufficiently small $t>0$.
Lemma 2.3.
If $\underline{u}$ is the subsolution of the critical LYZ equation (4), then there exists uniform positive constants $t_{0}$ and $C$ depending only on $\chi$ and $\omega$ such that $\hat{\theta}(t)\in(\pi-2Ct,\pi-Ct)$ for any $t\in(0,t_{0})$.
Proof.
By direct computation, we have
$(\chi+t\omega+\sqrt{-1}\omega)^{n}=(\chi+\sqrt{-1}\omega)^{n}+tn(\chi+\sqrt{-1}\omega)^{n-1}\wedge\omega+O(t^{2}).$
Since
$\mathrm{Re}\int_{M}(\chi+\sqrt{-1}\omega)^{n}<0,$
for sufficiently small $t$, we have
$\mathrm{Re}\int_{M}(\chi+t\omega+\sqrt{-1}\omega)^{n}<\mathrm{Re}\int_{M}(\chi+\sqrt{-1}\omega)^{n}+Ct<0.$ (12)
Next we want to prove the imaginary part is positive. The subsolution condition implies the positivity
$\mathrm{Im}(\chi_{\underline{u}}+\sqrt{-1}\omega)^{n-1}>0.$
Then for sufficiently small $t>0$, we get
$\mathrm{Im}\int_{M}(\chi+\sqrt{-1}\omega)^{n-1}\wedge\omega=\mathrm{Im}\int_{M}(\chi_{\underline{u}}+\sqrt{-1}\omega)^{n-1}\wedge\omega>0$ (13)
Thus we get
$\mathrm{Im}\int_{M}(\chi+t\omega+\sqrt{-1}\omega)^{n}=$ $\mathrm{Im}\int_{M}(\chi+\sqrt{-1}\omega)^{n}$
$+t\mathrm{Im}\int_{M}(\chi+\sqrt{-1}\omega)^{n-1}\wedge\omega+O(t^{2})$
$=$ $t\mathrm{Im}\int_{M}(\chi+\sqrt{-1}\omega)^{n-1}\wedge\omega+O(t^{2})$
$>$ $0,$
where in the last inequality we use (13). Then there exists a uniform positive constant $C$ depending only on $\chi$ and $\omega$ such that the principal argument of the complex number $\int_{M}(\chi+t\omega+\sqrt{-1}\omega)^{n}$ satisfies
$\hat{\theta}(t)=\mathrm{arccot}\Big(\frac{\mathrm{Re}\int_{M}(\chi+\sqrt{-1}\omega)^{n}+O(t)}{t\mathrm{Im}\int_{M}(\chi+\sqrt{-1}\omega)^{n-1}\wedge\omega+O(t^{2})}\Big)\in(\pi-2Ct,\pi-Ct).$
Hence by Lemma 2.3, we have, for sufficiently small $t>0$,
$\theta(t)\in\big((n-2)\frac{\pi}{2}+Ct,(n-2)\frac{\pi}{2}+2Ct\big),$ (14)
and then
$\min\limits_{1\leq j\leq n}\sum\limits_{\begin{subarray}{c}i=1\\ i\neq j\end{subarray}}^{n}\mathrm{arctan}\lambda_{i}(\chi_{\underline{u}}+t\omega)>(n-3)\frac{\pi}{2}+2Ct>\theta(t)-\frac{\pi}{2},$ (15)
which means that $\underline{u}$ is also a subsolution of the equations (6) for sufficiently small $t>0$. According to Lin [26] (see also Sun [35]), for every small $t>0$, there exists a smooth solution $u^{t}$ solving the supercritical LYZ equation (6). According to Proposition 3.6 in [7], we get the uniform $C^{0}$ estimates of $u^{t}$.
Lemma 2.4 (Collins-Jacob-Yau [7]).
Let $\underline{u}$ be the subsolution of the critical LYZ equation (4) and $u^{t}$ be the solution of the LYZ equation (6) with $\sup_{M}u^{t}=0$. Then there exist uniform constants $t_{0}$ and $C$ such that for any $t\in(0,t_{0})$, we have
$\sup_{M}|u^{t}|\leq C.$
3. Complex Hessian estimates
3.1. The linearized operator
We first give some useful formulae for the complex Hessian estimate in [11].
Lemma 3.1.
Let $u^{t}$ be the solution of the LYZ equation (6).
(i) The linearized operator $\mathcal{L}$ of the LYZ equation (6) has the form
$\mathcal{{L}}(v)=F^{i\bar{j}}v_{i\bar{j}},$
where $(F^{i\bar{j}})_{n\times n}$ is the inverse matrix of $wg^{-1}w+g$ with $g=(g_{i\bar{j}})_{n\times n}$, $w=(w_{i\bar{j}})_{n\times n}$ for $w_{i\bar{j}}=\chi_{i\bar{j}}+tg_{i\bar{j}}+u_{i\bar{j}}$.
(ii) Differentiating the LYZ equation (6), we have
$F^{i\bar{j}}w_{i\bar{j},p}=0,$ (16)
$F^{i\bar{j}}w_{i\bar{j},p\bar{p}}=F^{i\bar{l}}F^{k\bar{j}}w_{i\bar{j},p}(w_{k\bar{m},\bar{p}}g^{r\bar{m}}w_{r\bar{l}}+w_{k\bar{m}}g^{r\bar{m}}w_{r\bar{l},\bar{p}}),$ (17)
where we used covariant derivatives.
3.2. Key inequalities by the subsolution condition
The subsolution condition is very important to the $C^{2}$-estimates of solutions of fully nonlinear elliptic PDEs, one can see Guan [12] and Székelyhidi [36]. Similar to [12, 36], we prove key inequalities which will be used in the second order estimate. Fortunately, we can get a uniform constant $\delta_{0}$ in the following lemma.
Lemma 3.2.
Let $\mu\in\mathbb{R}^{n}$ satisfy $A(\mu):=\min_{1\leq j\leq n}\sum_{i\neq j}\arctan\mu_{i}>A_{0}>(n-3)\frac{\pi}{2}$. Then for all $\lambda\in\mathbb{R}^{n}$ such that $\theta(\lambda)=\theta(t)$, there exist positive constants $t_{0}$ and $\delta_{0}$ depending only on $n,\mu,A_{0}$ and $C$ in Lemma 2.3 such that for any $t\in(0,t_{0})$, one of the following holds:
  1. (i)
    $\sum_{i=1}^{n}\frac{\mu_{i}-\lambda_{i}}{1+\lambda_{i}^{2}}\geq\delta_{0}\sum_{i=1}^{n}\frac{1}{1+\lambda_{i}^{2}},\ \text{or}$
  2. (ii)
    $\frac{1}{1+\lambda_{j}^{2}}\geq\delta_{0}\sum_{i=1}^{n}\frac{1}{1+\lambda_{i}^{2}}\ \ \text{for}\ 1\leq j\leq n.$
Proof.
Since $A(\mu)>A_{0}>(n-3)\frac{\pi}{2}$, we can choose a small positive constant $\delta$ such that
$A(\mu-2\delta{{\bf{1}}})>\frac{A_{0}+(n-3)\frac{\pi}{2}}{2},$
where ${\bf{1}}=(1,\cdots,1)\in\mathbb{R}^{n}$.
For $\lambda\in\mathbb{R}^{n}$ satisfying $\theta(\lambda)=\theta(t)$, we assume $\lambda_{1}\geq\cdots\geq\lambda_{n}$. We know ${\bf{n}}=(\frac{1}{1+\lambda_{1}^{2}},\cdots,\frac{1}{1+\lambda_{n}^{2}})$ is an inward normal vector of $\Gamma^{\theta(t)}$ at $\lambda$. Let $T_{\lambda}$ be the tangent plane of the hypersurface $\partial\Gamma^{\theta(t)}$ at the point $\lambda$.
Case 1: $\big((\mu-\delta{{{\bf 1}}})-\lambda\big)\cdot{\bf{n}}\geq 0$. In this case, we have
$\sum_{i=1}^{n}\frac{\mu_{i}-\lambda_{i}}{1+\lambda_{i}^{2}}\geq\delta\sum_{i=1}^{n}\frac{1}{1+\lambda_{i}^{2}}.$ (18)
Case 2: $\big((\mu-\delta{{{{{\bf 1}}}}})-\lambda\big)\cdot{\bf{n}}<0$. We can show there exists a positive constant $s$ such that $\mu-\delta{{{{{\bf 1}}}}}+se_{1}\in{\Gamma^{\theta(t)}}.$ Indeed, we have
$\theta(\mu-\delta{{{\bf 1}}}+se_{1})=$ $\arctan(\mu_{1}-\delta+s)+\sum_{i=2}^{n}\arctan(\mu_{i}-\delta)$
$>$ $\arctan(\mu_{1}-\delta+s)+A(\mu-\delta{{{\bf 1}}})$
$\geq$ $\arctan(\mu_{1}-\delta+s)+\frac{1}{2}(A_{0}+(n-3)\frac{\pi}{2})$
$>$ $\theta(t)$
if we choose $s=|\delta-\mu_{1}|+\tan\big(\frac{\pi}{2}-\frac{1}{4}(A_{0}-(n-3)\frac{\pi}{2})\big)$ and assume $0<t<t_{0}$ for $t_{0}=\frac{1}{8C}(A_{0}-(n-3)\frac{\pi}{2})$.
The convexity of $\Gamma^{\theta(t)}$ implies $\mu-\delta{{{\bf 1}}}$ and $\mu-\delta{{{\bf 1}}}+se_{1}$ are on two sides of the tangent plane $T_{\lambda}$. Then there exists $a_{0}\in(0,1)$ such that $a_{0}(\mu-\delta{{{\bf 1}}}+se_{1})+(1-a_{0})(\mu-\delta{{{\bf 1}}})=\mu-\delta{{{\bf 1}}}+a_{0}se_{1}\in\mathrm{T_{\lambda}}$ . We also have $\mu-\delta{{{\bf 1}}}+a_{0}se_{1}+se_{1}-\delta e_{n}\in\Gamma^{\theta(t)}$ which follows from
$\theta(\mu-\delta{{{\bf 1}}}+a_{0}se_{1}+se_{1}-\delta e_{n})$
$=$ $\arctan(\mu_{1}-\delta+a_{0}s+s)$
$+\sum_{i=2}^{n-1}\arctan(\mu_{i}-\delta)+\arctan(\mu_{n}-2\delta)$
$\geq$ $\arctan(\mu_{1}-\delta+a_{0}s+s)+A(\mu-2\delta\bf{1})$
$\geq$ $\arctan(\mu_{1}-\delta+s)+\frac{1}{2}(A_{0}+(n-3)\frac{\pi}{2})$
$>$ $\theta(t).$
Thus we have
$0\leq$ $\big((\mu-\delta{{{\bf 1}}}+a_{0}se_{1}+se_{1}-\delta e_{n}\big)-(\mu-\delta{{{\bf 1}}}+a_{0}se_{1})\big)\cdot\bf{n}$
$=$ $(se_{1}-\delta e_{n})\cdot\bf n$
$=$ $\frac{s}{1+\lambda_{1}^{2}}-\frac{\delta}{1+\lambda_{n}^{2}},$
or $\frac{1}{1+\lambda_{1}^{2}}\geq\delta s^{-1}\frac{1}{1+\lambda_{n}^{2}}$. Since $\lambda_{1}\geq\cdots\geq\lambda_{n-1}\geq|\lambda_{n}|$ by Lemma 2.2, we get
$\frac{1}{1+\lambda_{j}^{2}}\geq\frac{\delta}{sn}\sum_{i=1}^{n}\frac{1}{1+\lambda_{i}^{2}},1\leq j\leq n.$ (19)
In view of (18) and (19), if we choose $\delta_{0}=\frac{\delta}{ns}$, we finish the proof. ∎
By the above lemma, we get the following important lemma and the estimate here does not depend on $(\theta(t)-(n-2)\frac{\pi}{2})^{-1}$.
Lemma 3.3.
Let $\underline{u}$ be the subsolution of the critical LYZ equation (4) and $u^{t}$ be the solution of the LYZ equation (6). Then there exist uniform positive constants $t_{0}$ and $\delta_{0}$ such that for any $t\in(0,t_{0})$ and for any $z\in M$, one of the following holds at $z$:
(i) $\sum\limits_{i,j}F^{i\bar{j}}({\underline{u}_{i\bar{j}}-u^{t}_{i\bar{j}}})\geq\delta_{0}\sum\limits_{i,j}^{n}F^{i\bar{j}}g_{i\bar{j}},\ \text{or}$
(ii) $\mathcal{F}_{min}\geq\delta_{0}\sum\limits_{i,j}^{n}F^{i\bar{j}}g_{i\bar{j}},$ where $\mathcal{F}_{min}$ is the smallest eigenvalue of the matrix $(F^{i\bar{k}}g_{j\bar{k}})_{n\times n}$.
Proof.
Let $\mu=(\mu_{1},\cdots,\mu_{n})$ be the eigenvalues of $\chi_{\underline{u}}$ with respect to $\omega$ with $\mu_{1}\geq\cdots\geq\mu_{n}$. For any $z\in M$, we choose the normal coordinates such that at $z$
$g_{i\bar{j}}=\delta_{ij},\ \chi_{t,u^{t}}=\sum\limits_{i=1}^{n}\lambda_{i}\sqrt{-1}dz^{i}\wedge d\bar{z}^{i}\ \text{with}\ \lambda_{1}\geq\cdots\geq\lambda_{n},$
where $\chi_{t,u^{t}}=\chi+t\omega+\sqrt{-1}\partial\bar{\partial}{u^{t}}$. Then at $z$, we have
$F^{i\bar{j}}=\frac{\delta_{i\bar{j}}}{1+\lambda_{i}^{2}}.$ (20)
Let $\delta_{0}$ and $t_{0}$ be the uniform constants in Lemma 3.2. We use Lemma 3.2 to get the proof.
Case 1: $\sum\limits_{i}F^{i\bar{i}}(\mu_{i}-\lambda_{i})\geq\delta_{0}\sum\limits_{i}F^{i\bar{i}}$. Since $F^{1\bar{1}}\leq\cdots\leq F^{n\bar{n}}$ and $\mu_{1}\geq\cdots\geq\mu_{n}$, we use the Schur-Horn theorem [16] to get
$\sum_{i,j}F^{i\bar{j}}(\underline{u}_{i\bar{j}}-u^{t}_{i\bar{j}})=$ $\sum_{i,j}F^{i\bar{j}}((\chi_{\underline{u}})_{i\bar{j}}-(\chi_{t,u^{t}})_{i\bar{j}})+t\sum_{i}F^{i\bar{i}}$
$\geq$ $F^{i\bar{j}}((\chi_{\underline{u}})_{i\bar{j}}-(\chi_{t,u^{t}})_{i\bar{j}})$
$\geq$ $\sum\limits_{i}F^{i\bar{i}}(\mu_{i}-\lambda_{i})$
$\geq$ $\delta_{0}\sum\limits_{i}F^{i\bar{i}}.$
Case 2: $\sum\limits_{i}F^{i\bar{i}}(\mu_{i}-\lambda_{i})<\delta_{0}\sum\limits_{i}F^{i\bar{i}}$. By Lemma 3.2, we get $\mathcal{F}_{min}=\frac{1}{1+\lambda_{1}^{2}}\geq\delta_{0}\sum\limits_{i}\frac{1}{1+\lambda_{i}^{2}}=\delta_{0}\sum\limits_{i,j}F^{i\bar{j}}g_{i\bar{j}}$. ∎
3.3. The complex Hessian estimate of Hou-Ma-Wu type
We use the auxiliary function introduced by Hou-Ma-Wu [17] who established complex Hessian estimates for the $k$-Hessian equation on compact Kähler manifold to derive the complex Hessian estimates for families of LYZ equations. Here we use the term $w_{i\bar{j}}\xi^{i}\bar{\xi}^{j}$ instead of the largest eigenvalue of $(g^{j\bar{k}}w_{i\bar{k}})_{n\times n}$ and thus the proof seems much more simple. The complex Hessian estimates here do not depend on $({\theta(t)-(n-2)\frac{\pi}{2}})^{-1}$.
Theorem 3.1.
Let $\underline{u}$ be the subsolution of the critical LYZ equation (4) and $u^{t}$ be the solution of the LYZ equation (6) with $\sup_{M}u^{t}=0$. Then there exists uniform positive constants $t_{0}$ and $C$ such that for any $t\in(0,t_{0})$, we have
$\sup_{M}{|\sqrt{-1}\partial\bar{\partial}u^{t}|_{\omega}}\leq C(1+\sup_{M}|\nabla u^{t}|_{\omega}^{2}).$ (21)
Proof.
Without loss of generality, we assume $\underline{u}=0$. Let $w_{i\bar{j}}=\chi_{i\bar{j}}+tg_{i\bar{j}}+u^{t}_{i\bar{j}}$. For simplicity we use $u$ instead of $u^{t}$ in the following proof. Denote $S(T^{1,0}M):=\bigcup_{z\in M}\{\xi\in T^{1,0}_{z}M\,|\,|\xi|_{\omega}=1\}$. We consider the auxiliary function on $S(T^{1,0}M)$
$H(z,\xi)=\log(w_{k\bar{l}}\xi^{k}\bar{\xi}^{l})+\varphi(|\nabla u|^{2}_{g})+\psi(u),$
where $\varphi$ and $\psi$ are given by
$\begin{split}\varphi\left(s\right)&=-\frac{1}{2}\log\left({1-\frac{s}{{2K}}}\right),\quad 0\leq s\leq K-1,\\ \psi\left(s\right)&=-A\log\left({1+\frac{s}{{2L}}}\right),\quad-L+1\leq s\leq 0,\end{split}$
with $K=\mathop{\sup}\limits_{M}|\nabla u|_{g}^{2}+1,\ L=\mathop{\sup}\limits_{M}\left|u\right|+1,\ A=2L(C_{0}+1)$ where $C_{0}$ is a uniform positive constant to be determined later. We have
$(4K)^{-1}$ $\leq\varphi^{\prime}\leq(2K)^{-1},\ \varphi^{\prime\prime}=2\big(\varphi^{\prime})^{2},$ (22)
$\frac{1}{2}{A}{L^{-1}}$ $\leq-\psi^{\prime}\leq{A}{L^{-1}},\ \psi^{\prime\prime}=A^{-1}(\psi^{\prime})^{2}.$ (23)
Suppose the function $H$ attains its maximum at $z_{0}$ along the direction $\xi_{0}$. We choose holomorphic coordinates near $z_{0}$ such that
$g_{i\bar{j}}(z_{0})=\delta_{i\bar{j}},\ \partial_{k}g_{i\bar{j}}(z_{0})=0,\text{and}$ (24)
$w_{i\bar{j}}(z_{0})=\lambda_{i}\delta_{i\bar{j}}\ \textup{with}\ \lambda_{1}\geq\lambda_{2}\geq\ldots\geq\lambda_{n}$
which forces $\xi_{0}=\frac{\partial}{\partial z_{1}}$. We extend $\xi_{0}$ near $z_{0}$ as $\tilde{\xi}_{0}(z)=(g_{1\bar{1}})^{-\frac{1}{2}}\frac{\partial}{\partial z_{1}}$. Then the function $Q(z)=H(z,\tilde{\xi}_{0}(z))$ attains its local maximum at $z_{0}$. By the maximum principle, we have at $z_{0}$
${Q_{i}}=\frac{{{w_{1\bar{1},i}}}}{{{w_{1\bar{1}}}}}+{\varphi_{i}}+{\psi_{i}}=0,$ (25)
and
$0\geq\mathcal{L}(Q)=F^{i\bar{i}}Q_{i\bar{i}}=$ $\lambda_{1}^{-1}F^{i\bar{i}}w_{1\bar{1},i\bar{i}}-\lambda_{1}^{-2}F^{i\bar{i}}{|w_{1\bar{1},i}|^{2}}$
$+\varphi^{\prime\prime}F^{i\bar{i}}\big|(|\nabla u|^{2})_{i}\big|^{2}+\varphi^{\prime}F^{i\bar{i}}\big(|\nabla u|^{2}\big)_{i\bar{i}}$
$+\psi^{\prime\prime}F^{i\bar{i}}|u_{i}|^{2}+\psi^{\prime}F^{i\bar{i}}u_{i\bar{i}}.$ (26)
We will estimate each term in (3.3). In the following, we always assume
$\lambda_{1}>80A\sqrt{A}K,$ (27)
otherwise we finish the proof.
Since $d\chi=0$, by covariant derivative formulae, we have
$w_{1\bar{1},i\bar{i}}=w_{i\bar{i},1\bar{1}}+(\lambda_{1}-\lambda_{i})R_{1\bar{1}i\bar{i}}.$ (28)
By (17) and $\lambda_{i}+\lambda_{j}\geq 0$ for $i\neq j$ by Lemma 2.2, we get
$F^{i\bar{i}}w_{i\bar{i},1\bar{1}}=$ $\sum_{i,j}\frac{\lambda_{i}+\lambda_{j}}{(1+\lambda_{i}^{2})(1+\lambda_{j}^{2})}|w_{i\bar{j},1}|^{2}$
$\geq$ $\sum_{i=2}^{n}\frac{\lambda_{1}+\lambda_{i}}{(1+\lambda_{1}^{2})(1+\lambda_{i}^{2})}|w_{1\bar{1},i}|^{2}+\sum_{i=1}^{n}\frac{2\lambda_{i}}{(1+\lambda_{i}^{2})^{2}}|w_{i\bar{i},1}|^{2},$ (29)
where we used $w_{i\bar{1},1}=w_{1\bar{1},i}$ since $d\chi=0$.
We prove
$\sum_{i=1}^{n}\frac{\lambda_{i}}{(1+\lambda_{i}^{2})^{2}}|w_{i\bar{i},1}|^{2}\geq 0.$ (30)
This is true if $\lambda_{n}\geq 0$. If $\lambda_{n}<0$, by (16), we have
$\big|\frac{w_{n\bar{n},1}}{1+\lambda_{n}^{2}}\big|=\big|\sum_{i=1}^{n-1}\frac{w_{i\bar{i},1}}{1+\lambda_{i}^{2}}\big|$
By the above equality and noting that $\lambda_{i}>0$ for $1\leq i\leq n-1$, we get
$\sum_{i=1}^{n}\frac{\lambda_{i}}{(1+\lambda_{i}^{2})^{2}}|w_{i\bar{i},1}|^{2}=$ $\sum_{i=1}^{n-1}\frac{\lambda_{i}|w_{i\bar{i},1}|^{2}}{(1+\lambda_{i}^{2})^{2}}+\lambda_{n}\big|\sum_{i=1}^{n-1}\frac{w_{i\bar{i},1}}{1+\lambda_{i}^{2}}\big|^{2}$
$\geq$ $\big(\sum_{i=1}^{n-1}\lambda_{i}^{-1}\big)^{-1}\big|\sum_{i=1}^{n-1}{\frac{w_{i\bar{i},1}}{1+\lambda_{i}^{2}}}\big|^{2}+\lambda_{n}\big|\sum_{i=1}^{n-1}{\frac{w_{i\bar{i},1}}{1+\lambda_{i}^{2}}}\big|^{2}$
$=$ $\frac{1+\lambda_{n}\sum\limits_{i=1}^{n-1}\lambda_{i}^{-1}}{\sum\limits_{i=1}^{n-1}\lambda_{i}^{-1}}\big|\sum_{i=1}^{n-1}{\frac{w_{i\bar{i},1}}{1+\lambda_{i}^{2}}}\big|^{2}$
$\geq$ $0,$
where in the last inequality we used
$1+\lambda_{n}\sum_{i=1}^{n-1}\lambda_{i}^{-1}=\lambda_{n}\frac{\sigma_{n-1}(\lambda)}{\sigma_{n}(\lambda)}\geq 0.$
Combining (3.3) with (30), we have
$F^{i\bar{i}}w_{i\bar{i},1\bar{1}}\geq\sum_{i=2}^{n}\frac{\lambda_{1}+\lambda_{i}}{(1+\lambda_{1}^{2})(1+\lambda_{i}^{2})}|w_{1\bar{1},i}|^{2}.$ (31)
Then we get
$\lambda_{1}^{-1}F^{i\bar{i}}w_{i\bar{i},1\bar{1}}-\lambda_{1}^{-2}F^{i\bar{i}}{|w_{1\bar{1},i}|^{2}}\geq$ $-\lambda_{1}^{-2}{(1+\lambda_{1}^{2})^{-1}}\big|w_{1\bar{1},1}|^{2}$
$-\sum_{i=2}^{n}\frac{|w_{1\bar{1},i}|^{2}}{\lambda_{1}^{2}(1+\lambda_{1}^{2})(1+\lambda_{i}^{2})}$
$+\sum_{i=2}^{n}\frac{\lambda_{i}}{\lambda_{1}(1+\lambda_{1}^{2})(1+\lambda_{i}^{2})}|w_{1\bar{1},i}|^{2}$
$=$ $\mathcal{I}+\mathcal{II}+\mathcal{III}.$ (32)
We first estimate $\mathcal{I}$ and $\mathcal{II}$ by (25):
$\mathcal{I}\geq$ $-2(\varphi^{\prime})^{2}F^{1\bar{1}}|(|\nabla u|^{2})_{1}|^{2}-2(\psi^{\prime})^{2}F^{1\bar{1}}|u_{1}|^{2}$
$\geq$ $-\varphi^{\prime\prime}F^{1\bar{1}}|(|\nabla u|^{2})_{1}|^{2}-2(\psi^{\prime})^{2}KF^{1\bar{1}}$ (33)
and
$\mathcal{II}=$ $-(1+\lambda_{1}^{2})^{-1}\sum_{i=2}^{n}\frac{F^{i\bar{i}}|w_{1\bar{1},i}|^{2}}{\lambda_{1}^{2}}$
$\geq$ $-2(1+\lambda_{1}^{2})^{-1}(\varphi^{\prime})^{2}\sum_{i=2}^{n}F^{i\bar{i}}|(|\nabla u|^{2})_{i}|^{2}-2(1+\lambda_{1}^{2})^{-1}(\psi^{\prime})^{2}\sum_{i=2}^{n}F^{i\bar{i}}|u_{i}|^{2}$
$\geq$ $-\frac{1}{12}\varphi^{\prime\prime}\sum_{i=2}^{n}F^{i\bar{i}}|(|\nabla u|^{2})_{i}|^{2}-\frac{1}{2}\psi^{\prime\prime}\sum_{i=2}^{n}F^{i\bar{i}}|u_{i}|^{2},$ (34)
where in the last inequality we used (27).
Next we estimate term $\mathcal{III}$ . Since $\lambda_{i}>0$ for $1\leq i\leq n-1$ by Lemma 2.2, we get
$\mathcal{III}\geq$ $\frac{\lambda_{n}|w_{1\bar{1},n}|^{2}}{\lambda_{1}(1+\lambda_{1}^{2})(1+\lambda_{n}^{2})},$
then by (25) and the Cauchy-Schwarz inequality, we obtain
$\mathcal{III}\geq$ $-\frac{4}{3}\frac{\lambda_{1}|\lambda_{n}|}{1+\lambda_{1}^{2}}(\varphi^{\prime})^{2}F^{n\bar{n}}|(|\nabla u|^{2})_{n}|^{2}-\frac{4\lambda_{1}|\lambda_{n}|}{1+\lambda_{1}^{2}}(\psi^{\prime})^{2}F^{n\bar{n}}|u_{n}|^{2}$
$\geq$ $-\frac{2}{3}\varphi^{\prime\prime}F^{n\bar{n}}|(|\nabla u|^{2})_{n}|^{2}-\frac{4\lambda_{1}|\lambda_{n}|}{1+\lambda_{1}^{2}}(\psi^{\prime})^{2}F^{n\bar{n}}|u_{n}|^{2},$ (35)
where we used (22). Let $\tau=\frac{1}{8A}$. We continue to estimate $\mathcal{III}$ by two cases.
Case 1: $|\lambda_{n}|\leq\tau\lambda_{1}$. In this case, we have
$\mathcal{III}\geq$ $-\frac{2}{3}\varphi^{\prime\prime}F^{n\bar{n}}|(|\nabla u|^{2})_{n}|^{2}-4\tau(\psi^{\prime})^{2}F^{n\bar{n}}|u_{n}|^{2}$
$\geq$ $-\frac{2}{3}\varphi^{\prime\prime}F^{n\bar{n}}|(|\nabla u|^{2})_{n}|^{2}-\frac{1}{2}\psi^{\prime\prime}F^{n\bar{n}}|u_{n}|^{2},$ (36)
where we used (23).
Case 2: $|\lambda_{n}|\geq\tau\lambda_{1}$. By (27), we have $|u_{n\bar{n}}|^{2}=(\lambda_{n}-\chi_{n\bar{n}}-t)^{2}\geq\frac{\tau^{2}}{2}\lambda_{1}^{2}$ and we obtain
$4(\psi^{\prime})^{2}\frac{\lambda_{1}|\lambda_{n}|}{1+\lambda_{1}^{2}}F^{n\bar{n}}|u_{n}|^{2}\leq$ $4AF^{n\bar{n}}K\leq\frac{1}{3}{\varphi^{\prime}}F^{n\bar{n}}u_{n\bar{n}}^{2}.$
Inserting the above inequality into (3.3), we have
$\mathcal{III}\geq-\frac{2}{3}\varphi^{\prime\prime}F^{n\bar{n}}|(|\nabla u|^{2})_{n}|^{2}-\frac{1}{3}\varphi^{\prime}F^{n\bar{n}}u_{n\bar{n}}^{2}.$ (37)
By (3.3) and (37), we obtain
$\mathcal{III}\geq-\frac{2}{3}\varphi^{\prime\prime}F^{n\bar{n}}|(|\nabla u|^{2})_{n}|^{2}-\frac{1}{2}\psi^{\prime\prime}F^{n\bar{n}}|u_{n}|^{2}-\frac{1}{3}{\varphi^{\prime}}F^{n\bar{n}}u_{n\bar{n}}^{2}.$ (38)
Inserting (3.3), (3.3) and (38) into (32), we obtain
$\lambda_{1}^{-1}F^{i\bar{i}}w_{i\bar{i},1\bar{1}}-\lambda_{1}^{-2}F^{i\bar{i}}{|w_{1\bar{1},i}|^{2}}+\frac{3}{4}\varphi^{\prime\prime}F^{i\bar{i}}|(|\nabla u|^{2})_{i}|^{2}+\psi^{\prime\prime}F^{i\bar{i}}|u_{i}|^{2}+\frac{1}{3}\varphi^{\prime}F^{i\bar{i}}u_{i\bar{i}}^{2}$
$\geq$ $\frac{1}{3}\varphi^{\prime}F^{1\bar{1}}u_{1\bar{1}}^{2}-2K(\psi^{\prime})^{2}F^{1\bar{1}}>0,$ (39)
where we used $u_{1\bar{1}}\geq\frac{\lambda_{1}}{2}$ which follows from (27).
Next we estimate $F^{i\bar{i}}(|\nabla u|^{2})_{i\bar{i}}$. By direct computation, we have
$F^{i\bar{i}}(|\nabla u|^{2})_{i\bar{i}}=F^{i\bar{i}}(u_{k\bar{i}}u_{\bar{k}i}+u_{ki}u_{\bar{k}\bar{i}})+F^{i\bar{i}}(u_{ki\bar{i}}u_{\bar{k}}+u_{k}u_{\bar{k}i\bar{i}}).$ (40)
By covariant derivative formulae and (16), we have
$F^{i\bar{i}}(u_{ki\bar{i}}u_{\bar{k}}+u_{k}u_{\bar{k}i\bar{i}})$
$\geq$ $-F^{i\bar{i}}(u_{i\bar{i}k}u_{\bar{k}}+u_{k}u_{i\bar{i}\bar{k}})-|Rm||\nabla u|^{2}\sum_{i}F^{i\bar{i}}$
$\geq$ $-2K(|\nabla\chi|+|Rm|)\sum_{i}F^{i\bar{i}}.$ (41)
Inserting the above estimate into (40) , we get
$F^{i\bar{i}}(|\nabla u|^{2})_{i\bar{i}}\geq F^{i\bar{i}}u_{i\bar{i}}^{2}-2K(|\nabla\chi|+|Rm|)\sum_{i=1}^{n}F^{i\bar{i}}.$ (42)
Inserting (28), (3.3) and (42) into (3.3), we obtain
$0\geq F^{i\bar{i}}Q_{i\bar{i}}\geq$ $\psi^{\prime}\sum_{i=1}^{n}F^{i\bar{i}}u_{i\bar{i}}+\frac{1}{3}\varphi^{\prime}F^{i\bar{i}}u_{i\bar{i}}^{2}-3K\varphi^{\prime}(|\nabla\chi|+|Rm|)\sum_{i}F^{i\bar{i}}$
$\geq$ $\psi^{\prime}\sum_{i=1}^{n}F^{i\bar{i}}u_{i\bar{i}}+\frac{1}{3}\varphi^{\prime}F^{i\bar{i}}u_{i\bar{i}}^{2}-2(|\nabla\chi|+|Rm|)\sum_{i}F^{i\bar{i}},$ (43)
By Lemma 3.3 and noting that we have assumed $\underline{u}=0$, one of the following holds
$(i)\ -\sum_{i=1}^{n}F^{i\bar{i}}u_{i\bar{i}}\geq\delta_{0}\sum_{i=1}^{n}F^{i\bar{i}},\ \text{or}$
$(ii)\ F^{1\bar{1}}\geq\delta_{0}\sum_{i=1}^{n}F^{i\bar{i}}.$
We choose $C_{0}=6\delta_{0}^{-1}(\max|\nabla\chi|+|Rm|)$. If $(i)$ holds, by (3.3) and $\psi^{\prime}\geq C_{0}+1$, we get a contradiction. Then $(ii)$ holds and we obtain $|\lambda_{n}|\geq\frac{\delta_{0}}{2}\lambda_{1}$. Hence we have
$0\geq$ $\frac{\lambda_{1}^{2}}{4}F^{n\bar{n}}-3K(|\nabla\chi|+|Rm|)\sum_{i}F^{i\bar{i}}$
$\geq$ $\big(\frac{\lambda_{1}^{2}}{4}-3nK(|\nabla\chi|+|Rm|)\big)F^{n\bar{n}}.$
Therefore we get $\lambda_{1}\leq CK$ and $C$ is a uniform positive constant which is independent of $t$. ∎
4. Gradient estimates and the proof of the main theorem
In this section, following the argument by Dinew-Kolodziej [10] and Szekelyhidi [36], we prove a Liouville theorem for the equation $\sigma_{n}+\sigma_{n-1}=0$. Then based on the uniform complex Hessian estimates of Hou-Ma-Wu type in Section 3, we get the gradient estimate by a blow-up argument. At last, we prove the main theorems.
4.1. Viscosity solutions
We begin by defining viscosity solutions for the equation
$\sigma_{n-1}(\sqrt{-1}\partial\bar{\partial}v)+\sigma_{n}(\sqrt{-1}\partial\bar{\partial}v)=0.$ (44)
A $C^{2}$ function $v:\Omega\rightarrow\mathbb{R}$ is said to be (n1)(n-1)-admissible in a domain $\Omega\subset\mathbb{C}^{n}$ if, for all $z\in\Omega$, the eigenvalue vector $\lambda(\sqrt{-1}\partial\bar{\partial}v(z))$ belongs to $\overline{\Gamma}_{n-1}$.
Definition 4.1.
$(i)$ A continuous function $v:\Omega\rightarrow\mathbb{R}$ is a viscosity subsolution of (44) if for every $C^{2}$ function $\varphi$ and every local minimum point $z_{0}\in\Omega$ of $\varphi-v$, the following inequality holds:
$\sigma_{n-1}(\sqrt{-1}\partial\bar{\partial}\varphi(z_{0}))+\sigma_{n}(\sqrt{-1}\partial\bar{\partial}\varphi(z_{0}))\geq 0.$
$(ii)$ A continuous function $v:\Omega\rightarrow\mathbb{R}$ is a viscosity supersolution of (44) if for every (n1)(n-1)-admissible function $\varphi$ and every local maximum point $z_{0}\in\Omega$ of $\varphi-v$, the following inequality holds:
$\sigma_{n-1}(\sqrt{-1}\partial\bar{\partial}\varphi(z_{0}))+\sigma_{n}(\sqrt{-1}\partial\bar{\partial}\varphi(z_{0}))\leq 0.$
$(iii)$ A continuous function $v:\Omega\rightarrow\mathbb{R}$ is a viscosity solution of (44) if it is both a viscosity subsolution and a viscosity supersolution.
If $v$ is a $C^{2}$ viscosity subsolution of (44), then it satisfies $\sigma_{n-1}(\sqrt{-1}\partial\bar{\partial}v)+\sigma_{n}(\sqrt{-1}\partial\bar{\partial}v)\geq 0$ in the classical sense.
We now establish some basic properties.
Lemma 4.1.
$(i)$ Let $\{v^{i}\}$ be a sequence of smooth viscosity subsolutions of (44) in $\Omega$ that converges locally uniformly to $v$. Then $v$ is also a viscosity subsolution of (44) in $\Omega$.
$(ii)$ Let $v$ be a viscosity subsolution of (44) in $\mathbb{C}^{n}$. If $v$ is independent of $z_{n}$, then its restriction to $\mathbb{C}^{n-1}$ is a viscosity subsolution of $\sigma_{n-1}=0$.
Proof.
For any $C^{2}$ function $\varphi$ and any $z_{0}\in\Omega^{n}$ of the local minimum point of $\varphi-v$, we need to show
$\sigma_{n-1}(\varphi_{\alpha\bar{\beta}}(z_{0}))+a_{0}\sigma_{n}(\varphi_{\alpha\bar{\beta}}(z_{0}))\geq 0.$ (45)
For any $\epsilon>0$, for all sufficiently large $i$, we have $\sup_{B_{1}(z_{0})}|v^{i}-v|<\frac{\epsilon^{3}}{3}$. Then the function $\Phi=\varphi+\epsilon|z-z_{0}|^{2}-v^{i}$ satisfies $\Phi|_{\partial B_{\epsilon}(z_{0})}>\Phi(z_{0})$, and thus $\Phi$ attains its minimum on $\overline{B_{\epsilon}(z_{0})}$ at some point $z_{i,\epsilon}\in B_{\epsilon}(z_{0})$. Since $v^{i}$ is a viscosity subsolution, we have
$\sigma_{n-1}(\varphi_{\alpha\bar{\beta}}(z_{i,\epsilon})+\epsilon\delta_{\alpha\beta})+a_{0}\sigma_{n}(\varphi_{\alpha\bar{\beta}}(z_{i,\epsilon})+\epsilon\delta_{\alpha\beta})\geq 0.$
Letting $\epsilon\to 0$ then implies (45). This proves $(i)$.
$(ii)$ follows from the definition of the viscosity subsolution.
We will need the following comparison principle in the proof of the Liouville Theorem 4.1.
Lemma 4.2.
Let $\Omega\subset\mathbb{C}^{n}$ be a bounded domain. Suppose $w$ is an (n1)(n-1)-admissible function and a viscosity subsolution of (44) in $\Omega$, and $v$ is a viscosity solution of (44) in $\Omega$. If $w\leq v$ on $\partial\Omega$, then $w\leq v$ in $\Omega$.
Proof.
Assume that there exists a point $x_{0}\in\Omega$ such that $w(x_{0})>v(x_{0})$. Choose $\epsilon>0$ sufficiently small so that
$\epsilon<\frac{w(x_{0})-v(x_{0})}{\max_{z\in\overline{\Omega}}|z|^{2}}.$
Since $w\leq v$ on $\partial\Omega$, we deduce the following chain of inequalities:
$\max_{z\in\overline{\Omega}}\,(w+\epsilon|z|^{2}-v)$ $\geq w(x_{0})-v(x_{0})$
$>\epsilon\cdot\max_{z\in\partial\Omega}|z|^{2}$
$\geq\max_{z\in\partial\Omega}\,(w+\epsilon|z|^{2}-v).$
This implies that the function $w+\epsilon|z|^{2}-v$ attains its maximum at an interior point $z_{0}\in\Omega$.
Now, since $w$ is (n1)(n-1)-admissible, the function $w+\epsilon|z|^{2}$ is also (n1)(n-1)-admissible. As $v$ is a viscosity solution (and hence a supersolution), and $w+\epsilon|z|^{2}$ touches $v$ from above at $z_{0}$, the supersolution property yields:
$\sigma_{n-1}((w_{\alpha\bar{\beta}}(z_{0}))+\epsilon I_{n})+a_{0}\sigma_{n}((w_{\alpha\bar{\beta}}(z_{0}))+\epsilon I_{n})\leq 0.$ (46)
However, since $\sigma_{k}(w_{\alpha\bar{\beta}}(z_{0}))\geq 0$ for $1\leq k\leq n-1$, expanding the left-hand side of (46) gives:
$\sigma_{n-1}(w_{\alpha\bar{\beta}}(z_{0}))+a_{0}\sigma_{n}(w_{\alpha\bar{\beta}}(z_{0}))\leq-\epsilon^{n-1}<0.$ (47)
This contradicts the assumption that $w$ is a viscosity subsolution of (44), thereby completing the proof. ∎
4.2. Weak solutions in the sense of currents
Let $\mathcal{E}=\sum_{\alpha}\sqrt{-1}dz_{\alpha}\wedge d\bar{z}_{\alpha}$ denote the standard Kähler metric on $\mathbb{C}^{n}$. We recall the definition of $k$-subharmonic functions following Blocki [1].
Definition 4.2.
For $1\leq k\leq n$, a continuous function $v:\Omega\rightarrow\mathbb{R}$ is called $k$-subharmonic if $v$ is subharmonic and
$\sqrt{-1}\partial\bar{\partial}v\wedge\gamma_{1}\wedge\cdots\wedge\gamma_{k-1}\wedge\mathcal{E}^{n-k}\geq 0\quad\text{in }\Omega$
for all real $(1,1)$-forms $\gamma_{1},\dots,\gamma_{k-1}$ with $\lambda(\gamma_{i})\in\overline{\Gamma}_{k}$ for $1\leq i\leq k-1$.
If $v$ is $C^{2}$ and $k$-subharmonic, then $v$ is $k$-admissible. More generally, by Lemma 3.7 in [29], $v$ is a viscosity subsolution of $\sigma_{k}=0$ if and only if $v$ is $k$-subharmonic.
The following lemma, which generalizes the Garding inequality, is particularly useful for $k$-subharmonic functions (see [1, page 1444, (3.4)]).
Lemma 4.3.
Let $v,v_{1},\dots,v_{k-1}$ be continuous $k$-subharmonic functions. Then
$(\sqrt{-1}\partial\bar{\partial}v)\wedge(\sqrt{-1}\partial\bar{\partial}v_{1})\wedge\cdots\wedge(\sqrt{-1}\partial\bar{\partial}v_{k-1})\wedge\mathcal{E}^{n-k}\geq 0$
in the sense of currents. In particular,
$\sqrt{-1}\partial v\wedge\bar{\partial}v\wedge(\sqrt{-1}\partial\bar{\partial}v)^{k-1}\wedge\mathcal{E}^{n-k}\geq 0$
in the sense of currents.
We now define weak solutions in the sense of currents.
Definition 4.3.
Let $v$ be a continuous (n1)(n-1)-subharmonic function in $\mathbb{C}^{n}$ such that $v+|z|^{2}$ is plurisubharmonic. We say that $v$ is a weak solution (respectively, weak subsolution) of the equation $\sigma_{n}+\sigma_{n-1}=0$ if it satisfies
[…]
in the sense of currents.
Remark 4.1.
The condition that $v+|z|^{2}$ is plurisubharmonic ensures that $(\sqrt{-1}\partial\bar{\partial}v)^{n}$ is well-defined as a current. Note that the blow-up solution constructed in Subsection 4.4 satisfies this condition. On the other hand, if $v$ is of class $C^{2}$, then $\sigma_{n}(\sqrt{-1}\partial\bar{\partial}v)+\sigma_{n-1}(\sqrt{-1}\partial\bar{\partial}v)=0$ (respectively, $\geq 0$) is equivalent to $(\sqrt{-1}\partial\bar{\partial}v)^{n}+n(\sqrt{-1}\partial\bar{\partial}v)^{n-1}\wedge\mathcal{E}=0$ (respectively, $\geq 0$).
Let $\eta\in C^{\infty}(\mathbb{C}^{n},\mathbb{R})$ be a mollifier, defined by
$\eta(z)=\begin{cases}C_{n}e^{\frac{1}{|z|^{2}-1}},&|z|<1,\\ 0,&|z|\geq 1,\end{cases}$ (48)
where the constant $C_{n}$ is chosen so that $\int_{\mathbb{C}^{n}}\eta(z)\,\mathcal{E}^{n}(z)=1$. Similarly, let $\eta_{0}$ be a mollifier on $\mathbb{C}$.
For a function $w:\mathbb{C}^{n}\rightarrow\mathbb{R}$, we define its mollification by
$[w]_{r}(z)=\int_{\mathbb{C}^{n}}w(z+ry)\eta(y)\,\mathcal{E}^{n}(y).$ (49)
For the function $\tilde{w}(\tilde{z})=w(z)+|z_{n+1}|^{2}$ defined on $\mathbb{C}^{n+1}$, where $\tilde{z}=(z,z_{n+1})$, we define its mollification as
[…]
where $\tilde{\eta}(\tilde{y})=\eta_{0}(y_{n+1})\eta(y)$ with $\tilde{y}=(y,y_{n+1})$. A direct computation shows that
$\langle\tilde{w}\rangle_{r}(\tilde{z})=[w]_{r}(z)+|z_{n+1}|^{2}+r^{2}M,$
where $M:=\int_{\mathbb{C}}|y_{n+1}|^{2}\eta_{0}(y_{n+1})\,\sqrt{-1}dy_{n+1}\wedge d\bar{y}_{n+1}$.
The following lemma establishes key properties of weak solutions to the equation $\sigma_{n}+\sigma_{n-1}=0$.
Lemma 4.4.
Let $v$ be a weak solution (in the sense of currents) to $\sigma_{n}+\sigma_{n-1}=0$. Then the following properties hold:
  1. (i)
    The functions $\tilde{v}(\tilde{z})=v(z)+|z_{n+1}|^{2}$ is a weak solution to $\sigma_{n}=0$ and $\langle\tilde{v}\rangle_{r}$ is $n$-subharmonic in $\mathbb{C}^{n+1}$.
  2. (ii)
    The mollification $v]_{r}$ is a smooth subsolution to the same equation.
  3. (iii)
    The function $\frac{1}{2}v^{2}$ is a subsolution to the same equation.
  4. (iv)
    If $w$ is another subsolution, then $\frac{1}{2}(v+w)$ is a weak subsolution.
Proof.
We first note that $v]_{r}$ is (n1)(n-1)-subharmonic since $v$ is (n1)(n-1)-subharmonic. To prove $(i)$, we will show that $\tilde{v}$ satisfies the equation: $\sigma_{n}=0$ on $\mathbb{C}^{n+1}$ in the current sense.
Let $e_{n+1}=\sqrt{-1}dy_{n+1}\wedge d\bar{y}_{n+1}$ and let $\widetilde{\mathcal{E}}=e_{n+1}+\mathcal{E}$ be the standard Kähler metric on $\mathbb{C}^{n+1}$. Since $v+|z|^{2}$ is plurisubharmonic in $\mathbb{C}^{n}$, it follows that $\tilde{v}+|z|^{2}$ is also plurisubharmonic in $\mathbb{C}^{n+1}$. Moreover, for any $1\leq k\leq n$, we have:
$(\sqrt{-1}\partial\bar{\partial}\tilde{v})^{k}\wedge\widetilde{\mathcal{E}}^{n+1-k}$
$=$ $C^{k}_{n+1}(\sqrt{-1}\partial\bar{\partial}v+e_{n+1})^{k}\wedge(\mathcal{E}+e_{n+1})^{n+1-k}$
$=$ $\left((\sqrt{-1}\partial\bar{\partial}v)^{k}+k(\sqrt{-1}\partial\bar{\partial}v)^{k-1}\wedge e_{n+1}\right)\wedge\left(\mathcal{E}^{n-k+1}+(n+1-k)\mathcal{E}^{n-k}\wedge e_{n+1}\right)$
$=$ $\left((n-k+1)(\sqrt{-1}\partial\bar{\partial}v)^{k}\wedge\mathcal{E}^{n-k}+k(\sqrt{-1}\partial\bar{\partial}v)^{k-1}\wedge\mathcal{E}^{n-k+1}\right)\wedge e_{n+1}.$
[…]
$(\sqrt{-1}\partial\bar{\partial}v)^{n}+n(\sqrt{-1}\partial\bar{\partial}v)^{n-1}\wedge\mathcal{E}=0.$
This proves that $\tilde{v}$ is $n$-subharmonic in $\mathbb{C}^{n+1}$ and satisfies $\sigma_{n}=0$ in the current sense. Consequently, $\langle\tilde{v}\rangle_{r}$ is also $n$-subharmonic in $\mathbb{C}^{n+1}$.
Since $\langle\tilde{v}\rangle_{r}=[v]_{r}(z)+|z_{n+1}|^{2}+r^{2}M$, it follows that $v]_{r}$ is a subsolution to the equation $\sigma_{n}+\sigma_{n-1}=0$. This establishes (ii).
To prove (iv), let $W=\frac{v+w}{2}$. Then $W+|z|^{2}$ is plurisubharmonic and (n1)(n-1)-subharmonic. Define $\widetilde{W}(\tilde{z})=W(z)+|z_{n+1}|^{2}$. By $(i)$, we have $\widetilde{W}=\frac{\tilde{v}+\tilde{w}}{2}$, which is $n$-subharmonic in $\mathbb{C}^{n+1}$. Therefore, $W$ is a subsolution, proving $(iv)$.
We now prove (iii) using (i). Define $\tilde{F}(\tilde{z})=\frac{1}{2}v(z)^{2}+|z_{n+1}|^{2}$. Then for any $1\leq k\leq n$, we have:
$(\sqrt{-1}\partial\bar{\partial}\tilde{F})^{k}\wedge\widetilde{\mathcal{E}}^{n+1-k}$
$=$ $\left(v\sqrt{-1}\partial\bar{\partial}\tilde{v}+(1-v)e_{n+1}+\sqrt{-1}\partial v\wedge\bar{\partial}v\right)^{k}\wedge\widetilde{\mathcal{E}}^{n+1-k}$
$=$ $v^{k}(\sqrt{-1}\partial\bar{\partial}\tilde{v})^{k}\wedge\widetilde{\mathcal{E}}^{n+1-k}$
$+kv^{k-1}(\sqrt{-1}\partial\bar{\partial}\tilde{v})^{k-1}\wedge\left((1-v)e_{n+1}+\sqrt{-1}\partial v\wedge\bar{\partial}v\right)\wedge\widetilde{\mathcal{E}}^{n+1-k}$
$+k(k-1)(1-v)v^{k-2}(\sqrt{-1}\partial\bar{\partial}\tilde{v})^{k-2}\wedge e_{n+1}\wedge(\sqrt{-1}\partial v\wedge\bar{\partial}v)\wedge\widetilde{\mathcal{E}}^{n+1-k}.$
This expression is nonnegative by Lemma 4.3, since $\tilde{v}$ is $n$-subharmonic in $\mathbb{C}^{n+1}$ and $0\leq v\leq 1$. ∎
We immediately conclude that a weak solution (subsolution) in the sense of currents is also a viscosity solution (subsolution).
Lemma 4.5.
Let $v$ be a weak solution (respectively, subsolution) in the sense of currents to the equation $\sigma_{n}+\sigma_{n-1}=0$. Then $v$ is also a viscosity solution (respectively, subsolution).
Proof.
By part (i) of Lemma 4.4, we know that $\tilde{v}=v+|z_{n+1}|^{2}$ is $n$-subharmonic in $\mathbb{C}^{n+1}$, and hence $\tilde{v}$ is a viscosity subsolution to $\sigma_{n}=0$ in $\mathbb{C}^{n+1}$.
Now, let $\varphi$ be any $C^{2}$ function and let $z_{0}$ be a local minimum point of $\varphi-v$. Consider the function $\tilde{\varphi}=\varphi+|z_{n+1}|^{2}$. Then $(z_{0},0)$ is a local minimum point of $\tilde{\varphi}-\tilde{v}$. Since $\tilde{v}$ is a viscosity subsolution to $\sigma_{n}=0$, we have
$\sigma_{n}(\sqrt{-1}\partial\bar{\partial}\tilde{\varphi})(0,z_{0})\geq 0.$
Observing that
$\sigma_{n}(\sqrt{-1}\partial\bar{\partial}\tilde{\varphi})(z_{0},0)=\sigma_{n}(\sqrt{-1}\partial\bar{\partial}\varphi)(z_{0})+\sigma_{n-1}(\sqrt{-1}\partial\bar{\partial}\varphi)(z_{0}),$
we conclude that $v$ is a viscosity subsolution.
The proof for supersolutions follows similarly, using the corresponding definition of viscosity supersolutions. ∎
4.3. A Liouville theorem to the equation $\sigma_{n}+\sigma_{n-1}=0$ in $\mathbb{C}^{n}$
In this subsection, based on the lemmas established in the preceding subsections and following arguments similar to those in [10, 36], we prove the following Liouville-type theorem.
Theorem 4.1.
Let $v:\mathbb{C}^{n}\rightarrow\mathbb{R}$ be a $C^{1}$ function that is (n1)(n-1)-subharmonic and such that $v+|z|^{2}$ is plurisubharmonic. Assume there exists a positive constant $C$ such that $\|v\|_{C^{1}(\mathbb{C}^{n})}\leq C$ and $\Delta v\leq C$ in the weak sense (see Remark 4.2 for details). If $v$ is a weak solution in the sense of currents to the equation
$\sigma_{n-1}(\sqrt{-1}\partial\bar{\partial}v)+\sigma_{n}(\sqrt{-1}\partial\bar{\partial}v)=0,$ (50)
then $v$ must be constant.
Remark 4.2.
The condition $\Delta v\leq C$ in the weak sense means that there exists a constant $C>0$ such that for every nonnegative test function $f\in C_{c}^{\infty}(\mathbb{C}^{n})$ with $\int_{\mathbb{C}^{n}}f(z)\,\mathcal{E}^{n}(z)=1$, we have
$\int_{\mathbb{C}^{n}}v\,\sqrt{-1}\partial\bar{\partial}f\wedge\mathcal{E}^{n-1}\leq C.$
From this condition, it follows immediately that the regularization $v]_{r}$ has bounded Laplacian and hence bounded complex Hessian.
Proof of Theorem 4.1.
Assume that $v$ is not constant. Then $\inf_{\mathbb{C}^{n}}v<\sup_{\mathbb{C}^{n}}v$. By scaling, we may assume that $v$ satisfies the following normalized conditions:
$\begin{cases}\sup_{\mathbb{C}^{n}}v=1,\quad\inf_{\mathbb{C}^{n}}v=0,\\ |\nabla v|\leq C_{1}:=(\sup_{\mathbb{C}^{n}}v-\inf_{\mathbb{C}^{n}}v)^{\frac{1}{2}},\quad\Delta v\leq C,\\ \sigma_{n-1}(\sqrt{-1}\partial\bar{\partial}v)+\sigma_{n}(\sqrt{-1}\partial\bar{\partial}v)=0.\end{cases}$ (51)
Indeed, the function $\tilde{v}(z)=C_{1}^{-2}({v(C_{1}z)-\inf_{\mathbb{C}^{n}}v})$ satisfies these conditions.
For sufficiently small $\epsilon>0$, let $v^{\epsilon}=[v]_{\epsilon}$. Consider the function
Case 1: There exist a positive constant $\rho$, sequences $\epsilon_{k}\to 0$, $x_{k}\in\mathbb{C}^{n}$, $r_{k}\to\infty$, and a sequence of unit $(1,0)$-vectors $\xi_{k}$ such that
$\frac{1}{2}\bigl(\bigl[\frac{(v^{\epsilon})^{2}}{3}\bigr]_{r}(z)+[v]_{\rho}(z)-\frac{7}{6}\bigr)$ $\geq v(x_{k}),$ (52)
$\lim_{k\to\infty}\int_{B_{r_{k}}(x_{k})}|v_{\xi_{k}}^{\epsilon}|^{2}(z)\,\mathcal{E}^{n}(z)$ $=0.$ (53)
Following the arguments in [36] and [10], we obtain an (n1)(n-1)-subharmonic function $v^{\infty}$ that is also a solution to the same equation and is independent of $z_{n}$. This implies that $v^{\infty}$ is plurisubharmonic in $\mathbb{C}^{n-1}$. Since $v^{\infty}$ is bounded, it must be constant, contradicting (52).
Case 2: Suppose Case 1 does not hold. Then for every $\rho>0$, there exists a constant $C_{\rho}>0$ with the following property: for all $\epsilon<C_{\rho}^{-1}$, $r>C_{\rho}$, $x\in\mathbb{C}^{n}$, and any unit $(1,0)$-vector $\xi$, if $\frac{1}{2}\bigl(\bigl[\frac{(v^{\epsilon})^{2}}{3}\bigr]_{r}(z)-{\tau_{0}}|z|^{2}+[v]_{\rho}(z)\bigr)-v(x)\geq 0$, then
$\int_{B_{r}(x)}|v^{\epsilon}_{\xi}|^{2}(z)\,\mathcal{E}^{n}(z)\geq c_{0}:=C_{\rho}^{-1}.$ (54)
Define the open set
$\Omega:=\bigl\{z\in\mathbb{C}^{n}:v(z)<\frac{1}{2}\bigl(\bigl[\frac{(v^{\epsilon})^{2}}{3}\bigr]_{r}(z)-{\tau_{0}}|z|^{2}+[v]_{\rho}(z)-\frac{7}{6}\bigr)\bigr\},$
where $\tau_{0}>0$ is a constant to be determined later.
Since $0\leq v\leq 1$, the set $\Omega$ is bounded. As $\inf_{\mathbb{C}^{n}}v=0$, we may assume (after translation) that $v(0)<\frac{1}{60}$. By Cartan’s lemma (see [10]), we choose $\rho$ large enough so that $v]_{\rho}(0)>\frac{9}{10}$, and choose $r$ large enough so that $\bigl[\frac{v^{2}}{3}\bigr]_{r}(0)>\frac{3}{10}$. Then $0\in\Omega$.
We now show that $\frac{1}{2}\bigl(\bigl[\frac{(v^{\epsilon})^{2}}{3}\bigr]_{r}(z)-{\tau_{0}}|z|^{2}+[v]_{\rho}(z)\bigr)$ is a subsolution to $\sigma_{n-1}+\sigma_{n}=0$. Assuming this, the comparison principle (Lemma 4.2) implies that $v\geq w$ in $\Omega$, contradicting the definition of $\Omega$.
By Lemma 4.4 $(ii)$, we know that $v]_{r}$, $v]_{\rho}$, $\frac{(v^{\epsilon})^{2}}{2}$, and $\bigl[\frac{(v^{\epsilon})^{2}}{2}\bigr]_{r}$ are all subsolutions to (50).
From (54) and following [10, 36], there exists a positive constant $b_{0}$, depending only on $c_{0}$ and $C_{n}$, such that $\bigl[\frac{(v^{\epsilon})^{2}}{2}\bigr]_{r}-b_{0}|z|^{2}$ is (n1)(n-1)-subharmonic. By the concavity of $\sigma_{n-1}^{1/(n-1)}$ in $\Gamma_{n-1}$, we have
$\sigma_{n-1}^{1/(n-1)}\bigl(\sqrt{-1}\partial\bar{\partial}\bigl[\frac{(v^{\epsilon})^{2}}{2}\bigr]_{r}\bigr)\geq(n-1)^{1/(n-1)}b_{0},$
and thus
$\sigma_{n-1}\bigl(\sqrt{-1}\partial\bar{\partial}\bigl[\frac{(v^{\epsilon})^{2}}{2}\bigr]_{r}\bigr)\geq(n-1)b_{0}^{n-1}.$ (55)
We now show that $F:=\bigl[\frac{(v^{\epsilon})^{2}}{3}\bigr]_{r}-\tau_{0}|z|^{2}$ is a subsolution to $\sigma_{n}+\sigma_{n-1}=0$ for sufficiently small $\tau_{0}$. Indeed, by (55), we have:
$\sigma_{n}(\sqrt{-1}\partial\bar{\partial}F)+\sigma_{n-1}(\sqrt{-1}\partial\bar{\partial}F)$
$=\sigma_{n}\bigl(\sqrt{-1}\partial\bar{\partial}\bigl[\frac{(v^{\epsilon})^{2}}{3}\bigr]_{r}\bigr)+\sigma_{n-1}\bigl(\sqrt{-1}\partial\bar{\partial}\bigl[\frac{(v^{\epsilon})^{2}}{3}\bigr]_{r}\bigr)+O(\tau_{0})$
$=\bigl(\frac{2}{3}\bigr)^{n}\bigl[\sigma_{n}\bigl(\sqrt{-1}\partial\bar{\partial}\bigl[\frac{(v^{\epsilon})^{2}}{3}\bigr]_{r}\bigr)+\frac{3}{2}\sigma_{n-1}\bigl(\sqrt{-1}\partial\bar{\partial}\bigl[\frac{(v^{\epsilon})^{2}}{3}\bigr]_{r}\bigr)\bigr]+O(\tau_{0})$
$\geq\bigl(\frac{2}{3}\bigr)^{n-1}\sigma_{n-1}\bigl(\sqrt{-1}\partial\bar{\partial}\bigl[\frac{v^{2}}{3}\bigr]_{r}\bigr)+O(\tau_{0})$
$\geq\bigl(\frac{2}{3}\bigr)^{n-1}(n-1)b_{0}^{n-1}-C\tau_{0}>0,$
provided $\tau_{0}<C^{-1}\left(\frac{2b_{0}}{3}\right)^{n-1}$. Here we have used the fact that $(v^{\epsilon})^{2}]_{r}$ has uniform $C^{1}$-bounds and complex Hessian bounds.
By Lemma 4.4 $(iv)$, we get $\frac{1}{2}\bigl(\bigl[\frac{(v^{\epsilon})^{2}}{3}\bigr]_{r}(z)-{\tau_{0}}|z|^{2}+[v]_{\rho}(z)\bigr)$ is a subsolution to the equation $\sigma_{n}+\sigma_{n-1}=0$. ∎
4.4. Gradient estimates
We prove the gradient estimate by contradiction, following the arguments in [10, 36]. Assume that there exist a sequence $t_{i}\to 0$ and corresponding solutions $u^{t_{i}}$ to the LYZ equation (6) such that for points $p_{i}\in M$, we have
$\sup_{M}|\nabla u^{t_{i}}|_{\omega}=|\nabla u^{t_{i}}|_{\omega}(p_{i})=M_{i}\to\infty\quad\text{as }i\to\infty.$ (56)
Since $M$ is compact, by passing to a subsequence, we may assume that $\{p_{i}\}$ converges to a point $p_{0}\in M$. We choose local coordinates $\{w_{\alpha}\}_{1\leq\alpha\leq n}$ in a ball $B_{r}(p_{0})$ such that $p_{i}\in B_{r/2}(p_{0})$ and $\omega(p_{0})=\sum_{\alpha}\sqrt{-1}dw_{\alpha}\wedge d\bar{w}_{\alpha}$.
For any $R>0$, we can choose $i$ sufficiently large such that for all $z\in B_{R}(0)$, the point $p_{i}+\frac{z}{M_{i}}$ lies in $B_{2r/3}(p_{0})$. We define the rescaled functions
$\hat{u}^{t_{i}}(z)=u^{t_{i}}\bigl(p_{i}+\frac{z}{M_{i}}\bigr).$
We adopt the following notation:
$\hat{u}^{t_{i}}_{\alpha}(z)=\frac{\partial\hat{u}^{t_{i}}}{\partial z_{\alpha}}(z),\quad u^{t_{i}}_{\alpha}(w)=\frac{\partial u^{t_{i}}}{\partial w_{\alpha}}(w),$
$\hat{u}^{t_{i}}_{\alpha\bar{\beta}}(z)=\frac{\partial^{2}\hat{u}^{t_{i}}}{\partial z_{\alpha}\partial\bar{z}_{\beta}}(z),\quad u^{t_{i}}_{\alpha\bar{\beta}}(w)=\frac{\partial^{2}u^{t_{i}}}{\partial w_{\alpha}\partial\bar{w}_{\beta}}(w).$
Then we have the scaling relations:
$u^{t_{i}}_{\alpha}(w)=M_{i}\hat{u}^{t_{i}}_{\alpha}(z),\quad u^{t_{i}}_{\alpha\bar{\beta}}(w)=M_{i}^{2}\hat{u}^{t_{i}}_{\alpha\bar{\beta}}(z).$ (57)
By the $C^{0}$-estimate and the complex Hessian estimate for $u^{t_{i}}$ (for large $i$), we obtain
$\sup_{B_{R}(0)}\left(|\hat{u}^{t_{i}}|+|\partial\hat{u}^{t_{i}}|_{\mathcal{E}}+|\Delta\hat{u}^{t_{i}}|_{\mathcal{E}}\right)\leq C.$ (58)
By standard elliptic theory, we have the estimates
$|\hat{u}^{t_{i}}|_{C^{1,\frac{2}{3}}(B_{2R/3}(0))}\leq C.$ (59)
Passing to a subsequence, we find that $\hat{u}^{t_{i}}\to u^{\infty}$ in $C^{1,\frac{1}{2}}(B_{R/2}(0))$. Using a diagonal subsequence argument, we conclude that $\hat{u}^{t_{i}}\to u^{\infty}$ uniformly on compact subsets of $\mathbb{C}^{n}$, and $u^{\infty}$ is a bounded $C^{1,\frac{1}{2}}$ function defined on $\mathbb{C}^{n}$.
Since $u^{t_{i}}$ solves (6), by Lemma 2.2, we have $\sigma_{k}(\chi+t_{i}\omega+\sqrt{-1}\partial\bar{\partial}u^{t_{i}})\geq 0$ for $1\leq k\leq n-1$. It follows that $u^{\infty}$ is (n1)(n-1)-subharmonic. By construction, we have $|\partial u^{\infty}(0)|=1$, so $u^{\infty}$ is nonconstant.
Now, let
$a_{0}=-\liminf_{i\to\infty}\tan\hat{\theta}(t_{i})M_{i}^{2}.$
Recall that $\hat{\theta}(t)\in(\pi-Ct,\pi)$, so $a_{0}\in[0,\infty]$. By passing to a further subsequence, we may assume
$a_{0}=-\lim_{i\to\infty}\tan\hat{\theta}(t_{i})M_{i}^{2}.$ (60)
Case 1: $a_{0}=+\infty$. In this case, we show that $u^{\infty}$ is plurisubharmonic in $\mathbb{C}^{n}$. Since $u^{t_{i}}$ solves the LYZ equation (6), we have
$\chi+t_{i}\omega+\sqrt{-1}\partial\bar{\partial}u^{t_{i}}>\cot\hat{\theta}(t_{i})\omega.$
Then in the coordinate ball $B_{r}(p_{0})$,
$(u^{t_{i}}_{\alpha\bar{\beta}})_{n\times n}\geq(\cot\hat{\theta}(t_{i})-C)g.$ (61)
Now fix $R>0$ and $\epsilon>0$. For sufficiently large $i$ and all $z\in B_{R}(0)$, we have
$(\hat{u}^{t_{i}}_{\alpha\bar{\beta}}(z))_{n\times n}\geq-M_{i}^{-2}(\cot\hat{\theta}(t_{i})-C)g\geq-\epsilon I_{n},$ (62)
where we have used the fact that $\lim_{i\to\infty}M_{i}^{-2}\cot\hat{\theta}(t_{i})=0$. Therefore, $\hat{u}^{t_{i}}+\epsilon|z|^{2}$ is plurisubharmonic in $B_{R}(0)$. Since $\hat{u}^{t_{i}}$ converges uniformly to $u^{\infty}$ on compact sets, it follows that $u^{\infty}+\epsilon|z|^{2}$ is plurisubharmonic in $B_{R}(0)$.
As $R>0$ and $\epsilon>0$ are arbitrary, we conclude that $u^{\infty}$ is plurisubharmonic in $\mathbb{C}^{n}$. However, since $u^{\infty}$ is bounded and plurisubharmonic on $\mathbb{C}^{n}$, it must be constant. This contradicts the fact that $|\partial u^{\infty}(0)|=1$.
Case 2: $0\leq a_{0}<\infty$. Define
$A^{i}(z)=\left(g^{-1}(p_{i}+M_{i}^{-1}z)-I_{n}\right)(\hat{u}^{t_{i}}_{\alpha\bar{\beta}}(z))_{n\times n}+M_{i}^{-2}g^{-1}\chi(p_{i}+M_{i}^{-1}z)+t_{i}M_{i}^{-2}I_{n}.$
Then $A^{i}$ converges uniformly to 0 on any compact subset of $\mathbb{C}^{n}$.
Since $u^{t_{i}}$ solves the LYZ equation (6), we have
$\sigma_{n-1}\left((\hat{u}_{\alpha\bar{\beta}}^{t_{i}})_{n\times n}+A^{i}\right)=M_{i}^{2}\tan\hat{\theta}(t_{i})\sigma_{n}\left((\hat{u}_{\alpha\bar{\beta}}^{t_{i}})_{n\times n}+A^{i}\right)+O(M_{i}^{-2}),$ (63)
and by Lemma 2.2, we have
$\sigma_{k}\left((\hat{u}_{\alpha\bar{\beta}}^{t_{i}})_{n\times n}+A^{i}\right)>0\quad\text{for any }1\leq k\leq n-1.$ (64)
Subcase 2.1: $a_{0}=0$. In this subcase, by (63), (64), and the fact that $\hat{u}^{t_{i}}$ converges locally uniformly to $u^{\infty}$, we conclude that $u^{\infty}$ is (n1)(n-1)-subharmonic and satisfies $\sigma_{n-1}(\sqrt{-1}\partial\bar{\partial}u^{\infty})=0$ in the sense of currents. Hence, applying the Liouville theorem of Dinew-Kolodziej [10], we find that $u^{\infty}$ is constant, which is a contradiction.
Subcase 2.2: $0<a_{0}<\infty$. Since $\hat{u}^{t_{i}}$ converges uniformly to $u^{\infty}$ on compact subsets of $\mathbb{C}^{n}$, we have that $u^{\infty}+a_{0}|z|^{2}$ is plurisubharmonic, (n1)(n-1)-subharmonic, and $u^{\infty}$ satisfies the equation $\sigma_{n}+a_{0}\sigma_{n-1}=0$ in the sense of currents.
Moreover, for any test function $f\in C_{c}^{\infty}(\mathbb{C}^{n})$ with $f\geq 0$ and $\int_{\mathbb{C}^{n}}f=1$, we have
$\int_{\mathbb{C}^{n}}u^{\infty}\sqrt{-1}\partial\bar{\partial}f\wedge\mathcal{E}^{n-1}$ $=\lim_{i\to\infty}\int_{\mathbb{C}^{n}}\hat{u}^{t_{i}}\sqrt{-1}\partial\bar{\partial}f\wedge\mathcal{E}^{n-1}$
$=\lim_{i\to\infty}\int_{\mathbb{C}^{n}}f\sqrt{-1}\partial\bar{\partial}\hat{u}^{t_{i}}\wedge\mathcal{E}^{n-1}$
$\leq C.$
This proves that $\Delta u^{\infty}\leq C$ in the weak sense.
We can then apply the Liouville Theorem 4.1 to $a_{0}^{-1}u^{\infty}$ to conclude that $a_{0}^{-1}u^{\infty}$ is constant, which again leads to a contradiction.
4.5. Proof of Theorems 1.1 and 1.2
Once the gradient estimates for $u^{t}$ are established, we obtain the uniform bound
$\|u^{t}\|_{C^{1}}+\max_{M}|\partial\bar{\partial}u^{t}|\leq C,$
where $C$ is independent of $t$. Following the argument of Collins–Jacob–Yau [7], we derive uniform $C^{2,\alpha}$ estimates for $u^{t}$, also independent of $t$. For any $\beta\in(0,\alpha)$, there exists a subsequence $\{u^{t_{i}}\}$ converging in $C^{2,\beta}$ to a limit $u^{0}\in C^{2,\alpha}$. Thus, $u^{0}$ solves the critical LYZ equation (4). By standard elliptic regularity theory, $u^{0}$ is smooth.
5. Examples: the $3$-dimensional case and $4$-dimensional case
In this section, as corollaries of our main theorem, we solve the $2$-Hessian equation in dimension 3 and a Hessian quotient equation in dimension 4 under weaker conditions than those in the previous works.
5.1. 3-dimensional case
In this case, since
$(\chi_{u}+\sqrt{-1}\omega)^{3}=$ $\chi_{u}^{3}-3\chi_{u}\wedge\omega^{2}+\sqrt{-1}(3\chi_{u}^{2}\wedge\omega-\omega^{3}),$ (65)
the critical LYZ equation is exactly the 2-Hessian equation
$3\chi_{u}^{2}\wedge\omega=\omega^{3}.$
Sun [34] solved the $2$-Hessian equation under the following conditions:
$3\int_{M}\chi^{2}\wedge\omega=\int_{M}\omega^{3},\quad\text{and}\quad\chi\in\Gamma_{2}(M),$
with the following continuity method
$3\chi_{u}^{2}\wedge\omega=\Big(t+(1-t)\frac{3\chi^{2}\wedge\omega}{\omega^{3}}\Big)\omega^{3}.$ (66)
We consider a family of LYZ equations (6)
$3(\chi+t\omega+\sqrt{-1}\partial\bar{\partial}u^{t})^{2}\wedge\omega=$ $\tan\hat{\theta}(t)(\chi+t\omega+\sqrt{-1}\partial\bar{\partial}u^{t})^{3}$
$-3\tan\hat{\theta}(t)(\chi+t\omega+\sqrt{-1}\partial\bar{\partial}u^{t})\wedge\omega^{2}+\omega^{3}.$
As a corollary of our theorem, we solve the 2-Hessian equation under weaker condition instead of the usual condition $\chi\in\Gamma_{2}(M)$.
Corollary 5.1.
Let $(M,\omega)$ be a compact Kähler manifold with dimension $n=3$ and $\chi$ be a real $(1,1)$ form satisfying the following conditions:
  1. (i)
    $3\int_{M}\chi^{2}\wedge\omega=\int_{M}\omega^{3}$, $\int_{M}\chi^{3}<3\int_{M}\chi\wedge\omega^{2}$, and
  2. (ii)
    the subsolution condition
    $\chi\wedge\omega>0\ \ \text{as a $(2,2)$ form}.$ (67)
Then there exists a unique smooth solution $u$ solving
$\sigma_{2}(\chi_{u})=1,$ (68)
$\chi_{u}\in\Gamma_{2}(M),\ \sup_{M}u=0.$ (69)
Proof.
Condition $(i)$ implies the principal argument of $\int_{M}(\chi+\sqrt{-1}\omega)^{3}$ is $\pi$. Condition $(ii)$ means that 0 is the subsolution of the critical LYZ equation in dimension 3. According to Theorem 1.1, there exists a smooth solution solving the critical LYZ equation in dimension 3 which is exactly the 2-Hessian equation $\sigma_{2}=1$.∎
Remark 5.1.
If $\chi\in\Gamma_{2}(M)$, then it satisfies the subsolution condition $(ii)$. But the subsolution condition is weaker than $\chi\in\Gamma_{2}(M)$. We will prove the integral condition $(i)$ is necessary.
Let $\lambda=(\lambda_{1},\lambda_{2},\lambda_{3})$ be the eigenvalues of $\chi_{u}$ with respect to $\omega$ such that $\lambda_{1}\geq\lambda_{2}\geq\lambda_{3}$. Since $\sigma_{2}(\chi_{u})=1$, we have
$\sigma_{3}(\chi_{u})\leq\frac{1}{9}\sigma_{1}(\chi_{u})\sigma_{2}(\chi_{u})=\frac{1}{9}\sigma_{1}(\chi_{u}).$ (70)
Integration by parts, we get
$\int_{M}\chi^{3}=$ $\int_{M}\chi_{u}^{3}\leq\frac{1}{3}\int_{M}{\chi_{u}\wedge\omega^{2}}=\frac{1}{3}\int_{M}{\chi\wedge\omega^{2}}$
$<$ $3\int_{M}{\chi\wedge\omega^{2}},$
which proves the inequality in condition $(i)$.
5.2. $4$-dimesional case
When $n=4$, the critical LYZ equation is the Hessian quotient equation. In this case the subsolution implies $\sigma_{2}(\mu|i)>1$ for $1\leq i\leq 4$, where $\mu=(\mu_{1},\mu_{2},\mu_{3},\mu_{4})$ are eigenvalues of $\chi$ with respect to $\omega$. Assuming further $\chi\in\Gamma_{3}(M)$, Sun [34] and Székelyhidi [36] proved the existence result using different continuity methods.
As a corollary of our result, we prove an existence result for the Hessian quotient equation $\sigma_{3}(\chi_{u})=\sigma_{1}(\chi_{u})$ under weaker conditions than before.
Corollary 5.2.
Let $(M,\omega)$ be a compact Kähler manifold with dimension $n=4$. Let $\chi$ be a real $(1,1)$-form satisfying the following conditions:
  1. (i)
    $\int_{M}\chi^{3}\wedge\omega=\int_{M}\chi\wedge\omega^{3},\int_{M}\chi\wedge\omega^{3}>0,\ \mathrm{Re}\int_{M}(\chi+\sqrt{-1}\omega)^{4}<0,$
  2. (ii)
    $3\chi^{2}\wedge\omega-\omega^{3}>0\ \text{as a $(3,3)$-form}.$
Then there exists a unique smooth function $u$ solving
$\sigma_{3}(\chi_{u})=\sigma_{1}(\chi_{u}),$ (71)
$\chi_{u}\in\Gamma_{3}(M),\sup_{M}u=0.$
Proof.
Condition $(i)$ implies the principal argument of $\int_{M}(\chi+\sqrt{-1}\omega)^{4}$ is $\pi$. To apply Theorem 1.1, we only need to show 0 is the subsolution of the critical LYZ equation. By condition $(ii)$, we have $\sigma_{2}(\chi)>2$ and then $\sigma_{1}^{2}(\chi)\geq\frac{8}{3}\sigma_{2}(\chi)>\frac{16}{3}$. Since there exists $z_{0}\in M$ such that $\sigma_{1}(\chi(z_{0}))=4\frac{\int_{M}\chi\wedge\omega^{3}}{\int_{M}\omega^{4}}>0$, we get $\sigma_{1}(\chi)>0$ on $M$.
Let $\mu=(\mu_{1},\mu_{2},\mu_{3},\mu_{4})$ be the eigenvalues of $\chi$ with respect to $\omega$ such that $\mu_{1}\geq\mu_{2}\geq\mu_{3}\geq\mu_{4}$. We need to prove
$\arctan\mu_{2}+\arctan\mu_{3}+\arctan\mu_{4}>\frac{\pi}{2}.$ (72)
In fact, by condition $(ii)$, we have $\sigma_{2}(\mu|1)=\mu_{2}\mu_{3}+\mu_{2}\mu_{4}+\mu_{3}\mu_{4}>1$. $\sigma_{2}(\mu|1)>0$ and $\mu\in\Gamma_{2}$ implies $\mu_{2}+\mu_{3}>0$ and then $\mu_{4}>\frac{1-\mu_{2}\mu_{3}}{\mu_{2}+\mu_{3}}$. Then we get
$\arctan\mu_{4}>\arctan\frac{1-\mu_{2}\mu_{3}}{\mu_{2}+\mu_{3}}=\frac{\pi}{2}-\arctan\mu_{2}-\arctan\mu_{3},$
which proves (72) and hence 0 is the subsolution of the critical LYZ equation. ∎
We next show condition $(i)$ is necessary.
Lemma 5.1.
Let $u$ be the unique smooth solution of the Hessian quotient equation (71). Then $\chi$ satisfies condition $(i)$.
Proof.
The equality in condition $(i)$ follows from the integration of the equation. Since $\sigma_{1}(\chi_{u})>0$, we have
$\int_{M}\chi\wedge\omega^{3}=\int_{M}\chi_{u}\wedge\omega^{3}>0.$ (73)
Let $\lambda=(\lambda_{1},\lambda_{2},\lambda_{3},\lambda_{4})$ be the eigenvalues of $\chi_{u}$ with respect to $\omega$ such that $\lambda_{1}\geq\lambda_{2}\geq\lambda_{3}\geq\lambda_{4}$. By the MacLarin inequality, we have
$\sigma_{2}\geq 6\frac{\sigma_{3}}{\sigma_{1}}=6.$ (74)
Since $\lambda_{4}\sigma_{1}(\lambda|4)=\lambda_{4}(\lambda_{1}+\lambda_{2}+\lambda_{3})\leq\lambda_{1}\lambda_{2}+\lambda_{2}\lambda_{3}+\lambda_{1}\lambda_{3}=\sigma_{2}(\lambda|4)$, we get
$\sigma_{2}=\sigma_{2}(\lambda|4)+\lambda_{4}\sigma_{1}(\lambda|1)\leq 2\sigma_{2}(\lambda|4),$ (75)
and then $\sigma_{2}(\lambda|4)\geq 3.$ Since $\lambda_{4}=\frac{\sigma_{3}(\lambda|4)-\sigma_{1}(\lambda|4)}{-\sigma_{2}(\lambda|4)+1}$, we get
$\sigma_{4}-\sigma_{2}+1=$ $\lambda_{4}\sigma_{3}(\lambda|4)-\lambda_{4}\sigma_{1}(\lambda|4)-\sigma_{2}(\lambda|4)+1$
$=$ $\frac{\Big(\sigma_{3}(\lambda|4)-\sigma_{1}(\lambda|4)\Big)^{2}}{-\sigma_{2}(\lambda|4)+1}-\sigma_{2}(\lambda|4)+1$
$\leq$ $-2.$
Hence we get
$\mathrm{Re}\int_{M}(\chi+\sqrt{-1}\omega)^{4}=$ $\mathrm{Re}\int_{M}(\chi_{u}+\sqrt{-1}\omega)^{4}$
$=$ $\int_{M}(\sigma_{4}-\sigma_{2}+1)\omega^{4}<0,$
which proves the second inequality in condition $(i)$. ∎

### v2

@@
 Collins-Jacob-Yau [7] suggested studying the LYZ equation when $\theta\leq(n-2)\frac{\pi}{2}$.
-This was later explicitly posed as a question by Li [25], as resolving it is essential for the LYZ equation to induce a Bridgeland stability condition on $D^{b}\text{Coh}(M)$.
+This was later explicitly posed as a question by Li [26], as resolving it is essential for the LYZ equation to induce a Bridgeland stability condition on $D^{b}\text{Coh}(M)$, the bounded derived category of coherent sheaves on $M$.
 In this paper, we solve the critical case; that is, we establish the existence of smooth solutions for the critical LYZ equation:
@@
 In this case, the principal argument of $\int_{M}(\chi+\sqrt{-1}\omega)^{n}$ is $\pi$, and the subsolution condition becomes
-$\min\limits_{1\leq j\leq n}\sum\limits_{\begin{subarray}{c}i=1\\ i\neq j\end{subarray}}^{n}\mathrm{arctan}\lambda_{i}(\chi_{\underline{u}})>\frac{n-3}{2}\pi.$ (5)
+$\min\limits_{1\leq j\leq n}\sum\limits_{\begin{subarray}{c}i=1\\ i\neq j\end{subarray}}^{n}\mathrm{arctan}\lambda_{i}(\chi_{\underline{u}})>{(n-3)\frac{\pi}{2}}.$ (5)
 Theorem 1.1.
@@
 Theorem 1.2.
-Assume the same conditions as in Theorem 1.1 hold.
+Assume the same conditions as in Theorem 1.1.
 Let $u^{t}$ be the smooth solution of (6) with $\sup_{M}u^{t}=0$.
-Then there exist uniform constants $C$ and $t_{0}$ depending only on $(M,\omega,\chi)$ but independent of $t$ such that for any $t\in(0,t_{0})$, we have
+Then there exist uniform constants $C$ and $t_{0}$, depending only on $(M,\omega,\chi)$ but independent of $t$, such that for any $t\in(0,t_{0})$,
 $|u^{t}|_{C^{2,\alpha}(M)}\leq C.$ (7)
-According to Collins-Jacob-Yau [7], $\sup_{M}|u^{t}|$ is uniform bounded.
-Their results then imply that $C^{2,\alpha}$ estimates can be reduced to prove the uniform complex Hessian estimates.
-Our main contribution is the proof of the uniform complex Hessian estimates of Hou-Ma-Wu type
-$\sup\limits_{M}|\sqrt{-1}\partial\bar{\partial}u|_{\omega}\leq C(1+\sup\limits_{M}|\partial u|^{2}_{\omega})$ (8)
-and the uniform gradient estimates via the Dinew-Kolodziej blow-up argument [10].
-However, the situation here is more complicated.
-If the gradient estimates fail, we obtain a $C^{1}$ bounded non-constant function $v$ on $\mathbb{C}^{n}$ which, in the sense of currents, is a weak solution to one of the following equations: $\sigma_{n}=0$, $\sigma_{n-1}=0$, or $\sigma_{n-1}+\sigma_{n}=0$.
-The last case is new and requires a new Liouville-type theorem.
+According to Collins–Jacob–Yau [7], $\sup_{M}|u^{t}|$ is uniformly bounded.
+Their results imply that the $C^{2,\alpha}$ estimates can be reduced to proving uniform complex Hessian estimates.
+Our main contribution is the proof of the uniform complex Hessian estimate of Hou–Ma–Wu type,
+$\sup\limits_{M}|\sqrt{-1}\partial\bar{\partial}u|_{\omega}\leq C\bigl(1+\sup\limits_{M}|\partial u|_{\omega}^{2}\bigr),$ (8)
+and the uniform gradient estimate via the Dinew–Kolodziej blow-up argument [10].
+However, the situation here is more subtle.
+If the gradient estimate fails, we obtain a bounded non-constant $C^{1}$ function $v$ on $\mathbb{C}^{n}$ that, in the sense of currents, is a weak solution to one of the following equations:
+$\sigma_{n}=0,\quad\sigma_{n-1}=0,\quad\text{or}\quad\sigma_{n-1}+\sigma_{n}=0.$
+The last case is new and requires a corresponding Liouville-type theorem.
 Note that $v+|z|^{2}$ is plurisubharmonic and $\Delta v$ is uniformly bounded in the weak sense (see Remark 4.2).
-When we follow the argument in [10, 36], we find the first case can be handled similarly.
-For the second case, we need to prove the standard mollification $v]_{r}$ of $v$ and $\frac{1}{3}[(v^{\epsilon})^{2}]_{r}-\tau_{0}|z|^{2}$ are subsolutions, where $v^{\epsilon}=[v]_{\epsilon}$ and $\tau_{0}>0$ is small enough.
-To address this issue, our starting point is that a smooth function $v$ satisfies equation $\sigma_{n-1}+\sigma_{n}=0$ (resp.
-$\geq 0$) on $\mathbb{C}^{n}$ is equivalent to the function $v+|z_{n+1}|^{2}$ satisfies equation $\sigma_{n}=0$ (resp.
+When following the argument in [10, 38], we find that the first case can be handled similarly.
+For the second case, we need to show that the standard mollification $v]_{r}$ of $v$, as well as $\frac{1}{3}[(v^{\epsilon})^{2}]_{r}-\tau_{0}|z|^{2}$, are subsolutions, where $v^{\epsilon}=[v]_{\epsilon}$ and $\tau_{0}>0$ is sufficiently small.
+To address this issue, our starting observation is that a smooth function $v$ satisfies $\sigma_{n-1}+\sigma_{n}=0$ (resp.
+$\geq 0$) on $\mathbb{C}^{n}$ if and only if the function $v+|z_{n+1}|^{2}$ satisfies $\sigma_{n}=0$ (resp.
 $\geq 0$) on $\mathbb{C}^{n+1}$, where $z_{n+1}$ is a new complex variable.
+Based on this observation, we use potential theory to obtain the required Liouville-type theorem (see Theorem 4.1).
 Finally, we find some applications of the main result. In general, the critical LYZ equation has the form
@@
 Corollary 1.1.
-Let $(M,\omega)$ be a compact Kähler manifolds with dimension $n=3$ and $\chi$ be a closed real $(1,1)$ form satisfying $\chi\wedge\omega>0$ as a positive $(2,2)$ form and
+Let $(M,\omega)$ be a compact Kähler manifolds with dimension $n=3$ and $\chi$ be a closed real $(1,1)$-form satisfying $\chi\wedge\omega>0$ as a positive $(2,2)$-form and
 $3\int_{M}\chi^{2}\wedge\omega=\int_{M}\omega^{3},\ \ \int_{M}\chi^{3}<3\int_{M}\chi\wedge\omega^{2}.$
@@
 $\sigma_{3}(\chi_{u})=\sigma_{1}(\chi_{u}).$
-Another consequence of our result is to solve the above Hessian quotient equation in dimension 4 under weaker conditions than the one by Sun [34] and Székelyhidi [36].
+Another consequence of our result is to solve the above Hessian quotient equation in dimension 4 under weaker conditions than those by Sun [36] and Székelyhidi [38].
 Corollary 1.2.
-Let $(M,\omega)$ be a compact Kähler manifold with dimension $n=4$. Let $\chi$ be a closed real $(1,1)$-form satisfying $3\chi^{2}\wedge\omega-\omega^{3}>0\ \text{as a positive $(3,3)$ form}$ and
+Let $(M,\omega)$ be a compact Kähler manifold with dimension $n=4$. Let $\chi$ be a closed real $(1,1)$-form satisfying $3\chi^{2}\wedge\omega-\omega^{3}>0\ \text{as a positive $(3,3)$-form}$ and
 $\int_{M}\chi^{3}\wedge\omega=$ $\int_{M}\chi\wedge\omega^{3},$
@@
 $\chi_{u}\in\Gamma_{3}(M),\ \sup_{M}u=0.$
-The Hessian quotient equation $\sigma_{k}(\chi_{u})=c\sigma_{l}(\chi_{u})$ with $0\leq l<k\leq n$ was solved by Sun [34] and Székelyhidi [36] by different continuity methods.
+We remark that the Hessian quotient equation is of the form $\sigma_{k}(\chi_{u})=c\sigma_{l}(\chi_{u})$ with $0\leq l<k\leq n$.
+It reduces to the $J$-equation when $k=n$ and l=n1l=n-1 which was solved by Song-Weinkove [35] using the flow method, and the case $k=n,1\leq l\leq n-1$ was solved by Fang-Lai-Ma [11].
+The general case was subsequently solved by Sun [36] and Székelyhidi [38] via different continuity methods.
 The organization of this paper is as follows.
@@
 We also give the $C^{0}$-estimate for $u^{t}$ using results from [7].
-In Section 3, we first proves a key inequality implied by the subsolution condition, then we establish the uniform complex Hessian estimate of Hou-Ma-Wu-type which is independent of $(\theta(t)-(n-2)\frac{\pi}{2})^{-1}$).
+In Section 3, we first prove a key inequality implied by the subsolution condition, then we establish the uniform complex Hessian estimate of Hou-Ma-Wu-type which is independent of $(\theta(t)-(n-2)\frac{\pi}{2})^{-1}$.
 In Section 4, based on the complex Hessian estimate, we prove the gradient estimate by a blow up argument.
@@
 2.1. The $k$-Hessian operators and the special Lagrangian operator
-For any vector $\lambda\in\mathbb{R}^{n}$ and $1\leq k\leq n$, we define
-$\sigma_{k}(\lambda)=\sum\limits_{1\leq i_{1}<\cdots<i_{k}\leq n}\lambda_{i_{1}}\cdots\lambda_{i_{k}}.$ (10)
-For a real $(1,1)$-form $\alpha$, the $k$-Hessian operator of $\alpha$ with respect to $\omega$ is defined by
+For an $n$-tuple $\lambda=(\lambda_{1},\dots,\lambda_{n})\in\mathbb{R}^{n}$ and $1\leq k\leq n$, we define
+$\sigma_{k}(\lambda)=\sum_{1\leq i_{1}<\cdots<i_{k}\leq n}\lambda_{i_{1}}\cdots\lambda_{i_{k}}.$ (10)
+For a real $(1,1)$-form $\alpha$, the $k$-Hessian operator of $\alpha$ with respect to the Kähler form $\omega$ is defined by
 $\sigma_{k}(\alpha):=C_{n}^{k}\frac{\alpha^{k}\wedge\omega^{n-k}}{\omega^{n}}.$ (11)
-In local coordinates, let $\alpha=\sqrt{-1}\alpha_{i\bar{j}}dz_{i}\wedge d\bar{z}_{j}$ and $\omega=\sqrt{-1}g_{i\bar{j}}dz_{i}\wedge d\bar{z}_{j}$, we have $\sigma_{k}{(\alpha)}=\sigma_{k}(\lambda(g^{j\bar{l}}\alpha_{i\bar{l}}))$.
-For simplicity, we also write $\sigma_{k}(g^{j\bar{l}}\alpha_{i\bar{l}})$ to represent $\sigma_{k}(\lambda(g^{j\bar{l}}\alpha_{i\bar{l}}))$ .
+In local coordinates, writing $\alpha=\sqrt{-1}\alpha_{i\bar{j}}\,dz_{i}\wedge d\bar{z}_{j}$ and $\omega=\sqrt{-1}g_{i\bar{j}}\,dz_{i}\wedge d\bar{z}_{j}$, we have
+$\sigma_{k}(\alpha)=\sigma_{k}\!\bigl(\lambda(g^{j\bar{l}}\alpha_{i\bar{l}})\bigr).$
+For simplicity, we also denote $\sigma_{k}(g^{j\bar{l}}\alpha_{i\bar{l}})$ for $\sigma_{k}\!\bigl(\lambda(g^{j\bar{l}}\alpha_{i\bar{l}})\bigr)$.
 We need the following convex cones:
@@
 […]
-[…]
-We have the concavity property for Hessian operators.
+$\mathcal{M}(\Gamma_{n-1})$ $:=\{A\mid\text{$A$ is an $n\times n$ Hermitian matrix and }\lambda(A)\in\Gamma_{n-1}\}.$
+The Hessian operators possess the following concavity property.
 Lemma 2.1.
-The function ${\sigma_{n-1}^{\frac{1}{n-1}}}(\lambda)$ is concave in $\Gamma_{n-1}$. The function ${\sigma_{n-1}^{\frac{1}{n-1}}}(A)$ is concave in $\mathcal{M}(\Gamma_{n-1})$.
+The function $\sigma_{n-1}^{\frac{1}{n-1}}(\lambda)$ is concave in $\Gamma_{n-1}$. Consequently, the function $\sigma^{\frac{1}{n-1}}_{n-1}(A)$ is concave in $\mathcal{M}(\Gamma_{n-1})$.
 Let
-$\theta(\lambda):=\sum_{i=1}^{n}\arctan\lambda_{i}\quad\textup{for}\ \lambda=(\lambda_{1},\dots,\lambda_{n})\in\mathbb{R}^{n}$
-and
-$\Gamma^{\tau}:=\{\lambda\in\mathbb{R}^{n}\mid\theta(\lambda)>\tau\}\subset\mathbb{R}^{n}\quad\textup{for $\tau\in[(n-2)\frac{\pi}{2},n\frac{\pi}{2})$}.$
-We have the following useful facts.
-Lemma 2.2 (Yuan [42], Wang-Yuan [40]).
-If $\theta(\lambda)\geq\tau\in[(n-2)\frac{\pi}{2},n\frac{\pi}{2})$ for $\lambda=(\lambda_{1},\dots,\lambda_{n})$ with $\lambda_{1}\geq\lambda_{2}\geq\cdots\geq\lambda_{n}$, then we have
+$\theta(\lambda):=\sum_{i=1}^{n}\arctan\lambda_{i}\quad\text{for }\lambda=(\lambda_{1},\dots,\lambda_{n})\in\mathbb{R}^{n},$ (12)
+and define the cones
+$\Gamma^{\tau}:=\{\lambda\in\mathbb{R}^{n}\mid\theta(\lambda)>\tau\}\subset\mathbb{R}^{n}\quad\text{for }\tau\in\bigl[(n-2)\tfrac{\pi}{2},n\tfrac{\pi}{2}\bigr).$
+The following lemma collects some key properties of these cones.
+Lemma 2.2 (Yuan [44], Wang–Yuan [42]).
+Assume that $\theta(\lambda)\geq\tau$ for some $\tau\in[(n-2)\frac{\pi}{2},n\frac{\pi}{2})$, and that the components of $\lambda=(\lambda_{1},\dots,\lambda_{n})$ satisfy $\lambda_{1}\geq\lambda_{2}\geq\cdots\geq\lambda_{n}$.
+Then the following hold:
   1. (1)
-    $\lambda_{1}\geq\lambda_{2}\geq\cdots\geq\lambda_{n-1}>0$, and $\lambda_{n-1}\geq|\lambda_{n}|$,
+    $\lambda_{1}\geq\lambda_{2}\geq\cdots\geq\lambda_{n-1}>0$ and $\lambda_{n-1}\geq|\lambda_{n}|$;
   2. (2)
-    $\lambda_{1}+(n-1)\lambda_{n}\geq 0$,
+    $\lambda_{1}+(n-1)\lambda_{n}\geq 0$;
   3. (3)
-    $\lambda\in\Gamma_{n-1}$, and
+    $\lambda\in\Gamma_{n-1}$;
   4. (4)
@@
 $2.$ $C^{0}$-estimates
-Let $\hat{\theta}(t)$ be the principal argument of the complex number $\int_{M}(\chi+t\omega+\sqrt{-1}\omega)^{n}$ and $\theta(t)=(n-2)\frac{\pi}{2}-\hat{\theta}(t)$.
-We first prove $\hat{\theta}(t)$ is sufficiently close to $\pi$ for sufficiently small $t>0$.
+Let $\hat{\theta}(t)$ denote the principal argument of the complex number $\int_{M}(\chi+t\omega+\sqrt{-1}\omega)^{n}$, and set $\theta(t)=(n-2)\frac{\pi}{2}-\hat{\theta}(t)$.
+We first show that $\hat{\theta}(t)$ is close to $\pi$ when $t>0$ is sufficiently small.
 Lemma 2.3.
-If $\underline{u}$ is the subsolution of the critical LYZ equation (4), then there exists uniform positive constants $t_{0}$ and $C$ depending only on $\chi$ and $\omega$ such that $\hat{\theta}(t)\in(\pi-2Ct,\pi-Ct)$ for any $t\in(0,t_{0})$.
+Assume that $\underline{u}$ is a subsolution of the critical LYZ equation (4).
+Then there exist uniform positive constants $t_{0}$ and $C$, depending only on $\chi$ and $\omega$, such that for all $t\in(0,t_{0})$,
+$\hat{\theta}(t)\in\bigl(\pi-2Ct,\;\pi-Ct\bigr).$
 Proof.
-By direct computation, we have
+By a direct computation, we have
 $(\chi+t\omega+\sqrt{-1}\omega)^{n}=(\chi+\sqrt{-1}\omega)^{n}+t\,n(\chi+\sqrt{-1}\omega)^{n-1}\wedge\omega+O(t^{2}).$
-Since
-$\mathrm{Re}\int_{M}(\chi+\sqrt{-1}\omega)^{n}<0,$
-for sufficiently small $t$, we have
-$\mathrm{Re}\int_{M}(\chi+t\omega+\sqrt{-1}\omega)^{n}<\mathrm{Re}\int_{M}(\chi+\sqrt{-1}\omega)^{n}+Ct<0.$ (12)
-Next we want to prove the imaginary part is positive. The subsolution condition implies the positivity
+Since $\mathrm{Re}\int_{M}(\chi+\sqrt{-1}\omega)^{n}<0$, it follows that for sufficiently small $t$,
+$\mathrm{Re}\int_{M}(\chi+t\omega+\sqrt{-1}\omega)^{n}<\mathrm{Re}\int_{M}(\chi+\sqrt{-1}\omega)^{n}+Ct<0.$
+We now show that the imaginary part is positive. The subsolution condition implies the positivity
 $\mathrm{Im}\,(\chi_{\underline{u}}+\sqrt{-1}\omega)^{n-1}>0.$
-Then for sufficiently small $t>0$, we get
-$\mathrm{Im}\int_{M}(\chi+\sqrt{-1}\omega)^{n-1}\wedge\omega=\mathrm{Im}\int_{M}(\chi_{\underline{u}}+\sqrt{-1}\omega)^{n-1}\wedge\omega>0$ (13)
-Thus we get
+Hence, for sufficiently small $t>0$,
+$\mathrm{Im}\int_{M}(\chi+\sqrt{-1}\omega)^{n-1}\wedge\omega=\mathrm{Im}\int_{M}(\chi_{\underline{u}}+\sqrt{-1}\omega)^{n-1}\wedge\omega>0.$ (13)
+Consequently,
 $\mathrm{Im}\int_{M}(\chi+t\omega+\sqrt{-1}\omega)^{n}$ $=\mathrm{Im}\int_{M}(\chi+\sqrt{-1}\omega)^{n}$
-$+t\mathrm{Im}\int_{M}(\chi+\sqrt{-1}\omega)^{n-1}\wedge\omega+O(t^{2})$
+$\quad+t\,\mathrm{Im}\int_{M}(\chi+\sqrt{-1}\omega)^{n-1}\wedge\omega+O(t^{2})$
 $=t\,\mathrm{Im}\int_{M}(\chi+\sqrt{-1}\omega)^{n-1}\wedge\omega+O(t^{2})$
 $>0,$
-where in the last inequality we use (13).
-Then there exists a uniform positive constant $C$ depending only on $\chi$ and $\omega$ such that the principal argument of the complex number $\int_{M}(\chi+t\omega+\sqrt{-1}\omega)^{n}$ satisfies
-$\hat{\theta}(t)=\mathrm{arccot}\Big(\frac{\mathrm{Re}\int_{M}(\chi+\sqrt{-1}\omega)^{n}+O(t)}{t\mathrm{Im}\int_{M}(\chi+\sqrt{-1}\omega)^{n-1}\wedge\omega+O(t^{2})}\Big)\in(\pi-2Ct,\pi-Ct).$
-Hence by Lemma 2.3, we have, for sufficiently small $t>0$,
-$\theta(t)\in\big((n-2)\frac{\pi}{2}+Ct,(n-2)\frac{\pi}{2}+2Ct\big),$ (14)
-and then
-$\min\limits_{1\leq j\leq n}\sum\limits_{\begin{subarray}{c}i=1\\ i\neq j\end{subarray}}^{n}\mathrm{arctan}\lambda_{i}(\chi_{\underline{u}}+t\omega)>(n-3)\frac{\pi}{2}+2Ct>\theta(t)-\frac{\pi}{2},$ (15)
-which means that $\underline{u}$ is also a subsolution of the equations (6) for sufficiently small $t>0$.
-According to Lin [26] (see also Sun [35]), for every small $t>0$, there exists a smooth solution $u^{t}$ solving the supercritical LYZ equation (6).
-According to Proposition 3.6 in [7], we get the uniform $C^{0}$ estimates of $u^{t}$.
-Lemma 2.4 (Collins-Jacob-Yau [7]).
-Let $\underline{u}$ be the subsolution of the critical LYZ equation (4) and $u^{t}$ be the solution of the LYZ equation (6) with $\sup_{M}u^{t}=0$.
-Then there exist uniform constants $t_{0}$ and $C$ such that for any $t\in(0,t_{0})$, we have
-$\sup_{M}|u^{t}|\leq C.$
+where the last inequality uses (13).
+Therefore, there exists a uniform constant $C>0$, depending only on $\chi$ and $\omega$, such that the principal argument $\hat{\theta}(t)$ of $\int_{M}(\chi+t\omega+\sqrt{-1}\omega)^{n}$ satisfies
+$\hat{\theta}(t)=\pi-\arctan\!\left(\frac{t\,\mathrm{Im}\int_{M}(\chi+\sqrt{-1}\omega)^{n-1}\wedge\omega+O(t^{2})}{-\mathrm{Re}\int_{M}(\chi+\sqrt{-1}\omega)^{n}+O(t)}\right)\in(\pi-2Ct,\;\pi-Ct)$
+for all sufficiently small $t>0$. ∎
+Hence, for sufficiently small $t>0$, Lemma 2.3 yields
+$\theta(t)\in\bigl((n-2)\tfrac{\pi}{2}+Ct,\;(n-2)\tfrac{\pi}{2}+2Ct\bigr),$ (14)
+and consequently,
+$\min_{1\leq j\leq n}\sum_{\begin{subarray}{c}i=1\\ i\neq j\end{subarray}}^{n}\arctan\lambda_{i}(\chi_{\underline{u}}+t\omega)>(n-3)\frac{\pi}{2}+2Ct>\theta(t)-\frac{\pi}{2}.$ (15)
+This shows that $\underline{u}$ remains a subsolution of equation (6) for all sufficiently small $t>0$.
+By the result of Lin [27] (see also Sun [37]), for each such $t$, there exists a smooth solution $u^{t}$ to the supercritical LYZ equation (6).
+Moreover, Proposition 3.6 of [7] provides a uniform $C^{0}$ estimate for these solutions.
+Lemma 2.4 (Collins–Jacob–Yau [7]).
+Let $\underline{u}$ be a subsolution of the critical LYZ equation (4), and let $u^{t}$ be the solution of (6) normalized by $\sup_{M}u^{t}=0$.
+Then there exist constants $t_{0}>0$ and $C>0$, depending only on the background data $(\chi,\omega)$ and $\underline{u}$, such that for all $t\in(0,t_{0})$,
+$\|u^{t}\|_{C^{0}(M)}\leq C.$
 3. Complex Hessian estimates
 3.1. The linearized operator
-We first give some useful formulae for the complex Hessian estimate in [11].
+We first collect some useful formulas for the complex Hessian estimate from [12].
 Lemma 3.1.
-Let $u^{t}$ be the solution of the LYZ equation (6).
-(i) The linearized operator $\mathcal{L}$ of the LYZ equation (6) has the form
-$\mathcal{{L}}(v)=F^{i\bar{j}}v_{i\bar{j}},$
-where $(F^{i\bar{j}})_{n\times n}$ is the inverse matrix of $wg^{-1}w+g$ with $g=(g_{i\bar{j}})_{n\times n}$, $w=(w_{i\bar{j}})_{n\times n}$ for $w_{i\bar{j}}=\chi_{i\bar{j}}+tg_{i\bar{j}}+u_{i\bar{j}}$.
-(ii) Differentiating the LYZ equation (6), we have
-$F^{i\bar{j}}w_{i\bar{j},p}=0,$ (16)
-$F^{i\bar{j}}w_{i\bar{j},p\bar{p}}=F^{i\bar{l}}F^{k\bar{j}}w_{i\bar{j},p}(w_{k\bar{m},\bar{p}}g^{r\bar{m}}w_{r\bar{l}}+w_{k\bar{m}}g^{r\bar{m}}w_{r\bar{l},\bar{p}}),$ (17)
-where we used covariant derivatives.
-3.2. Key inequalities by the subsolution condition
-The subsolution condition is very important to the $C^{2}$-estimates of solutions of fully nonlinear elliptic PDEs, one can see Guan [12] and Székelyhidi [36].
-Similar to [12, 36], we prove key inequalities which will be used in the second order estimate.
-Fortunately, we can get a uniform constant $\delta_{0}$ in the following lemma.
+Let $u^{t}$ be a solution of the LYZ equation (6). Then:
+  1. (i)
+    The linearized operator $\mathcal{L}$ of (6) takes the form
+    $\mathcal{L}(v)=\sum_{i,j}F^{i\bar{j}}v_{i\bar{j}},$
+    where the matrix $(F^{i\bar{j}})$ is the inverse of $wg^{-1}w+g$. Here $g=(g_{i\bar{j}})$ and $w=(w_{i\bar{j}})$ with $w_{i\bar{j}}=\chi_{i\bar{j}}+tg_{i\bar{j}}+u^{t}_{i\bar{j}}$.
+  2. (ii)
+    Differentiating the LYZ equation (6) (using covariant derivatives) yields
+    $\sum_{i,j}F^{i\bar{j}}w_{i\bar{j},p}=0,$ (16)
+    $\sum_{i,j}F^{i\bar{j}}w_{i\bar{j},p\bar{p}}=\sum_{i,j,k,l}F^{i\bar{l}}F^{k\bar{j}}w_{i\bar{j},p}\bigl(w_{k\bar{m},\bar{p}}g^{r\bar{m}}w_{r\bar{l}}+w_{k\bar{m}}g^{r\bar{m}}w_{r\bar{l},\bar{p}}\bigr).$ (17)
+3.2. Key inequalities from the subsolution condition
+The subsolution condition plays a crucial role in the $C^{2}$-estimates for solutions of fully nonlinear elliptic PDEs, as demonstrated in the works of Guan [13] and Székelyhidi [38].
+Following the approach in these references, we establish key inequalities that will be essential for the second-order estimate.
+A favorable feature is that the constant $\delta_{0}$ appearing below can be chosen uniformly.
 Lemma 3.2.
-Let $\mu\in\mathbb{R}^{n}$ satisfy $A(\mu):=\min_{1\leq j\leq n}\sum_{i\neq j}\arctan\mu_{i}>A_{0}>(n-3)\frac{\pi}{2}$.
-Then for all $\lambda\in\mathbb{R}^{n}$ such that $\theta(\lambda)=\theta(t)$, there exist positive constants $t_{0}$ and $\delta_{0}$ depending only on $n,\mu,A_{0}$ and $C$ in Lemma 2.3 such that for any $t\in(0,t_{0})$, one of the following holds:
+Let $\mu\in\mathbb{R}^{n}$ satisfy
+$A(\mu):=\min_{1\leq j\leq n}\sum_{i\neq j}\arctan\mu_{i}\;>\;A_{0}\;>\;(n-3)\frac{\pi}{2}.$
+Then, for every $\lambda\in\mathbb{R}^{n}$ with $\theta(\lambda)=\theta(t)$, there exist positive constants $t_{0}$ and $\delta_{0}$, depending only on $n$, $\mu$, $A_{0}$, and the constant $C$ from Lemma 2.3, such that for any $t\in(0,t_{0})$, one of the following statements holds:
   1. (i)
-    $\sum_{i=1}^{n}\frac{\mu_{i}-\lambda_{i}}{1+\lambda_{i}^{2}}\geq\delta_{0}\sum_{i=1}^{n}\frac{1}{1+\lambda_{i}^{2}},\ \text{or}$
+    $\sum_{i=1}^{n}\frac{\mu_{i}-\lambda_{i}}{1+\lambda_{i}^{2}}\geq\delta_{0}\sum_{i=1}^{n}\frac{1}{1+\lambda_{i}^{2}}$; or
   2. (ii)
-    $\frac{1}{1+\lambda_{j}^{2}}\geq\delta_{0}\sum_{i=1}^{n}\frac{1}{1+\lambda_{i}^{2}}\ \ \text{for}\ 1\leq j\leq n.$
+    $\frac{1}{1+\lambda_{j}^{2}}\geq\delta_{0}\sum_{i=1}^{n}\frac{1}{1+\lambda_{i}^{2}}$ for $1\leq j\leq n$.
 Proof.
-Since $A(\mu)>A_{0}>(n-3)\frac{\pi}{2}$, we can choose a small positive constant $\delta$ such that
-$A(\mu-2\delta{{\bf{1}}})>\frac{A_{0}+(n-3)\frac{\pi}{2}}{2},$
-where ${\bf{1}}=(1,\cdots,1)\in\mathbb{R}^{n}$.
-For $\lambda\in\mathbb{R}^{n}$ satisfying $\theta(\lambda)=\theta(t)$, we assume $\lambda_{1}\geq\cdots\geq\lambda_{n}$.
-We know ${\bf{n}}=(\frac{1}{1+\lambda_{1}^{2}},\cdots,\frac{1}{1+\lambda_{n}^{2}})$ is an inward normal vector of $\Gamma^{\theta(t)}$ at $\lambda$.
-Let $T_{\lambda}$ be the tangent plane of the hypersurface $\partial\Gamma^{\theta(t)}$ at the point $\lambda$.
-Case 1: $\big((\mu-\delta{{{\bf 1}}})-\lambda\big)\cdot{\bf{n}}\geq 0$. In this case, we have
+Given that $A(\mu)>A_{0}>(n-3)\frac{\pi}{2}$, we can choose a small positive constant $\delta>0$ such that
+$A(\mu-2\delta\mathbf{1})>\frac{A_{0}+(n-3)\frac{\pi}{2}}{2},$
+where $\mathbf{1}=(1,\ldots,1)\in\mathbb{R}^{n}$.
+For $\lambda\in\mathbb{R}^{n}$ satisfying $\theta(\lambda)=\theta(t)$, assume without loss of generality that $\lambda_{1}\geq\cdots\geq\lambda_{n}$. Recall that the vector
+$\mathbf{n}=\Bigl(\frac{1}{1+\lambda_{1}^{2}},\cdots,\frac{1}{1+\lambda_{n}^{2}}\Bigr)$
+is an inward normal to $\Gamma^{\theta(t)}$ at $\lambda$. Let $T_{\lambda}$ denote the tangent plane to the hypersurface $\partial\Gamma^{\theta(t)}$ at $\lambda$.
+Case 1: $\bigl((\mu-\delta\mathbf{1})-\lambda\bigr)\cdot\mathbf{n}\geq 0$. In this case, we have
 $\sum_{i=1}^{n}\frac{\mu_{i}-\lambda_{i}}{1+\lambda_{i}^{2}}\geq\delta\sum_{i=1}^{n}\frac{1}{1+\lambda_{i}^{2}}.$ (18)
-Case 2: $\big((\mu-\delta{{{{{\bf 1}}}}})-\lambda\big)\cdot{\bf{n}}<0$.
-We can show there exists a positive constant $s$ such that $\mu-\delta{{{{{\bf 1}}}}}+se_{1}\in{\Gamma^{\theta(t)}}.$ Indeed, we have
-$\theta(\mu-\delta{{{\bf 1}}}+se_{1})=$ $\arctan(\mu_{1}-\delta+s)+\sum_{i=2}^{n}\arctan(\mu_{i}-\delta)$
-$>$ $\arctan(\mu_{1}-\delta+s)+A(\mu-\delta{{{\bf 1}}})$
-$\geq$ $\arctan(\mu_{1}-\delta+s)+\frac{1}{2}(A_{0}+(n-3)\frac{\pi}{2})$
-$>$ $\theta(t)$
-if we choose $s=|\delta-\mu_{1}|+\tan\big(\frac{\pi}{2}-\frac{1}{4}(A_{0}-(n-3)\frac{\pi}{2})\big)$ and assume $0<t<t_{0}$ for $t_{0}=\frac{1}{8C}(A_{0}-(n-3)\frac{\pi}{2})$.
-The convexity of $\Gamma^{\theta(t)}$ implies $\mu-\delta{{{\bf 1}}}$ and $\mu-\delta{{{\bf 1}}}+se_{1}$ are on two sides of the tangent plane $T_{\lambda}$.
-Then there exists $a_{0}\in(0,1)$ such that $a_{0}(\mu-\delta{{{\bf 1}}}+se_{1})+(1-a_{0})(\mu-\delta{{{\bf 1}}})=\mu-\delta{{{\bf 1}}}+a_{0}se_{1}\in\mathrm{T_{\lambda}}$ .
-We also have $\mu-\delta{{{\bf 1}}}+a_{0}se_{1}+se_{1}-\delta e_{n}\in\Gamma^{\theta(t)}$ which follows from
-$\theta(\mu-\delta{{{\bf 1}}}+a_{0}se_{1}+se_{1}-\delta e_{n})$
-$=$ $\arctan(\mu_{1}-\delta+a_{0}s+s)$
-$+\sum_{i=2}^{n-1}\arctan(\mu_{i}-\delta)+\arctan(\mu_{n}-2\delta)$
-$\geq$ $\arctan(\mu_{1}-\delta+a_{0}s+s)+A(\mu-2\delta\bf{1})$
-$\geq$ $\arctan(\mu_{1}-\delta+s)+\frac{1}{2}(A_{0}+(n-3)\frac{\pi}{2})$
+Case 2: $\bigl((\mu-\delta\mathbf{1})-\lambda\bigr)\cdot\mathbf{n}<0$.
+We claim that there exists a positive constant $s$ such that $\mu-\delta\mathbf{1}+se_{1}\in\Gamma^{\theta(t)}$.
+Indeed, observe that
+$\theta(\mu-\delta\mathbf{1}+se_{1})$ $=\arctan(\mu_{1}-\delta+s)+\sum_{i=2}^{n}\arctan(\mu_{i}-\delta)$
+$>\arctan(\mu_{1}-\delta+s)+A(\mu-\delta\mathbf{1})$
+$\geq\arctan(\mu_{1}-\delta+s)+\frac{1}{2}\Bigl(A_{0}+(n-3)\frac{\pi}{2}\Bigr).$
+To ensure $\theta(\mu-\delta\mathbf{1}+se_{1})>\theta(t)$, it suffices to choose $s$ satisfying
+$\arctan(\mu_{1}-\delta+s)>\theta(t)-\frac{1}{2}\Bigl(A_{0}+(n-3)\frac{\pi}{2}\Bigr).$
+Take
+$s=|\delta-\mu_{1}|+\tan\Bigl(\frac{\pi}{2}-\frac{1}{4}\bigl(A_{0}-(n-3)\frac{\pi}{2}\bigr)\Bigr),$
+and assume $0<t<t_{0}$ with $t_{0}=\frac{1}{8C}\bigl(A_{0}-(n-3)\frac{\pi}{2}\bigr)$. Then the desired inequality holds.
+The convexity of $\Gamma^{\theta(t)}$ implies that the points $\mu-\delta\mathbf{1}$ and $\mu-\delta\mathbf{1}+s\mathbf{e}_{1}$ lie on opposite sides of the tangent plane $T_{\lambda}$.
+Hence, there exists $a_{0}\in(0,1)$ such that the convex combination
+$a_{0}(\mu-\delta\mathbf{1}+s\mathbf{e}_{1})+(1-a_{0})(\mu-\delta\mathbf{1})=\mu-\delta\mathbf{1}+a_{0}s\mathbf{e}_{1}$
+lies on $T_{\lambda}$.
+We now show that $\mu-\delta\mathbf{1}+a_{0}s\mathbf{e}_{1}+s\mathbf{e}_{1}-\delta\mathbf{e}_{n}\in\Gamma^{\theta(t)}$. Indeed,
+$\theta\bigl(\mu-\delta\mathbf{1}+a_{0}s\mathbf{e}_{1}+s\mathbf{e}_{1}-\delta\mathbf{e}_{n}\bigr)$
+$=\arctan(\mu_{1}-\delta+a_{0}s+s)+\sum_{i=2}^{n-1}\arctan(\mu_{i}-\delta)+\arctan(\mu_{n}-2\delta)$
+$\geq\arctan(\mu_{1}-\delta+a_{0}s+s)+A(\mu-2\delta\mathbf{1})$
+$\geq\arctan(\mu_{1}-\delta+s)+\frac{1}{2}\Bigl(A_{0}+(n-3)\frac{\pi}{2}\Bigr)$
 $>\theta(t).$
-Thus we have
-$0\leq$ $\big((\mu-\delta{{{\bf 1}}}+a_{0}se_{1}+se_{1}-\delta e_{n}\big)-(\mu-\delta{{{\bf 1}}}+a_{0}se_{1})\big)\cdot\bf{n}$
-$=$ $(se_{1}-\delta e_{n})\cdot\bf n$
+Thus, we have
+$0$ $\leq\bigl((\mu-\delta\mathbf{1}+a_{0}s\mathbf{e}_{1}+s\mathbf{e}_{1}-\delta\mathbf{e}_{n})-(\mu-\delta\mathbf{1}+a_{0}s\mathbf{e}_{1})\bigr)\cdot\mathbf{n}$
+$=(s\mathbf{e}_{1}-\delta\mathbf{e}_{n})\cdot\mathbf{n}$
 $=\frac{s}{1+\lambda_{1}^{2}}-\frac{\delta}{1+\lambda_{n}^{2}},$
-or $\frac{1}{1+\lambda_{1}^{2}}\geq\delta s^{-1}\frac{1}{1+\lambda_{n}^{2}}$. Since $\lambda_{1}\geq\cdots\geq\lambda_{n-1}\geq|\lambda_{n}|$ by Lemma 2.2, we get
-$\frac{1}{1+\lambda_{j}^{2}}\geq\frac{\delta}{sn}\sum_{i=1}^{n}\frac{1}{1+\lambda_{i}^{2}},1\leq j\leq n.$ (19)
-In view of (18) and (19), if we choose $\delta_{0}=\frac{\delta}{ns}$, we finish the proof. ∎
-By the above lemma, we get the following important lemma and the estimate here does not depend on $(\theta(t)-(n-2)\frac{\pi}{2})^{-1}$.
+which is equivalent to
+$\frac{1}{1+\lambda_{1}^{2}}\geq\frac{\delta}{s}\cdot\frac{1}{1+\lambda_{n}^{2}}.$ (19)
+By Lemma 2.2, we have $\lambda_{1}\geq\cdots\geq\lambda_{n-1}\geq|\lambda_{n}|$. Consequently, for each $j=1,\dots,n$,
+$\frac{1}{1+\lambda_{n}^{2}}\geq\frac{1}{1+\lambda_{j}^{2}}.$
+Combining with (19), we obtain
+$\frac{1}{1+\lambda_{1}^{2}}\geq\frac{\delta}{s}\cdot\frac{1}{1+\lambda_{j}^{2}},\quad\textup{for all $1\leq j\leq n$.}$
+Summing over $j=1,\dots,n$ yields
+$n\cdot\frac{1}{1+\lambda_{1}^{2}}\geq\frac{\delta}{s}\sum_{j=1}^{n}\frac{1}{1+\lambda_{j}^{2}},$
+and therefore
+$\frac{1}{1+\lambda_{1}^{2}}\geq\frac{\delta}{ns}\sum_{j=1}^{n}\frac{1}{1+\lambda_{j}^{2}}.$
+Since $\lambda_{1}^{2}\geq\lambda_{j}^{2}$ for all $j$, we have
+$\frac{1}{1+\lambda_{j}^{2}}\geq\frac{1}{1+\lambda_{1}^{2}}\geq\frac{\delta}{ns}\sum_{i=1}^{n}\frac{1}{1+\lambda_{i}^{2}}.$ (20)
+In view of (18) and (20), choosing $\delta_{0}=\frac{\delta}{ns}$ completes the proof. ∎
+By the preceding lemma, we obtain the following key result, noting that the estimate does not depend on $(\theta(t)-(n-2)\frac{\pi}{2})^{-1}$.
 Lemma 3.3.
-Let $\underline{u}$ be the subsolution of the critical LYZ equation (4) and $u^{t}$ be the solution of the LYZ equation (6).
-Then there exist uniform positive constants $t_{0}$ and $\delta_{0}$ such that for any $t\in(0,t_{0})$ and for any $z\in M$, one of the following holds at $z$:
-(i) $\sum\limits_{i,j}F^{i\bar{j}}({\underline{u}_{i\bar{j}}-u^{t}_{i\bar{j}}})\geq\delta_{0}\sum\limits_{i,j}^{n}F^{i\bar{j}}g_{i\bar{j}},\ \text{or}$
-(ii) $\mathcal{F}_{min}\geq\delta_{0}\sum\limits_{i,j}^{n}F^{i\bar{j}}g_{i\bar{j}},$ where $\mathcal{F}_{min}$ is the smallest eigenvalue of the matrix $(F^{i\bar{k}}g_{j\bar{k}})_{n\times n}$.
+Let $\underline{u}$ be a subsolution of the critical LYZ equation (4) and $u^{t}$ the solution of the LYZ equation (6).
+Then there exist uniform positive constants $t_{0}$ and $\delta_{0}$ such that for any $t\in(0,t_{0})$ and any $z\in M$, one of the following holds at $z$:
+  1. (i)
+    $\sum_{i,j}F^{i\bar{j}}(\underline{u}_{i\bar{j}}-u^{t}_{i\bar{j}})\geq\delta_{0}\sum_{i,j}F^{i\bar{j}}g_{i\bar{j}}$, or
+  2. (ii)
+    $\mathcal{F}_{\min}\geq\delta_{0}\sum_{i,j}F^{i\bar{j}}g_{i\bar{j}}$, where $\mathcal{F}_{\min}$ denotes the smallest eigenvalue of the Hermitian matrix $\bigl(F^{i\bar{k}}g_{j\bar{k}}\bigr)_{n\times n}$.
 Proof.
-Let $\mu=(\mu_{1},\cdots,\mu_{n})$ be the eigenvalues of $\chi_{\underline{u}}$ with respect to $\omega$ with $\mu_{1}\geq\cdots\geq\mu_{n}$.
-For any $z\in M$, we choose the normal coordinates such that at $z$
-$g_{i\bar{j}}=\delta_{ij},\ \chi_{t,u^{t}}=\sum\limits_{i=1}^{n}\lambda_{i}\sqrt{-1}dz^{i}\wedge d\bar{z}^{i}\ \text{with}\ \lambda_{1}\geq\cdots\geq\lambda_{n},$
-where $\chi_{t,u^{t}}=\chi+t\omega+\sqrt{-1}\partial\bar{\partial}{u^{t}}$. Then at $z$, we have
-$F^{i\bar{j}}=\frac{\delta_{i\bar{j}}}{1+\lambda_{i}^{2}}.$ (20)
-Let $\delta_{0}$ and $t_{0}$ be the uniform constants in Lemma 3.2. We use Lemma 3.2 to get the proof.
-Case 1: $\sum\limits_{i}F^{i\bar{i}}(\mu_{i}-\lambda_{i})\geq\delta_{0}\sum\limits_{i}F^{i\bar{i}}$.
-Since $F^{1\bar{1}}\leq\cdots\leq F^{n\bar{n}}$ and $\mu_{1}\geq\cdots\geq\mu_{n}$, we use the Schur-Horn theorem [16] to get
-$\sum_{i,j}F^{i\bar{j}}(\underline{u}_{i\bar{j}}-u^{t}_{i\bar{j}})=$ $\sum_{i,j}F^{i\bar{j}}((\chi_{\underline{u}})_{i\bar{j}}-(\chi_{t,u^{t}})_{i\bar{j}})+t\sum_{i}F^{i\bar{i}}$
-$\geq$ $F^{i\bar{j}}((\chi_{\underline{u}})_{i\bar{j}}-(\chi_{t,u^{t}})_{i\bar{j}})$
-$\geq$ $\sum\limits_{i}F^{i\bar{i}}(\mu_{i}-\lambda_{i})$
-$\geq$ $\delta_{0}\sum\limits_{i}F^{i\bar{i}}.$
-Case 2: $\sum\limits_{i}F^{i\bar{i}}(\mu_{i}-\lambda_{i})<\delta_{0}\sum\limits_{i}F^{i\bar{i}}$.
-By Lemma 3.2, we get $\mathcal{F}_{min}=\frac{1}{1+\lambda_{1}^{2}}\geq\delta_{0}\sum\limits_{i}\frac{1}{1+\lambda_{i}^{2}}=\delta_{0}\sum\limits_{i,j}F^{i\bar{j}}g_{i\bar{j}}$.
-∎
+Let $\mu=(\mu_{1},\dots,\mu_{n})$ be the eigenvalues of $\chi_{\underline{u}}$ with respect to $\omega$, arranged so that $\mu_{1}\geq\cdots\geq\mu_{n}$.
+Fix $z\in M$. We choose normal coordinates at $z$ such that
+$g_{i\bar{j}}=\delta_{ij},\quad\chi_{t,u^{t}}=\sum_{i=1}^{n}\lambda_{i}\sqrt{-1}\,dz^{i}\wedge d\bar{z}^{i}\quad\text{with}\quad\lambda_{1}\geq\cdots\geq\lambda_{n},$
+where $\chi_{t,u^{t}}=\chi+t\omega+\sqrt{-1}\partial\bar{\partial}u^{t}$. In these coordinates, at $z$ we have
+$F^{i\bar{j}}=\frac{\delta_{ij}}{1+\lambda_{i}^{2}}.$ (21)
+Let $\delta_{0}$ and $t_{0}$ be the uniform constants given by Lemma 3.2. We now prove the lemma by distinguishing two cases.
+Case 1: $\sum_{i}F^{i\bar{i}}(\mu_{i}-\lambda_{i})\geq\delta_{0}\sum_{i}F^{i\bar{i}}$.
+Since $F^{1\bar{1}}\leq\cdots\leq F^{n\bar{n}}$ and $\mu_{1}\geq\cdots\geq\mu_{n}$, the Schur–Horn theorem [17] implies
+$\sum_{i,j}F^{i\bar{j}}(\underline{u}_{i\bar{j}}-u^{t}_{i\bar{j}})$ $=\sum_{i,j}F^{i\bar{j}}\bigl((\chi_{\underline{u}})_{i\bar{j}}-(\chi_{t,u^{t}})_{i\bar{j}}\bigr)+t\sum_{i}F^{i\bar{i}}$
+$\geq\sum_{i,j}F^{i\bar{j}}\bigl((\chi_{\underline{u}})_{i\bar{j}}-(\chi_{t,u^{t}})_{i\bar{j}}\bigr)$
+$=\sum_{i}F^{i\bar{i}}(\mu_{i}-\lambda_{i})$
+$\geq\delta_{0}\sum_{i}F^{i\bar{i}}.$
+Case 2: $\sum_{i}F^{i\bar{i}}(\mu_{i}-\lambda_{i})<\delta_{0}\sum_{i}F^{i\bar{i}}$.
+By Lemma 3.3, we obtain
+$\mathcal{F}_{\min}=\frac{1}{1+\lambda_{1}^{2}}\geq\delta_{0}\sum_{i}\frac{1}{1+\lambda_{i}^{2}}=\delta_{0}\sum_{i,j}F^{i\bar{j}}g_{i\bar{j}}.$
+This completes the proof. ∎
 3.3. The complex Hessian estimate of Hou-Ma-Wu type
-We use the auxiliary function introduced by Hou-Ma-Wu [17] who established complex Hessian estimates for the $k$-Hessian equation on compact Kähler manifold to derive the complex Hessian estimates for families of LYZ equations.
-Here we use the term $w_{i\bar{j}}\xi^{i}\bar{\xi}^{j}$ instead of the largest eigenvalue of $(g^{j\bar{k}}w_{i\bar{k}})_{n\times n}$ and thus the proof seems much more simple.
-The complex Hessian estimates here do not depend on $({\theta(t)-(n-2)\frac{\pi}{2}})^{-1}$.
+We use the auxiliary function introduced by Hou-Ma-Wu [18], who established complex Hessian estimates for the $k$-Hessian equation on compact Kähler manifolds, to derive the complex Hessian estimates for families of LYZ equations.
+Here, instead of using the largest eigenvalue of $(g^{j\bar{k}}w_{i\bar{k}})_{n\times n}$, we employ the term $w_{i\bar{j}}\xi^{i}\bar{\xi}^{j}$, which simplifies the proof considerably.
 Theorem 3.1.
-Let $\underline{u}$ be the subsolution of the critical LYZ equation (4) and $u^{t}$ be the solution of the LYZ equation (6) with $\sup_{M}u^{t}=0$.
-Then there exists uniform positive constants $t_{0}$ and $C$ such that for any $t\in(0,t_{0})$, we have
-$\sup_{M}{|\sqrt{-1}\partial\bar{\partial}u^{t}|_{\omega}}\leq C(1+\sup_{M}|\nabla u^{t}|_{\omega}^{2}).$ (21)
+Let $\underline{u}$ be a subsolution of the critical LYZ equation (4) and let $u^{t}$ be the solution of the LYZ equation (6) with $\sup_{M}u^{t}=0$.
+Then there exist uniform positive constants $t_{0}$ and $C$, independent of $\bigl(\theta(t)-(n-2)\frac{\pi}{2}\bigr)^{-1}$, such that for any $t\in(0,t_{0})$,
+$\sup_{M}|\sqrt{-1}\partial\bar{\partial}u^{t}|_{\omega}\leq C\bigl(1+\sup_{M}|\nabla u^{t}|_{\omega}^{2}\bigr).$ (22)
 Proof.
@@
 Let $w_{i\bar{j}}=\chi_{i\bar{j}}+tg_{i\bar{j}}+u^{t}_{i\bar{j}}$.
-For simplicity we use $u$ instead of $u^{t}$ in the following proof.
-Denote $S(T^{1,0}M):=\bigcup_{z\in M}\{\xi\in T^{1,0}_{z}M\,|\,|\xi|_{\omega}=1\}$.
-We consider the auxiliary function on $S(T^{1,0}M)$
-$H(z,\xi)=\log(w_{k\bar{l}}\xi^{k}\bar{\xi}^{l})+\varphi(|\nabla u|^{2}_{g})+\psi(u),$
-where $\varphi$ and $\psi$ are given by
-$\begin{split}\varphi\left(s\right)&=-\frac{1}{2}\log\left({1-\frac{s}{{2K}}}\right),\quad 0\leq s\leq K-1,\\ \psi\left(s\right)&=-A\log\left({1+\frac{s}{{2L}}}\right),\quad-L+1\leq s\leq 0,\end{split}$
-with $K=\mathop{\sup}\limits_{M}|\nabla u|_{g}^{2}+1,\ L=\mathop{\sup}\limits_{M}\left|u\right|+1,\ A=2L(C_{0}+1)$ where $C_{0}$ is a uniform positive constant to be determined later. We have
-$(4K)^{-1}$ $\leq\varphi^{\prime}\leq(2K)^{-1},\ \varphi^{\prime\prime}=2\big(\varphi^{\prime})^{2},$ (22)
-$\frac{1}{2}{A}{L^{-1}}$ $\leq-\psi^{\prime}\leq{A}{L^{-1}},\ \psi^{\prime\prime}=A^{-1}(\psi^{\prime})^{2}.$ (23)
-Suppose the function $H$ attains its maximum at $z_{0}$ along the direction $\xi_{0}$. We choose holomorphic coordinates near $z_{0}$ such that
-$g_{i\bar{j}}(z_{0})=\delta_{i\bar{j}},\ \partial_{k}g_{i\bar{j}}(z_{0})=0,\text{and}$ (24)
-$w_{i\bar{j}}(z_{0})=\lambda_{i}\delta_{i\bar{j}}\ \textup{with}\ \lambda_{1}\geq\lambda_{2}\geq\ldots\geq\lambda_{n}$
-which forces $\xi_{0}=\frac{\partial}{\partial z_{1}}$.
-We extend $\xi_{0}$ near $z_{0}$ as $\tilde{\xi}_{0}(z)=(g_{1\bar{1}})^{-\frac{1}{2}}\frac{\partial}{\partial z_{1}}$.
-Then the function $Q(z)=H(z,\tilde{\xi}_{0}(z))$ attains its local maximum at $z_{0}$.
-By the maximum principle, we have at $z_{0}$
+For simplicity, we write $u$ instead of $u^{t}$ in the following proof.
+Denote $S(T^{1,0}M):=\bigcup_{z\in M}\{\xi\in T^{1,0}_{z}M\mid|\xi|_{\omega}=1\}$.
+We consider the auxiliary function on $S(T^{1,0}M)$:
+$H(z,\xi)=\log\bigl(w_{k\bar{l}}\xi^{k}\bar{\xi}^{l}\bigr)+\varphi\bigl(|\nabla u|_{g}^{2}\bigr)+\psi(u),$
+where $\varphi$ and $\psi$ are defined by
+$\varphi(s)$ $=-\frac{1}{2}\log\!\left(1-\frac{s}{2K}\right),\quad 0\leq s\leq K-1,$
+$\psi(s)$ $=-A\log\!\left(1+\frac{s}{2L}\right),\quad-L+1\leq s\leq 0,$
+with
+$K=\sup_{M}|\nabla u|_{g}^{2}+1,\quad L=\sup_{M}|u|+1,\quad A=2L(C_{0}+1),$
+and $C_{0}$ is a uniform positive constant to be determined later. We have the following derivatives:
+$(4K)^{-1}$ $\leq\varphi^{\prime}\leq(2K)^{-1},\quad\varphi^{\prime\prime}=2(\varphi^{\prime})^{2},$ (23)
+$\frac{1}{2}AL^{-1}$ $\leq-\psi^{\prime}\leq AL^{-1},\quad\psi^{\prime\prime}=A^{-1}(\psi^{\prime})^{2}.$ (24)
+Suppose $H$ attains its maximum at $(z_{0},\xi_{0})$. We choose holomorphic coordinates near $z_{0}$ such that
+$g_{i\bar{j}}(z_{0})=\delta_{ij},\quad\partial_{k}g_{i\bar{j}}(z_{0})=0,\quad\text{and}$
+$w_{i\bar{j}}(z_{0})=\lambda_{i}\delta_{ij}\quad\text{with}\quad\lambda_{1}\geq\lambda_{2}\geq\dots\geq\lambda_{n},$
+which forces $\xi_{0}=\frac{\partial}{\partial z_{1}}$. Extend $\xi_{0}$ near $z_{0}$ by $\tilde{\xi}_{0}(z)=(g_{1\bar{1}})^{-\frac{1}{2}}\frac{\partial}{\partial z_{1}}$. Then the function
+$Q(z)=H(z,\tilde{\xi}_{0}(z))$
+attains a local maximum at $z_{0}$.
+By the maximum principle, at $z_{0}$ we have
 $Q_{i}=\frac{w_{1\bar{1},i}}{w_{1\bar{1}}}+\varphi_{i}+\psi_{i}=0,$ (25)
 and
-$0\geq\mathcal{L}(Q)=F^{i\bar{i}}Q_{i\bar{i}}=$ $\lambda_{1}^{-1}F^{i\bar{i}}w_{1\bar{1},i\bar{i}}-\lambda_{1}^{-2}F^{i\bar{i}}{|w_{1\bar{1},i}|^{2}}$
-$+\varphi^{\prime\prime}F^{i\bar{i}}\big|(|\nabla u|^{2})_{i}\big|^{2}+\varphi^{\prime}F^{i\bar{i}}\big(|\nabla u|^{2}\big)_{i\bar{i}}$
-$+\psi^{\prime\prime}F^{i\bar{i}}|u_{i}|^{2}+\psi^{\prime}F^{i\bar{i}}u_{i\bar{i}}.$ (26)
-We will estimate each term in (3.3). In the following, we always assume
+$0\geq\mathcal{L}(Q)=\sum_{i}F^{i\bar{i}}Q_{i\bar{i}}=$ $\sum_{i}\lambda_{1}^{-1}F^{i\bar{i}}w_{1\bar{1},i\bar{i}}-\lambda_{1}^{-2}\sum_{i}F^{i\bar{i}}{|w_{1\bar{1},i}|^{2}}$
+$+\sum_{i}\varphi^{\prime\prime}F^{i\bar{i}}\big|(|\nabla u|^{2})_{i}\big|^{2}+\varphi^{\prime}\sum_{i}F^{i\bar{i}}\big(|\nabla u|^{2}\big)_{i\bar{i}}$
+$+\psi^{\prime\prime}\sum_{i}F^{i\bar{i}}|u_{i}|^{2}+\psi^{\prime}\sum_{i}F^{i\bar{i}}u_{i\bar{i}}.$ (26)
+We will estimate each term in (3.3). In what follows, we always assume
 $\lambda_{1}>80A\sqrt{A}K,$ (27)
-otherwise we finish the proof.
-Since $d\chi=0$, by covariant derivative formulae, we have
+otherwise the desired estimate holds trivially.
+Since $d\chi=0$, the commutation formula for covariant derivatives yields
 $w_{1\bar{1},i\bar{i}}=w_{i\bar{i},1\bar{1}}+(\lambda_{1}-\lambda_{i})R_{1\bar{1}i\bar{i}}.$ (28)
-By (17) and $\lambda_{i}+\lambda_{j}\geq 0$ for $i\neq j$ by Lemma 2.2, we get
-$F^{i\bar{i}}w_{i\bar{i},1\bar{1}}=$ $\sum_{i,j}\frac{\lambda_{i}+\lambda_{j}}{(1+\lambda_{i}^{2})(1+\lambda_{j}^{2})}|w_{i\bar{j},1}|^{2}$
+By (17) and the fact that $\lambda_{i}+\lambda_{j}\geq 0$ for $i\neq j$ (Lemma 2.2), we obtain
+$\sum_{i}F^{i\bar{i}}w_{i\bar{i},1\bar{1}}$ $=\sum_{i,j}\frac{\lambda_{i}+\lambda_{j}}{(1+\lambda_{i}^{2})(1+\lambda_{j}^{2})}|w_{i\bar{j},1}|^{2}$
 $\geq\sum_{i=2}^{n}\frac{\lambda_{1}+\lambda_{i}}{(1+\lambda_{1}^{2})(1+\lambda_{i}^{2})}|w_{1\bar{1},i}|^{2}+\sum_{i=1}^{n}\frac{2\lambda_{i}}{(1+\lambda_{i}^{2})^{2}}|w_{i\bar{i},1}|^{2},$ (29)
-where we used $w_{i\bar{1},1}=w_{1\bar{1},i}$ since $d\chi=0$.
-We prove
-$\sum_{i=1}^{n}\frac{\lambda_{i}}{(1+\lambda_{i}^{2})^{2}}|w_{i\bar{i},1}|^{2}\geq 0.$ (30)
-This is true if $\lambda_{n}\geq 0$. If $\lambda_{n}<0$, by (16), we have
-$\big|\frac{w_{n\bar{n},1}}{1+\lambda_{n}^{2}}\big|=\big|\sum_{i=1}^{n-1}\frac{w_{i\bar{i},1}}{1+\lambda_{i}^{2}}\big|$
-By the above equality and noting that $\lambda_{i}>0$ for $1\leq i\leq n-1$, we get
-$\sum_{i=1}^{n}\frac{\lambda_{i}}{(1+\lambda_{i}^{2})^{2}}|w_{i\bar{i},1}|^{2}=$ $\sum_{i=1}^{n-1}\frac{\lambda_{i}|w_{i\bar{i},1}|^{2}}{(1+\lambda_{i}^{2})^{2}}+\lambda_{n}\big|\sum_{i=1}^{n-1}\frac{w_{i\bar{i},1}}{1+\lambda_{i}^{2}}\big|^{2}$
-$\geq$ $\big(\sum_{i=1}^{n-1}\lambda_{i}^{-1}\big)^{-1}\big|\sum_{i=1}^{n-1}{\frac{w_{i\bar{i},1}}{1+\lambda_{i}^{2}}}\big|^{2}+\lambda_{n}\big|\sum_{i=1}^{n-1}{\frac{w_{i\bar{i},1}}{1+\lambda_{i}^{2}}}\big|^{2}$
-$=$ $\frac{1+\lambda_{n}\sum\limits_{i=1}^{n-1}\lambda_{i}^{-1}}{\sum\limits_{i=1}^{n-1}\lambda_{i}^{-1}}\big|\sum_{i=1}^{n-1}{\frac{w_{i\bar{i},1}}{1+\lambda_{i}^{2}}}\big|^{2}$
-$\geq$ $0,$
-where in the last inequality we used
+where we have used $w_{i\bar{1},1}=w_{1\bar{1},i}$, which follows from $d\chi=0$.
+We now prove the following inequality
+$\sum_{i}\frac{\lambda_{i}}{(1+\lambda_{i}^{2})^{2}}|w_{i\bar{i},1}|^{2}\geq 0.$ (30)
+This is clearly true if $\lambda_{n}\geq 0$. Assume now that $\lambda_{n}<0$. By (16), we have
+$\Bigl|\frac{w_{n\bar{n},1}}{1+\lambda_{n}^{2}}\Bigr|=\Bigl|\sum_{i=1}^{n-1}\frac{w_{i\bar{i},1}}{1+\lambda_{i}^{2}}\Bigr|.$
+Using this relation and the fact that $\lambda_{i}>0$ for $1\leq i\leq n-1$, we obtain
+$\sum_{i=1}^{n}\frac{\lambda_{i}}{(1+\lambda_{i}^{2})^{2}}|w_{i\bar{i},1}|^{2}$ $=\sum_{i=1}^{n-1}\frac{\lambda_{i}|w_{i\bar{i},1}|^{2}}{(1+\lambda_{i}^{2})^{2}}+\lambda_{n}\Bigl|\sum_{i=1}^{n-1}\frac{w_{i\bar{i},1}}{1+\lambda_{i}^{2}}\Bigr|^{2}$
+$\geq\Bigl(\sum_{i=1}^{n-1}\lambda_{i}^{-1}\Bigr)^{-1}\Bigl|\sum_{i=1}^{n-1}\frac{w_{i\bar{i},1}}{1+\lambda_{i}^{2}}\Bigr|^{2}+\lambda_{n}\Bigl|\sum_{i=1}^{n-1}\frac{w_{i\bar{i},1}}{1+\lambda_{i}^{2}}\Bigr|^{2}$
+$=\frac{1+\lambda_{n}\sum\limits_{i=1}^{n-1}\lambda_{i}^{-1}}{\sum\limits_{i=1}^{n-1}\lambda_{i}^{-1}}\Bigl|\sum_{i=1}^{n-1}\frac{w_{i\bar{i},1}}{1+\lambda_{i}^{2}}\Bigr|^{2}$
+$\geq 0,$
+where the last inequality follows from
 $1+\lambda_{n}\sum_{i=1}^{n-1}\lambda_{i}^{-1}=\lambda_{n}\,\frac{\sigma_{n-1}(\lambda)}{\sigma_{n}(\lambda)}\geq 0.$
-Combining (3.3) with (30), we have
-$F^{i\bar{i}}w_{i\bar{i},1\bar{1}}\geq\sum_{i=2}^{n}\frac{\lambda_{1}+\lambda_{i}}{(1+\lambda_{1}^{2})(1+\lambda_{i}^{2})}|w_{1\bar{1},i}|^{2}.$ (31)
-Then we get
-$\lambda_{1}^{-1}F^{i\bar{i}}w_{i\bar{i},1\bar{1}}-\lambda_{1}^{-2}F^{i\bar{i}}{|w_{1\bar{1},i}|^{2}}\geq$ $-\lambda_{1}^{-2}{(1+\lambda_{1}^{2})^{-1}}\big|w_{1\bar{1},1}|^{2}$
-$-\sum_{i=2}^{n}\frac{|w_{1\bar{1},i}|^{2}}{\lambda_{1}^{2}(1+\lambda_{1}^{2})(1+\lambda_{i}^{2})}$
+Combining (3.3) with (30), we obtain
+$\sum_{i}F^{i\bar{i}}w_{i\bar{i},1\bar{1}}\geq\sum_{i=2}^{n}\frac{\lambda_{1}+\lambda_{i}}{(1+\lambda_{1}^{2})(1+\lambda_{i}^{2})}|w_{1\bar{1},i}|^{2}.$ (31)
+Consequently,
+$\lambda_{1}^{-1}\sum_{i}F^{i\bar{i}}w_{i\bar{i},1\bar{1}}-\lambda_{1}^{-2}\sum_{i}F^{i\bar{i}}|w_{1\bar{1},i}|^{2}$
+$\geq$ $-\lambda_{1}^{-2}(1+\lambda_{1}^{2})^{-1}|w_{1\bar{1},1}|^{2}-\sum_{i=2}^{n}\frac{|w_{1\bar{1},i}|^{2}}{\lambda_{1}^{2}(1+\lambda_{1}^{2})(1+\lambda_{i}^{2})}$
 $+\sum_{i=2}^{n}\frac{\lambda_{i}}{\lambda_{1}(1+\lambda_{1}^{2})(1+\lambda_{i}^{2})}|w_{1\bar{1},i}|^{2}$
 $=$ $\mathcal{I}+\mathcal{II}+\mathcal{III}.$ (32)
-We first estimate $\mathcal{I}$ and $\mathcal{II}$ by (25):
+We first estimate $\mathcal{I}$ and $\mathcal{II}$ using (25):
 $\mathcal{I}$ $\geq-2(\varphi^{\prime})^{2}F^{1\bar{1}}|(|\nabla u|^{2})_{1}|^{2}-2(\psi^{\prime})^{2}F^{1\bar{1}}|u_{1}|^{2}$
-$\geq$ $-\varphi^{\prime\prime}F^{1\bar{1}}|(|\nabla u|^{2})_{1}|^{2}-2(\psi^{\prime})^{2}KF^{1\bar{1}}$ (33)
+$\geq-\varphi^{\prime\prime}F^{1\bar{1}}|(|\nabla u|^{2})_{1}|^{2}-2(\psi^{\prime})^{2}KF^{1\bar{1}},$ (33)
 and
 $\mathcal{II}$ $=-(1+\lambda_{1}^{2})^{-1}\sum_{i=2}^{n}\frac{F^{i\bar{i}}|w_{1\bar{1},i}|^{2}}{\lambda_{1}^{2}}$
-$\geq$ $-2(1+\lambda_{1}^{2})^{-1}(\varphi^{\prime})^{2}\sum_{i=2}^{n}F^{i\bar{i}}|(|\nabla u|^{2})_{i}|^{2}-2(1+\lambda_{1}^{2})^{-1}(\psi^{\prime})^{2}\sum_{i=2}^{n}F^{i\bar{i}}|u_{i}|^{2}$
+$\geq-2(1+\lambda_{1}^{2})^{-1}(\varphi^{\prime})^{2}\sum_{i=2}^{n}F^{i\bar{i}}|(|\nabla u|^{2})_{i}|^{2}$
+$\quad-2(1+\lambda_{1}^{2})^{-1}(\psi^{\prime})^{2}\sum_{i=2}^{n}F^{i\bar{i}}|u_{i}|^{2}$
 $\geq-\frac{1}{12}\varphi^{\prime\prime}\sum_{i=2}^{n}F^{i\bar{i}}|(|\nabla u|^{2})_{i}|^{2}-\frac{1}{2}\psi^{\prime\prime}\sum_{i=2}^{n}F^{i\bar{i}}|u_{i}|^{2},$ (34)
-where in the last inequality we used (27).
-Next we estimate term $\mathcal{III}$ . Since $\lambda_{i}>0$ for $1\leq i\leq n-1$ by Lemma 2.2, we get
-$\mathcal{III}\geq$ $\frac{\lambda_{n}|w_{1\bar{1},n}|^{2}}{\lambda_{1}(1+\lambda_{1}^{2})(1+\lambda_{n}^{2})},$
-then by (25) and the Cauchy-Schwarz inequality, we obtain
+where the last inequality follows from (27).
+Next we estimate the term $\mathcal{III}$. Since $\lambda_{i}>0$ for $1\leq i\leq n-1$ by Lemma 2.2, we obtain
+$\mathcal{III}\geq\frac{\lambda_{n}|w_{1\bar{1},n}|^{2}}{\lambda_{1}(1+\lambda_{1}^{2})(1+\lambda_{n}^{2})}.$
+Then, using (25) and the Cauchy-Schwarz inequality, we have
 $\mathcal{III}$ $\geq-\frac{4}{3}\frac{\lambda_{1}|\lambda_{n}|}{1+\lambda_{1}^{2}}(\varphi^{\prime})^{2}F^{n\bar{n}}|(|\nabla u|^{2})_{n}|^{2}-\frac{4\lambda_{1}|\lambda_{n}|}{1+\lambda_{1}^{2}}(\psi^{\prime})^{2}F^{n\bar{n}}|u_{n}|^{2}$
 $\geq-\frac{2}{3}\varphi^{\prime\prime}F^{n\bar{n}}|(|\nabla u|^{2})_{n}|^{2}-\frac{4\lambda_{1}|\lambda_{n}|}{1+\lambda_{1}^{2}}(\psi^{\prime})^{2}F^{n\bar{n}}|u_{n}|^{2},$ (35)
-where we used (22). Let $\tau=\frac{1}{8A}$. We continue to estimate $\mathcal{III}$ by two cases.
-Case 1: $|\lambda_{n}|\leq\tau\lambda_{1}$. In this case, we have
+where we used (23).
+Let $\tau=\frac{1}{8A}$. We proceed to estimate $\mathcal{III}$ in two cases.
+Case 1: $|\lambda_{n}|\leq\tau\lambda_{1}$. In this case, we obtain
 $\mathcal{III}$ $\geq-\frac{2}{3}\varphi^{\prime\prime}F^{n\bar{n}}|(|\nabla u|^{2})_{n}|^{2}-4\tau(\psi^{\prime})^{2}F^{n\bar{n}}|u_{n}|^{2}$
@@
 where we used (24).
-Case 2: $|\lambda_{n}|\geq\tau\lambda_{1}$. By (27), we have $|u_{n\bar{n}}|^{2}=(\lambda_{n}-\chi_{n\bar{n}}-t)^{2}\geq\frac{\tau^{2}}{2}\lambda_{1}^{2}$ and we obtain
-$4(\psi^{\prime})^{2}\frac{\lambda_{1}|\lambda_{n}|}{1+\lambda_{1}^{2}}F^{n\bar{n}}|u_{n}|^{2}\leq$ $4AF^{n\bar{n}}K\leq\frac{1}{3}{\varphi^{\prime}}F^{n\bar{n}}u_{n\bar{n}}^{2}.$
-Inserting the above inequality into (3.3), we have
+Case 2: $|\lambda_{n}|\geq\tau\lambda_{1}$. By (27), we have $|u_{n\bar{n}}|^{2}=(\lambda_{n}-\chi_{n\bar{n}}-t)^{2}\geq\frac{\tau^{2}}{2}\lambda_{1}^{2}$, and consequently
+$4(\psi^{\prime})^{2}\frac{\lambda_{1}|\lambda_{n}|}{1+\lambda_{1}^{2}}F^{n\bar{n}}|u_{n}|^{2}\leq 4AF^{n\bar{n}}K\leq\frac{1}{3}\varphi^{\prime}F^{n\bar{n}}u_{n\bar{n}}^{2}.$
+Inserting this inequality into (3.3) yields
 $\mathcal{III}\geq-\frac{2}{3}\varphi^{\prime\prime}F^{n\bar{n}}|(|\nabla u|^{2})_{n}|^{2}-\frac{1}{3}\varphi^{\prime}F^{n\bar{n}}u_{n\bar{n}}^{2}.$ (37)
-By (3.3) and (37), we obtain
+Combining (3.3) and (37), we obtain
 $\mathcal{III}\geq-\frac{2}{3}\varphi^{\prime\prime}F^{n\bar{n}}|(|\nabla u|^{2})_{n}|^{2}-\frac{1}{2}\psi^{\prime\prime}F^{n\bar{n}}|u_{n}|^{2}-\frac{1}{3}\varphi^{\prime}F^{n\bar{n}}u_{n\bar{n}}^{2}.$ (38)
 Inserting (3.3), (3.3) and (38) into (32), we obtain
-$\lambda_{1}^{-1}F^{i\bar{i}}w_{i\bar{i},1\bar{1}}-\lambda_{1}^{-2}F^{i\bar{i}}{|w_{1\bar{1},i}|^{2}}+\frac{3}{4}\varphi^{\prime\prime}F^{i\bar{i}}|(|\nabla u|^{2})_{i}|^{2}+\psi^{\prime\prime}F^{i\bar{i}}|u_{i}|^{2}+\frac{1}{3}\varphi^{\prime}F^{i\bar{i}}u_{i\bar{i}}^{2}$
+$\lambda_{1}^{-1}\sum_{i}F^{i\bar{i}}w_{i\bar{i},1\bar{1}}-\lambda_{1}^{-2}\sum_{i}F^{i\bar{i}}|w_{1\bar{1},i}|^{2}$ (39)
+$+\frac{3}{4}\varphi^{\prime\prime}\sum_{i}F^{i\bar{i}}|(|\nabla u|^{2})_{i}|^{2}+\psi^{\prime\prime}\sum_{i}F^{i\bar{i}}|u_{i}|^{2}+\frac{1}{3}\varphi^{\prime}\sum_{i}F^{i\bar{i}}u_{i\bar{i}}^{2}$
 $\geq$ $\frac{1}{3}\varphi^{\prime}F^{1\bar{1}}u_{1\bar{1}}^{2}-2K(\psi^{\prime})^{2}F^{1\bar{1}}>0,$ (40)
-where we used $u_{1\bar{1}}\geq\frac{\lambda_{1}}{2}$ which follows from (27).
-Next we estimate $F^{i\bar{i}}(|\nabla u|^{2})_{i\bar{i}}$. By direct computation, we have
-$F^{i\bar{i}}(|\nabla u|^{2})_{i\bar{i}}=F^{i\bar{i}}(u_{k\bar{i}}u_{\bar{k}i}+u_{ki}u_{\bar{k}\bar{i}})+F^{i\bar{i}}(u_{ki\bar{i}}u_{\bar{k}}+u_{k}u_{\bar{k}i\bar{i}}).$ (40)
-By covariant derivative formulae and (16), we have
-$F^{i\bar{i}}(u_{ki\bar{i}}u_{\bar{k}}+u_{k}u_{\bar{k}i\bar{i}})$
-$\geq$ $-F^{i\bar{i}}(u_{i\bar{i}k}u_{\bar{k}}+u_{k}u_{i\bar{i}\bar{k}})-|Rm||\nabla u|^{2}\sum_{i}F^{i\bar{i}}$
-$\geq$ $-2K(|\nabla\chi|+|Rm|)\sum_{i}F^{i\bar{i}}.$ (41)
+where we used $u_{1\bar{1}}\geq\frac{\lambda_{1}}{2}$, which follows from (27).
+Next we estimate $\sum\limits_{i}F^{i\bar{i}}(|\nabla u|^{2})_{i\bar{i}}$. By direct computation, we have
+$\sum_{i}F^{i\bar{i}}(|\nabla u|^{2})_{i\bar{i}}=\sum_{i}F^{i\bar{i}}(u_{k\bar{i}}u_{\bar{k}i}+u_{ki}u_{\bar{k}\bar{i}})+\sum_{i,k}F^{i\bar{i}}(u_{ki\bar{i}}u_{\bar{k}}+u_{k}u_{\bar{k}i\bar{i}}).$ (41)
+Using the covariant derivative formulae and (16), we obtain
+$\sum_{i,k}F^{i\bar{i}}(u_{ki\bar{i}}u_{\bar{k}}+u_{k}u_{\bar{k}i\bar{i}})$
+$\geq$ $-\sum_{i,k}F^{i\bar{i}}(u_{i\bar{i}k}u_{\bar{k}}+u_{k}u_{i\bar{i}\bar{k}})-|\text{Rm}||\nabla u|^{2}\sum_{i}F^{i\bar{i}}$
+$\geq$ $-2K(|\nabla\chi|+|\text{Rm}|)\sum_{i}F^{i\bar{i}}.$ (42)
 Inserting the above estimate into (41), we get
-$F^{i\bar{i}}(|\nabla u|^{2})_{i\bar{i}}\geq F^{i\bar{i}}u_{i\bar{i}}^{2}-2K(|\nabla\chi|+|Rm|)\sum_{i=1}^{n}F^{i\bar{i}}.$ (42)
-Inserting (28), (3.3) and (42) into (3.3), we obtain
-$0\geq F^{i\bar{i}}Q_{i\bar{i}}\geq$ $\psi^{\prime}\sum_{i=1}^{n}F^{i\bar{i}}u_{i\bar{i}}+\frac{1}{3}\varphi^{\prime}F^{i\bar{i}}u_{i\bar{i}}^{2}-3K\varphi^{\prime}(|\nabla\chi|+|Rm|)\sum_{i}F^{i\bar{i}}$
-$\geq$ $\psi^{\prime}\sum_{i=1}^{n}F^{i\bar{i}}u_{i\bar{i}}+\frac{1}{3}\varphi^{\prime}F^{i\bar{i}}u_{i\bar{i}}^{2}-2(|\nabla\chi|+|Rm|)\sum_{i}F^{i\bar{i}},$ (43)
-By Lemma 3.3 and noting that we have assumed $\underline{u}=0$, one of the following holds
-$(i)\ -\sum_{i=1}^{n}F^{i\bar{i}}u_{i\bar{i}}\geq\delta_{0}\sum_{i=1}^{n}F^{i\bar{i}},\ \text{or}$
-$(ii)\ F^{1\bar{1}}\geq\delta_{0}\sum_{i=1}^{n}F^{i\bar{i}}.$
-We choose $C_{0}=6\delta_{0}^{-1}(\max|\nabla\chi|+|Rm|)$.
-If $(i)$ holds, by (3.3) and $\psi^{\prime}\geq C_{0}+1$, we get a contradiction.
-Then $(ii)$ holds and we obtain $|\lambda_{n}|\geq\frac{\delta_{0}}{2}\lambda_{1}$.
-Hence we have
-$0\geq$ $\frac{\lambda_{1}^{2}}{4}F^{n\bar{n}}-3K(|\nabla\chi|+|Rm|)\sum_{i}F^{i\bar{i}}$
-$\geq$ $\big(\frac{\lambda_{1}^{2}}{4}-3nK(|\nabla\chi|+|Rm|)\big)F^{n\bar{n}}.$
-Therefore we get $\lambda_{1}\leq CK$ and $C$ is a uniform positive constant which is independent of $t$. ∎
-4. Gradient estimates and the proof of the main theorem
+$\sum_{i}F^{i\bar{i}}(|\nabla u|^{2})_{i\bar{i}}\geq\sum_{i}F^{i\bar{i}}u_{i\bar{i}}^{2}-2K(|\nabla\chi|+|\text{Rm}|)\sum_{i}F^{i\bar{i}}.$ (43)
+Inserting (28), (39) and (43) into (3.3), we obtain
+$0\geq\sum_{i}F^{i\bar{i}}Q_{i\bar{i}}$ $\geq\psi^{\prime}\sum_{i}F^{i\bar{i}}u_{i\bar{i}}+\frac{1}{3}\varphi^{\prime}\sum_{i}F^{i\bar{i}}u_{i\bar{i}}^{2}-3K\varphi^{\prime}(|\nabla\chi|+|\text{Rm}|)\sum_{i}F^{i\bar{i}}$
+$\geq\psi^{\prime}\sum_{i=1}^{n}F^{i\bar{i}}u_{i\bar{i}}+\frac{1}{3}\varphi^{\prime}\sum_{i}F^{i\bar{i}}u_{i\bar{i}}^{2}-2(|\nabla\chi|+|\text{Rm}|)\sum_{i}F^{i\bar{i}}.$ (44)
+By Lemma 3.3 and noting that we have assumed $\underline{u}=0$, one of the following holds:
+$(i)\ -\sum_{i}F^{i\bar{i}}u_{i\bar{i}}\geq\delta_{0}\sum_{i}F^{i\bar{i}},\quad\text{or}$
+$(ii)\ F^{1\bar{1}}\geq\delta_{0}\sum_{i}F^{i\bar{i}}.$
+We choose $C_{0}=6\delta_{0}^{-1}(\max|\nabla\chi|+|\text{Rm}|)$.
+If (i) holds, then by (3.3) and $\psi^{\prime}\geq C_{0}+1$, we obtain a contradiction. Thus (ii) holds, and we get $|\lambda_{n}|\geq\frac{\delta_{0}}{2}\lambda_{1}$. Consequently, we have
+$0$ $\geq\frac{\lambda_{1}^{2}}{4}F^{n\bar{n}}-3K(|\nabla\chi|+|\text{Rm}|)\sum_{i}F^{i\bar{i}}$
+$\geq\Bigl(\frac{\lambda_{1}^{2}}{4}-3nK(|\nabla\chi|+|\text{Rm}|)\Bigr)F^{n\bar{n}}.$
+Therefore, $\lambda_{1}\leq CK$, where $C$ is a uniform positive constant independent of $t$. ∎
+4. Gradient estimates and the proof of the main theorems
 In this section, following the argument by Dinew-Kolodziej [10] and Szekelyhidi [38], we prove a Liouville theorem for the equation $\sigma_{n}+\sigma_{n-1}=0$.
@@
 For any $C^{2}$ function $\varphi$ and any $z_{0}\in\Omega^{n}$ of the local minimum point of $\varphi-v$, we need to show
-$\sigma_{n-1}(\varphi_{\alpha\bar{\beta}}(z_{0}))+a_{0}\sigma_{n}(\varphi_{\alpha\bar{\beta}}(z_{0}))\geq 0.$ (45)
+$\sigma_{n-1}(\varphi_{\alpha\bar{\beta}}(z_{0}))+\sigma_{n}(\varphi_{\alpha\bar{\beta}}(z_{0}))\geq 0.$ (46)
 For any $\epsilon>0$, for all sufficiently large $i$, we have $\sup_{B_{1}(z_{0})}|v^{i}-v|<\frac{\epsilon^{3}}{3}$.
@@
 Since $v^{i}$ is a viscosity subsolution, we have
-$\sigma_{n-1}(\varphi_{\alpha\bar{\beta}}(z_{i,\epsilon})+\epsilon\delta_{\alpha\beta})+a_{0}\sigma_{n}(\varphi_{\alpha\bar{\beta}}(z_{i,\epsilon})+\epsilon\delta_{\alpha\beta})\geq 0.$
+$\sigma_{n-1}(\varphi_{\alpha\bar{\beta}}(z_{i,\epsilon})+\epsilon\delta_{\alpha\beta})+\sigma_{n}(\varphi_{\alpha\bar{\beta}}(z_{i,\epsilon})+\epsilon\delta_{\alpha\beta})\geq 0.$
 Letting $\epsilon\to 0$ then implies (46). This proves $(i)$.
@@
 Now, since $w$ is (n1)(n-1)-admissible, the function $w+\epsilon|z|^{2}$ is also (n1)(n-1)-admissible.
-As $v$ is a viscosity solution (and hence a supersolution), and $w+\epsilon|z|^{2}$ touches $v$ from above at $z_{0}$, the supersolution property yields:
-$\sigma_{n-1}((w_{\alpha\bar{\beta}}(z_{0}))+\epsilon I_{n})+a_{0}\sigma_{n}((w_{\alpha\bar{\beta}}(z_{0}))+\epsilon I_{n})\leq 0.$ (46)
+As $v$ is a viscosity solution (and hence a supersolution), by the definition of the supersolution, we have
+$\sigma_{n-1}((w_{\alpha\bar{\beta}}(z_{0}))+\epsilon I_{n})+\sigma_{n}((w_{\alpha\bar{\beta}}(z_{0}))+\epsilon I_{n})\leq 0.$ (47)
 However, since $\sigma_{k}(w_{\alpha\bar{\beta}}(z_{0}))\geq 0$ for $1\leq k\leq n-1$, expanding the left-hand side of (47) gives:
-$\sigma_{n-1}(w_{\alpha\bar{\beta}}(z_{0}))+a_{0}\sigma_{n}(w_{\alpha\bar{\beta}}(z_{0}))\leq-\epsilon^{n-1}<0.$ (47)
+$\sigma_{n-1}(w_{\alpha\bar{\beta}}(z_{0}))+\sigma_{n}(w_{\alpha\bar{\beta}}(z_{0}))\leq-\epsilon^{n}<0.$ (48)
 This contradicts the assumption that $w$ is a viscosity subsolution of (45), thereby completing the proof. ∎
@@
 $[w]_{r}(z)=\int_{\mathbb{C}^{n}}w(z+ry)\eta(y)\,\mathcal{E}^{n}(y).$ (50)
-For the function $\tilde{w}(\tilde{z})=w(z)+|z_{n+1}|^{2}$ defined on $\mathbb{C}^{n+1}$, where $\tilde{z}=(z,z_{n+1})$, we define its mollification as
+For the function $\tilde{w}(\tilde{z})=w(z)+|z_{n+1}|^{2}$ defined on $\mathbb{C}^{n+1}$, where $\tilde{z}=(z,z_{n+1})$, we define
 […]
@@
   4. (iv)
-    If $w$ is another subsolution, then $\frac{1}{2}(v+w)$ is a weak subsolution.
+    If $w$ is a subsolution to the same equation, then $\frac{1}{2}(v+w)$ is also a subsolution.
 Proof.
@@
 Since $\tilde{v}$ is a viscosity subsolution to $\sigma_{n}=0$, we have
-$\sigma_{n}(\sqrt{-1}\partial\bar{\partial}\tilde{\varphi})(0,z_{0})\geq 0.$
+$\sigma_{n}(\sqrt{-1}\partial\bar{\partial}\tilde{\varphi})(z_{0},0)\geq 0.$
 Observing that
@@
 Case 1: There exist a positive constant $\rho$, sequences $\epsilon_{k}\to 0$, $x_{k}\in\mathbb{C}^{n}$, $r_{k}\to\infty$, and a sequence of unit $(1,0)$-vectors $\xi_{k}$ such that
-$\frac{1}{2}\bigl(\bigl[\frac{(v^{\epsilon})^{2}}{3}\bigr]_{r}(z)+[v]_{\rho}(z)-\frac{7}{6}\bigr)$ $\geq v(x_{k}),$ (52)
+$\frac{1}{2}\bigl(\bigl[\frac{(v^{\epsilon})^{2}}{3}\bigr]_{r}(x_{k})+[v]_{\rho}(x_{k})-\frac{7}{6}\bigr)$ $\geq v(x_{k}),$ (53)
 $\lim_{k\to\infty}\int_{B_{r_{k}}(x_{k})}|v_{\xi_{k}}^{\epsilon}|^{2}(z)\,\mathcal{E}^{n}(z)$ $=0.$ (54)
@@
 $a_{0}=-\liminf_{i\to\infty}\tan\hat{\theta}(t_{i})M_{i}^{2}.$
-Recall that $\hat{\theta}(t)\in(\pi-Ct,\pi)$, so $a_{0}\in[0,\infty]$. By passing to a further subsequence, we may assume
+Recall that $\hat{\theta}(t)\in(\pi-2Ct,\pi-Ct)$, so $a_{0}\in[0,\infty]$. By passing to a further subsequence, we may assume
 $a_{0}=-\lim_{i\to\infty}\tan\hat{\theta}(t_{i})M_{i}^{2}.$ (61)
@@
 $-3\tan\hat{\theta}(t)(\chi+t\omega+\sqrt{-1}\partial\bar{\partial}u^{t})\wedge\omega^{2}+\omega^{3}.$
-As a corollary of our theorem, we solve the 2-Hessian equation under weaker condition instead of the usual condition $\chi\in\Gamma_{2}(M)$.
+As a corollary of our theorem, we solve the 2-Hessian equation under weaker conditions instead of the usual condition $\chi\in\Gamma_{2}(M)$.
 Corollary 5.1.
-Let $(M,\omega)$ be a compact Kähler manifold with dimension $n=3$ and $\chi$ be a real $(1,1)$ form satisfying the following conditions:
+Let $(M,\omega)$ be a compact Kähler manifold with dimension $n=3$ and $\chi$ be a real $(1,1)$-form satisfying the following conditions:
   1. (i)
-    $3\int_{M}\chi^{2}\wedge\omega=\int_{M}\omega^{3}$, $\int_{M}\chi^{3}<3\int_{M}\chi\wedge\omega^{2}$, and
+    $3\int_{M}\chi^{2}\wedge\omega=\int_{M}\omega^{3}$, $\int_{M}\chi^{3}<3\int_{M}\chi\wedge\omega^{2}$;
   2. (ii)
     the subsolution condition
-    $\chi\wedge\omega>0\ \ \text{as a $(2,2)$ form}.$ (67)
+    $\chi\wedge\omega>0\ \ \text{as a $(2,2)$-form}.$ (68)
 Then there exists a unique smooth solution $u$ solving
@@
 Proof.
-Condition $(i)$ implies the principal argument of $\int_{M}(\chi+\sqrt{-1}\omega)^{3}$ is $\pi$.
-Condition $(ii)$ means that 0 is the subsolution of the critical LYZ equation in dimension 3.
-According to Theorem 1.1, there exists a smooth solution solving the critical LYZ equation in dimension 3 which is exactly the 2-Hessian equation $\sigma_{2}=1$.∎
+Condition $(i)$ gives that the principal argument of $\int_{M}(\chi+\sqrt{-1}\omega)^{3}$ is $\pi$.
+Condition $(ii)$ means that 0 is a subsolution of the critical LYZ equation in dimension $3$.
+According to Theorem 1.1, there exists a smooth solution to the critical LYZ equation in dimension $3$, which is exactly the $2$-Hessian equation $\sigma_{2}(\chi_{u})=1$.
+∎
 Remark 5.1.
 If $\chi\in\Gamma_{2}(M)$, then it satisfies the subsolution condition $(ii)$.
-But the subsolution condition is weaker than $\chi\in\Gamma_{2}(M)$.
-We will prove the integral condition $(i)$ is necessary.
+However, the subsolution condition is strictly weaker than $\chi\in\Gamma_{2}(M)$.
+We will prove that the integral condition $(i)$ is necessary.
 Let $\lambda=(\lambda_{1},\lambda_{2},\lambda_{3})$ be the eigenvalues of $\chi_{u}$ with respect to $\omega$ such that $\lambda_{1}\geq\lambda_{2}\geq\lambda_{3}$.
@@
 $\sigma_{3}(\chi_{u})\leq\frac{1}{9}\sigma_{1}(\chi_{u})\sigma_{2}(\chi_{u})=\frac{1}{9}\sigma_{1}(\chi_{u}).$ (71)
-Integration by parts, we get
+Integrating by parts, we obtain
 $\int_{M}\chi^{3}$ $=\int_{M}\chi_{u}^{3}\leq\frac{1}{3}\int_{M}{\chi_{u}\wedge\omega^{2}}=\frac{1}{3}\int_{M}{\chi\wedge\omega^{2}}$
@@
 When $n=4$, the critical LYZ equation is the Hessian quotient equation.
-In this case the subsolution implies $\sigma_{2}(\mu|i)>1$ for $1\leq i\leq 4$, where $\mu=(\mu_{1},\mu_{2},\mu_{3},\mu_{4})$ are eigenvalues of $\chi$ with respect to $\omega$.
-Assuming further $\chi\in\Gamma_{3}(M)$, Sun [34] and Székelyhidi [36] proved the existence result using different continuity methods.
-As a corollary of our result, we prove an existence result for the Hessian quotient equation $\sigma_{3}(\chi_{u})=\sigma_{1}(\chi_{u})$ under weaker conditions than before.
+In this case, the subsolution condition yields $3\chi^{2}\wedge\omega-\omega^{3}>0$ as a $(3,3)$-form.
+Under the additional assumption that $\chi\in\Gamma_{3}(M)$, Sun [36] and Székelyhidi [38] proved the existence result using different continuity methods.
+As a corollary of our result, we prove an existence result for the Hessian quotient equation $\sigma_{3}(\chi_{u})=\sigma_{1}(\chi_{u})$ under conditions weaker than those previously required.
 Corollary 5.2.
-Let $(M,\omega)$ be a compact Kähler manifold with dimension $n=4$. Let $\chi$ be a real $(1,1)$-form satisfying the following conditions:
+Let $(M,\omega)$ be a compact Kähler manifold of dimension $n=4$. Let $\chi$ be a real $(1,1)$-form satisfying the following conditions:
   1. (i)
-    $\int_{M}\chi^{3}\wedge\omega=\int_{M}\chi\wedge\omega^{3},\int_{M}\chi\wedge\omega^{3}>0,\ \mathrm{Re}\int_{M}(\chi+\sqrt{-1}\omega)^{4}<0,$
+    $\int_{M}\chi^{3}\wedge\omega=\int_{M}\chi\wedge\omega^{3}$, $\int_{M}\chi\wedge\omega^{3}>0$, and $\mathrm{Re}\int_{M}(\chi+\sqrt{-1}\omega)^{4}<0$;
   2. (ii)
-    $3\chi^{2}\wedge\omega-\omega^{3}>0\ \text{as a $(3,3)$-form}.$
+    $3\chi^{2}\wedge\omega-\omega^{3}>0$ as a $(3,3)$-form.
 Then there exists a unique smooth function $u$ solving
 $\sigma_{3}(\chi_{u})=\sigma_{1}(\chi_{u}),$ (72)
-$\chi_{u}\in\Gamma_{3}(M),\sup_{M}u=0.$
+$\chi_{u}\in\Gamma_{3}(M),\quad\sup_{M}u=0.$
 Proof.
-Condition $(i)$ implies the principal argument of $\int_{M}(\chi+\sqrt{-1}\omega)^{4}$ is $\pi$.
-To apply Theorem 1.1, we only need to show 0 is the subsolution of the critical LYZ equation.
-By condition $(ii)$, we have $\sigma_{2}(\chi)>2$ and then $\sigma_{1}^{2}(\chi)\geq\frac{8}{3}\sigma_{2}(\chi)>\frac{16}{3}$.
-Since there exists $z_{0}\in M$ such that $\sigma_{1}(\chi(z_{0}))=4\frac{\int_{M}\chi\wedge\omega^{3}}{\int_{M}\omega^{4}}>0$, we get $\sigma_{1}(\chi)>0$ on $M$.
-Let $\mu=(\mu_{1},\mu_{2},\mu_{3},\mu_{4})$ be the eigenvalues of $\chi$ with respect to $\omega$ such that $\mu_{1}\geq\mu_{2}\geq\mu_{3}\geq\mu_{4}$. We need to prove
+Condition $(i)$ implies that the principal argument of $\int_{M}(\chi+\sqrt{-1}\omega)^{4}$ is $\pi$. To apply Theorem 1.1, we only need to show that 0 is a subsolution of the critical LYZ equation.
+By condition $(ii)$, we have $\sigma_{2}(\chi)>2$, which implies $\sigma_{1}^{2}(\chi)\geq\frac{8}{3}\sigma_{2}(\chi)>\frac{16}{3}$.
+Moreover, since there exists a point $z_{0}\in M$ such that $\sigma_{1}(\chi(z_{0}))=4\frac{\int_{M}\chi\wedge\omega^{3}}{\int_{M}\omega^{4}}>0$, we conclude that $\sigma_{1}(\chi)>0$ on $M$.
+Let $\mu=(\mu_{1},\mu_{2},\mu_{3},\mu_{4})$ be the eigenvalues of $\chi$ with respect to $\omega$, ordered such that $\mu_{1}\geq\mu_{2}\geq\mu_{3}\geq\mu_{4}$. We need to prove
 $\arctan\mu_{2}+\arctan\mu_{3}+\arctan\mu_{4}>\frac{\pi}{2}.$ (73)
-In fact, by condition $(ii)$, we have $\sigma_{2}(\mu|1)=\mu_{2}\mu_{3}+\mu_{2}\mu_{4}+\mu_{3}\mu_{4}>1$.
-$\sigma_{2}(\mu|1)>0$ and $\mu\in\Gamma_{2}$ implies $\mu_{2}+\mu_{3}>0$ and then $\mu_{4}>\frac{1-\mu_{2}\mu_{3}}{\mu_{2}+\mu_{3}}$.
-Then we get
-$\arctan\mu_{4}>\arctan\frac{1-\mu_{2}\mu_{3}}{\mu_{2}+\mu_{3}}=\frac{\pi}{2}-\arctan\mu_{2}-\arctan\mu_{3},$
-which proves (72) and hence 0 is the subsolution of the critical LYZ equation. ∎
-We next show condition $(i)$ is necessary.
+Indeed, condition $(ii)$ gives $\sigma_{2}(\mu|1)=\mu_{2}\mu_{3}+\mu_{2}\mu_{4}+\mu_{3}\mu_{4}>1$.
+Since $\sigma_{2}(\mu|1)>0$ and $\mu\in\Gamma_{2}$, it follows that $\mu_{2}+\mu_{3}>0$, and consequently
+$\mu_{4}>\frac{1-\mu_{2}\mu_{3}}{\mu_{2}+\mu_{3}}.$
+Therefore,
+$\arctan\mu_{4}>\arctan\left(\frac{1-\mu_{2}\mu_{3}}{\mu_{2}+\mu_{3}}\right)=\frac{\pi}{2}-\arctan\mu_{2}-\arctan\mu_{3},$
+which proves (73). Hence, 0 is a subsolution of the critical LYZ equation. ∎
+We next show that condition $(i)$ is necessary.
 Lemma 5.1.
-Let $u$ be the unique smooth solution of the Hessian quotient equation (71). Then $\chi$ satisfies condition $(i)$.
+Let $u$ be the unique smooth solution to the Hessian quotient equation (72). Then $\chi$ satisfies condition $(i)$.
 Proof.
-The equality in condition $(i)$ follows from the integration of the equation. Since $\sigma_{1}(\chi_{u})>0$, we have
+The equality in condition $(i)$ follows directly from integrating the equation. Since $\sigma_{1}(\chi_{u})>0$, we have
 $\int_{M}\chi\wedge\omega^{3}=\int_{M}\chi_{u}\wedge\omega^{3}>0.$ (74)
-Let $\lambda=(\lambda_{1},\lambda_{2},\lambda_{3},\lambda_{4})$ be the eigenvalues of $\chi_{u}$ with respect to $\omega$ such that $\lambda_{1}\geq\lambda_{2}\geq\lambda_{3}\geq\lambda_{4}$.
-By the MacLarin inequality, we have
+Let $\lambda=(\lambda_{1},\lambda_{2},\lambda_{3},\lambda_{4})$ be the eigenvalues of $\chi_{u}$ with respect to $\omega$, ordered such that $\lambda_{1}\geq\lambda_{2}\geq\lambda_{3}\geq\lambda_{4}$.
+By the MacLaurin inequality, we obtain
 $\sigma_{2}\geq 6\frac{\sigma_{3}}{\sigma_{1}}=6.$ (75)
-Since $\lambda_{4}\sigma_{1}(\lambda|4)=\lambda_{4}(\lambda_{1}+\lambda_{2}+\lambda_{3})\leq\lambda_{1}\lambda_{2}+\lambda_{2}\lambda_{3}+\lambda_{1}\lambda_{3}=\sigma_{2}(\lambda|4)$, we get
-$\sigma_{2}=\sigma_{2}(\lambda|4)+\lambda_{4}\sigma_{1}(\lambda|1)\leq 2\sigma_{2}(\lambda|4),$ (75)
-and then $\sigma_{2}(\lambda|4)\geq 3.$ Since $\lambda_{4}=\frac{\sigma_{3}(\lambda|4)-\sigma_{1}(\lambda|4)}{-\sigma_{2}(\lambda|4)+1}$, we get
+The inequality $\lambda_{4}\sigma_{1}(\lambda|4)\leq\sigma_{2}(\lambda|4)$ yields
+$\sigma_{2}=\sigma_{2}(\lambda|4)+\lambda_{4}\sigma_{1}(\lambda|4)\leq 2\sigma_{2}(\lambda|4),$ (76)
+and consequently,
+$\sigma_{2}(\lambda|4)\geq 3.$ (77)
+Using the identity $\lambda_{4}=\dfrac{\sigma_{3}(\lambda|4)-\sigma_{1}(\lambda|4)}{-\sigma_{2}(\lambda|4)+1}$, we compute
 $\sigma_{4}-\sigma_{2}+1$ $=\lambda_{4}\sigma_{3}(\lambda|4)-\lambda_{4}\sigma_{1}(\lambda|4)-\sigma_{2}(\lambda|4)+1$
-$=$ $\frac{\Big(\sigma_{3}(\lambda|4)-\sigma_{1}(\lambda|4)\Big)^{2}}{-\sigma_{2}(\lambda|4)+1}-\sigma_{2}(\lambda|4)+1$
+$=\frac{\bigl(\sigma_{3}(\lambda|4)-\sigma_{1}(\lambda|4)\bigr)^{2}}{-\sigma_{2}(\lambda|4)+1}-\sigma_{2}(\lambda|4)+1$
 $\leq-2.$
-Hence we get
+Therefore,
 $\mathrm{Re}\int_{M}(\chi+\sqrt{-1}\omega)^{4}$ $=\mathrm{Re}\int_{M}(\chi_{u}+\sqrt{-1}\omega)^{4}$
 $=\int_{M}(\sigma_{4}-\sigma_{2}+1)\,\omega^{4}<0,$
-which proves the second inequality in condition $(i)$. ∎
+which establishes the second inequality in condition $(i)$. ∎

### v3

@@
 Keywords: critical LYZ equation; subsolution; uniform estimates; blow-up argument
+footnotetext: 2020 Mathematics Subject Classification. Primary 32W20; Secondary 53C55.
 1. Introduction
@@
 The following lemma collects some key properties of these cones.
-Lemma 2.2 (Yuan [44], Wang–Yuan [42]).
+Lemma 2.2 (Yuan [43], Wang-Yuan [41]).
 Assume that $\theta(\lambda)\geq\tau$ for some $\tau\in[(n-2)\frac{\pi}{2},n\frac{\pi}{2})$, and that the components of $\lambda=(\lambda_{1},\dots,\lambda_{n})$ satisfy $\lambda_{1}\geq\lambda_{2}\geq\cdots\geq\lambda_{n}$.
@@
 Let $\Omega\subset\mathbb{C}^{n}$ be a bounded domain.
-Suppose $w$ is an (n1)(n-1)-admissible function and a viscosity subsolution of (45) in $\Omega$, and $v$ is a viscosity solution of (45) in $\Omega$.
+Suppose $w$ is an (n1)(n-1)-admissible function and a viscosity subsolution of (45) in $\Omega$, and $v$ is a viscosity supersolution of (45) in $\Omega$.
 If $w\leq v$ on $\partial\Omega$, then $w\leq v$ in $\Omega$.
@@
 This implies that the function $w+\epsilon|z|^{2}-v$ attains its maximum at an interior point $z_{0}\in\Omega$.
-Now, since $w$ is (n1)(n-1)-admissible, the function $w+\epsilon|z|^{2}$ is also (n1)(n-1)-admissible.
-As $v$ is a viscosity solution (and hence a supersolution), by the definition of the supersolution, we have
+Now, since $w$ is (n1)(n-1)-admissible, the function $w+\epsilon|z|^{2}$ is also (n1)(n-1)-admissible. As $v$ is a viscosity supersolution, by the definition of the supersolution, we have
 $\sigma_{n-1}((w_{\alpha\bar{\beta}}(z_{0}))+\epsilon I_{n})+\sigma_{n}((w_{\alpha\bar{\beta}}(z_{0}))+\epsilon I_{n})\leq 0.$ (47)
@@
   3. (iii)
-    The function $\frac{1}{2}v^{2}$ is a subsolution to the same equation.
+    If $0\leq v\leq 1$, $\frac{1}{2}v^{2}$ is a subsolution to the same equation.
   4. (iv)
@@
 Lemma 4.5.
-Let $v$ be a weak solution (respectively, subsolution) in the sense of currents to the equation $\sigma_{n}+\sigma_{n-1}=0$. Then $v$ is also a viscosity solution (respectively, subsolution).
+Let $v$ be a weak subsolution in the sense of currents to the equation $\sigma_{n}+\sigma_{n-1}=0$. Then $v$ is also a viscosity subsolution.
 Proof.
@@
 we conclude that $v$ is a viscosity subsolution.
-The proof for supersolutions follows similarly, using the corresponding definition of viscosity supersolutions. ∎
 4.3. A Liouville theorem to the equation $\sigma_{n}+\sigma_{n-1}=0$ in $\mathbb{C}^{n}$
@@
 Assume there exists a positive constant $C$ such that $\|v\|_{C^{1}(\mathbb{C}^{n})}\leq C$ and $\Delta v\leq C$ in the weak sense (see Remark 4.2 for details).
-If $v$ is a weak solution in the sense of currents to the equation
+If $v$ is a weak solution in the sense of currents and a viscosity supersolution to the equation
 $\sigma_{n-1}(\sqrt{-1}\partial\bar{\partial}v)+\sigma_{n}(\sqrt{-1}\partial\bar{\partial}v)=0,$ (51)
-then $v$ must be constant.
+then $v$ must be a constant.
 Remark 4.2.
@@
 $\sigma_{n-1}\left((\hat{u}_{\alpha\bar{\beta}}^{t_{i}})_{n\times n}+A^{i}\right)=M_{i}^{2}\tan\hat{\theta}(t_{i})\sigma_{n}\left((\hat{u}_{\alpha\bar{\beta}}^{t_{i}})_{n\times n}+A^{i}\right)+O(M_{i}^{-2}),$ (64)
-and by Lemma 2.2, we have
+where the matrix $A^{i}$ is uniformly bounded and by Lemma 2.2, we have
 $\sigma_{k}\left((\hat{u}_{\alpha\bar{\beta}}^{t_{i}})_{n\times n}+A^{i}\right)>0\quad\text{for any }1\leq k\leq n-1.$ (65)
@@
 This proves that $\Delta u^{\infty}\leq C$ in the weak sense.
+Next, we prove $u^{\infty}$ is a supersolution to the equation $\sigma_{n-1}+a_{0}\sigma_{n}=0$ in the viscosity sense.
+For any $C^{2}$ function $\varphi$ with $\lambda(\varphi_{\alpha\bar{\beta}})\in\overline{\Gamma}_{n-1}$ and for any $z_{0}$ of the local maximum point of $\varphi-u^{\infty}$, we need to show
+$\sigma_{n-1}\big(\varphi_{\alpha\bar{\beta}}(z_{0})\big)+a_{0}\sigma_{n}\big(\varphi_{\alpha\bar{\beta}}(z_{0})\big)\leq 0.$ (66)
+If $\sigma_{n-1}(\varphi_{\alpha\bar{\beta}}(z_{0}))=0$, then $\sigma_{n}(\varphi_{\alpha\bar{\beta}}(z_{0}))\leq 0$ and thus (66) holds.
+So we assume $c_{0}=\sigma_{n-1}(\varphi_{\alpha\bar{\beta}}(z_{0}))>0$. Then for any $\epsilon>0$ small enough, we have
+$\lambda(\varphi_{\alpha\bar{\beta}}-2\epsilon\delta_{\alpha\beta})\in\Gamma_{n-1}\ \ \text{and}\ \ \sigma_{n-1}(\varphi_{\alpha\bar{\beta}}-2\epsilon\delta_{\alpha\beta})>\frac{c_{0}}{2}\ \text{in}\ \overline{B}_{\epsilon}(z_{0}).$ (67)
+For sufficiently large $i$ such that $|\hat{u}^{t_{i}}-u^{\infty}|<\frac{\epsilon^{3}}{2}$, consider the function
+$\Phi=\varphi-\epsilon|z-z_{0}|^{2}-\hat{u}^{t_{i}}.$
+Since for any $z\in\partial B_{\epsilon}(z_{0})$,
+$\Phi(z)=\varphi(z)-\hat{u}^{t_{i}}(z)-\epsilon^{3}<$ $\varphi(z)-u^{\infty}(z)-\frac{\epsilon^{3}}{2}$
+$\leq$ $\varphi(z_{0})-u^{\infty}(z_{0})-\frac{\epsilon^{3}}{2}<\varphi(z_{0})-\hat{u}^{t_{i}}(z_{0})=\Phi(z_{0}).$
+then we know that $\Phi$ attains its maximum on $\overline{B}_{\epsilon}(z_{0})$ at an interior point $\tilde{z}_{0}\in B_{\epsilon}(z_{0})$.
+Then $(\Phi_{\alpha\bar{\beta}}(\tilde{z}_{0}))_{n\times n}\leq 0$, namely $(\varphi_{\alpha\bar{\beta}}(\tilde{z}_{0})-\epsilon\delta_{\alpha\beta})_{n\times n}\leq(\hat{u}^{t_{i}}_{\alpha\bar{\beta}}(\tilde{z}_{0}))_{n\times n}$.
+Thus $(\hat{u}^{t_{i}}_{\alpha\bar{\beta}}(\tilde{z}_{0})+A^{i}_{\alpha\bar{\beta}}(\tilde{z}_{0}))_{n\times n}\geq(\varphi_{\alpha\bar{\beta}}(\tilde{z}_{0})-2\epsilon\delta_{\alpha\beta})_{n\times n}$ and we obtain
+$\sigma_{n-1}(\hat{u}^{t_{i}}_{\alpha\bar{\beta}}(\tilde{z}_{0})+A_{\alpha\bar{\beta}}(\tilde{z}_{0}))\geq\sigma_{n-1}\big(\varphi_{\alpha\bar{\beta}}(\tilde{z}_{0})-2\epsilon\delta_{\alpha\beta}\big)\geq\frac{c_{0}}{2}.$
+By the concavity of $\frac{\sigma_{n}}{\sigma_{n-1}}$ in $\Gamma_{n-1}$, we get
+$1-\tan\hat{\theta}(t_{i})M^{2}_{i}\frac{\sigma_{n}\big(\varphi_{\alpha\bar{\beta}}(\tilde{z}_{0})-2\epsilon\delta_{\alpha\beta}\big)}{\sigma_{n-1}\big(\varphi_{\alpha\bar{\beta}}(\tilde{z}_{0})-2\epsilon\delta_{\alpha\beta}\big)}$
+$\leq$ $1-\tan\hat{\theta}(t_{i})M^{2}_{i}\frac{\sigma_{n}(\hat{u}^{t_{i}}_{\alpha\bar{\beta}}(\tilde{z}_{0})+A^{i}_{\alpha\bar{\beta}}(\tilde{z}_{0}))}{\sigma_{n-1}(\hat{u}^{t_{i}}_{\alpha\bar{\beta}}(\tilde{z}_{0})+A^{i}_{\alpha\bar{\beta}}(\tilde{z}_{0}))}$
+$\leq$ $c_{0}^{-1}CM_{i}^{-2},$ (68)
+where in the last inequality we use (64). Then we get
+${\sigma_{n-1}\big(\varphi_{\alpha\bar{\beta}}(\tilde{z}_{0})-2\epsilon\delta_{\alpha\beta}\big)}-\tan\hat{\theta}(t_{i})M^{2}_{i}{\sigma_{n}\big(\varphi_{\alpha\bar{\beta}}(\tilde{z}_{0})-2\epsilon\delta_{\alpha\beta}\big)}\leq c_{0}^{-1}CM_{i}^{-2}\sup|D^{2}\varphi|.$ (69)
+Let $i\rightarrow\inftyand$ then $\epsilon\rightarrow 0$ in the above, we obtain (66). Hence $u^{\infty}$ is a supersolution to the equation $\sigma_{n}+\sigma_{n-1}=0$.
 We can then apply the Liouville Theorem 4.1 to $a_{0}^{-1}u^{\infty}$ to conclude that $a_{0}^{-1}u^{\infty}$ is constant, which again leads to a contradiction.